### v2

# Fundamental limitations on the recoverability of quantum processes

Abstract
Quantum information processing and computing tasks can be understood as quantum networks, comprising quantum states and channels and possible physical transformations on them. It is hence pertinent to estimate the change in informational content of quantum processes due to physical transformations they undergo. The physical transformations of quantum states are described by quantum channels, while the transformations of quantum channels are described by quantum superchannels. In this work, we determine fundamental limitations on how well the physical transformation on quantum channels can be undone or reversed, which are of crucial interest to design and benchmark quantum information and computation devices. In particular, we refine (strengthen) the quantum data processing inequality for quantum channels under the action of quantum superchannels. We identify a class of quantum superchannels, which appears to be the superchannel analogue of subunital quantum channels, under the action of which the entropy of an arbitrary quantum channel is nondecreasing. We also provide a refined inequality for the entropy change of quantum channels under the action of an arbitrary quantum superchannel.
I Introduction
I.1 Motivation and background
Quantum theory lies at the core of our understanding of the natural phenomena at the small scale and plays a fundamental role in the rapidly growing area of quantum information processing and computing. Traditionally, quantum states are used as the information carriers and they need to be appropriately manipulated to perform the desired protocols or algorithms, e.g., quantum key distribution XMZ+ (20); DBWH (21) and protocols implying the advantage in quantum computation Sho (94); Gro (96); BGK (18); these manipulations are quantum state transformations. The physical transformations of quantum states of systems are described by completely positive (CP), trace preserving (TP) maps in quantum theory and are referred to as quantum channels. Achieving desired manipulations of quantum states require precise application of specific quantum channels. A thorough understanding of quantum channels, therefore, is not only essential for the theoretical insights but also pivotal for the realization of quantum computers, sensing and information processing devices, and quantum Internet NC (00); DM (03); Hol (12); SSD+ (15); Wil (13); SSHD (23).
However, with the rapid increase in our capabilities to handle and manipulate small scale systems, there is a resounding global effort to realize existing quantum technologies in laboratories and invent new ones. This necessitates a convenient and overarching framework to handle quantum states, measurements, and quantum channels succinctly and efficiently in a unified way. Quantum networks comprising quantum states, measurements, and quantum channels provide such a framework, where information can be encoded in quantum channels and/or quantum states depending on the desired tasks SNB+ (06); Das (19); DW19b ; PLY+ (23); then the manipulation of information can be treated in a unified way by the higher order transformations of quantum networks. Considering quantum networks as elementary quantum objects, their transformations would then describe the most general kind of physical transformations of quantum information. Examples of transformations of quantum channels include among others the programmable quantum channels (taking quantum state as input and outputting a quantum channel) DP (05); GBW (21) and the cloning of unknown unitaries CDP08c . The transformations of quantum channels are achieved by the action of quantum superchannels–linear transformations that map quantum channels into quantum channels CDP08a ; CDP08b ; Życ (08); Per (17); BP (19). In this work, we shed light on the irreversibility of the transformation of quantum channels under the action of quantum superchannels. In particular, we derive entropic inequalities that provide non-trivial lower bounds on the entropy change and data processing inequality pertaining to the transformation of quantum channels under the action of quantum superchannels.
The entropic functionals attributed to quantum channels, together with their inherent properties, furnishes both qualitative and quantitative methodologies to analyse quantum processes from the perspective of quantum information theory. Entropic inequalities, like uncertainty relations CBTW (17), second laws of thermodynamics BHN+ (15), data processing inequality Rus (02), etc. have been instrumental in devising quantum communication and computation protocols. The quantum relative entropy $D\left(\cdot\;\middle\|\;\cdot\right)$, a quantum generalization of Kullback-Leibler divergence, is one of the core entropic functions and is indispensable to analysis of distinguishability problems in the quantum information theory. $D\left(\rho\;\middle\|\;\sigma\right)$ between a state $\rho$ of an arbitrary dimensional quantum system, finite or infinite, and a positive semidefinite operator $\sigma$ is defined as Ume (62) [^1] The logarithm can be taken with respect to the base $2$ without loss of generality.
$D\left(\rho\;\middle\|\;\sigma\right):=\begin{cases}\operatorname{tr}(\rho(\log{\rho}-\log{\sigma}))&\text{if}\ \operatorname{supp}(\rho)\subseteq\operatorname{supp}(\sigma),\\ +\infty&\text{otherwise}.\end{cases}$ (1)
[…]
$D\left(\rho\;\middle\|\;\sigma\right)-D\left(\mathcal{N}(\rho)\;\middle\|\;\mathcal{N}(\sigma)\right)\geq 0.$ (2)
The quantum data processing inequality has several ramifications in the quantum information theory and is one of the widely invoked inequalities to derive limitations in quantum processing and computing tasks Rus (02); Hol (12); BSF+ (17); DBWH (21); SSC (23); KW (24). A consequence of the above inequality is a second law of thermodynamics like inequality which states that the von Neumann entropy  [^2] We simply refer to the von Neumann entropy functional as the entropy. We also discuss generalized entropy functions. The use of the entropy for the von Neumann entropy will be clear from the context. of a quantum state is nondecreasing under the action of a unital channel AU (77, 82), i.e., $S(\mathcal{N}(\rho))-S(\rho)\geq 0$ for any quantum channel $\mathcal{N}$ that preserves identity operator, i.e., $\mathcal{N}(\mathbbm{1})=\mathbbm{1}$, and arbitrary quantum state $\rho$. It is known that the von Neumann entropy of any quantum state is also nondecreasing for subunital channels AU (77, 82), meaning that $S(\mathcal{N}(\rho))-S(\rho)\geq 0$ for quantum channels $\mathcal{N}$ satisfying $\mathcal{N}(\mathbbm{1})\leq\mathbbm{1}$. The input and output dimensions of unital channels are the same whereas the output dimension of a subunital channel is greater than or equal to the input dimension.
The quantum data processing inequality (Eq. (2)) for states is saturated, i.e., $D\left(\rho\;\middle\|\;\sigma\right)=D\left(\mathcal{N}(\rho)\;\middle\|\;\mathcal{N}(\sigma)\right)$, if and only if there exists a completely positive, trace preserving map $\mathcal{P}$, called the Petz recovery map, such that $\mathcal{P}\circ\mathcal{N}(\rho)=\rho$ and $\mathcal{P}\circ\mathcal{N}(\sigma)=\sigma$ Pet (86, 88). This result has found crucial applications in quantum error correction BK (02); MN (12) among others. The quantum data processing inequality has been recently improved to the following. For arbitrary quantum states $\rho$ and $\sigma$, and an arbitrary quantum channel $\mathcal{N}$, there exists a universal recovery map $\mathcal{P}^{\mathrm{R}}_{\sigma,\mathcal{N}}$ (a completely positive and trace nonincreasing map that depends only on $\sigma$ and $\mathcal{N}$ with no dependence on $\rho$) such that JRS+ (18)
$D\left(\rho\;\middle\|\;\sigma\right)-D\left(\mathcal{N}(\rho)\;\middle\|\;\mathcal{N}(\sigma)\right)\geq-\log F\left(\rho,\mathcal{P}^{\mathrm{R}}_{\sigma,\mathcal{N}}\circ\mathcal{N}(\rho)\right)\geq 0,$ (3)
where $F(\rho,\sigma):=\norm{\sqrt{\rho}\sqrt{\sigma}}_{1}^{2}$ is the fidelity between $\rho$ and $\sigma$. $\norm{\mathsf{X}}_{1}=\operatorname{tr}{\sqrt{\mathsf{X}^{\ast}\mathsf{X}}}$ is the trace-norm of $\mathsf{X}$. There is also a strengthened (refined) inequality BDW (16) for entropy change of an arbitrary state under the action of an arbitrary positive and trace preserving map $\mathcal{N}$,
$S\left(\mathcal{N}(\rho)\right)-S(\rho)\geq D\left(\rho\;\middle\|\;\mathcal{N}^{\ast}\circ\mathcal{N}(\rho)\right),$ (4)
where $\mathcal{N}^{\ast}$ denotes the adjoint map of the linear map $\mathcal{N}$. The above inequality is a strengthened version of entropy change under quantum channels as quantum channels are special cases of positive and trace preserving maps. Notice that $D\left(\rho\;\middle\|\;\mathcal{N}^{\ast}\circ\mathcal{N}(\rho)\right)\geq 0$ for all subunital channels $\mathcal{N}$ and quantum states $\rho$, thus $\mathcal{N}^{\ast}$ can be deemed as a universal recovery map (channel) for unital channels $\mathcal{N}$ in the context of entropy change.
The relative entropy between quantum channels CMW (16); LKDW (18); Yua (19) is also defined in a way similar to the diamond norm Wat (04) of two quantum channels where the trace-distance is replaced by the relative entropy. Such a notion of relative entropy between quantum channels is nondecreasing under the action of quantum superchannels LKDW (18); Hir (23). Motivated by the definition of the entropy for quantum states, Gou19a proposed axioms that the definition of the entropy of quantum channel should satisfy. In Yua (19); GW (21), the definition of quantum entropy was introduced which satisfies the physically motivated axioms mentioned in Gou19a . This notion of the entropy of a quantum channel finds operational meaning in the information-theoretic tasks of hypothesis testing of quantum channels Yua (19) and quantum channel merging GW (21).
I.2 Our contribution and main results
[…]
We introduce $\mathcal{R}$-subpreserving supermaps $\hat{\Theta}:\mathcal{L}(A,B)\to\mathcal{L}(C,D)$ that are defined as supermap such that $\hat{\Theta}(\mathcal{R}_{A\to B})\leq\mathcal{R}_{C\to D}$, where $\mathcal{R}_{A\to B}$ and $\mathcal{R}_{C\to D}$ are completely depolarizing maps (Definition 2). The entropy $S[\mathcal{N}]$ of an arbitrary quantum channel $\mathcal{N}_{A\to B}$ is then defined as $S[\mathcal{N}]:=-D\left[\mathcal{N}\;\middle\|\;\mathcal{R}\right]$, i.e., negative of the relative entropy between a quantum channel $\mathcal{N}_{A\to B}$ and the completely deploarizing map $\mathcal{R}_{A\to B}$ (see Definition 5). We prove that the entropy of a quantum channel is nondecreasing under the action of an $\mathcal{R}$-subpreserving superchannel $\hat{\Theta}:\mathcal{L}(A,B)\to\mathcal{L}(C,D)$ (Proposition 8), i.e.,
$S[\hat{\Theta}({\mathcal{N}})]-S[\mathcal{N}]\geq 0.$ (5)
[…]
We show that the quantum relative entropy of quantum channels is nonincreasing under the action of a wide class of CP-preserving and TP-preserving supermaps (see Proposition 2). We strengthen the data processing inequality for quantum channels by deriving nontrivial lower bound to the monotonicity of relative entropy of quantum channels under the action of quantum superchannels. That is, the following lower bound holds for the monotonicity of the relative entropy $D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]$ between quantum channels $\mathcal{N}$ and $\mathcal{M}$ with a desirable property under the action of a quantum superchannel $\Theta$,
$D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]-D\left[\Theta\left(\mathcal{N}\right)\;\middle\|\;\Theta\left(\mathcal{M}\right)\right]\geq-\log F\left(\mathsf{C}^{\Psi}_{\mathcal{N}},\left(\mathcal{P}^{\mathrm{R}}\circ\mathfrak{T}^{\prime}\right)\left(\mathsf{C}^{\Psi}_{\mathcal{N}}\right)\right)\geq 0,$ (6)
where the fidelity $F$ is taken between the operators related to the quantum channel $\mathcal{N}$ and the operators related to the reversibility of the action of the quantum superchannel $\Theta$ on $\mathcal{N}$ and $\mathcal{M}$. For a precise statement and a condition on the pair of channels $\mathcal{N}$ and $\mathcal{M}$ for which the above bound holds, see Theorem 2. An important consequence of this result is that for a specific class of quantum channels and a quantum superchannel that preserves this class, the saturation of the data processing inequality implies the existence of a completely CP-preserving supermap that undoes (or reverses) the action of the superchannel on given quantum channels, see Remark 5.
Noticing the similarity between the definitions of entropy for quantum states and quantum channels, we define the generalized divergence and the entropy of quantum superchannels, and subsequently generalize these notions to higher-order quantum processes. By defining the notion of a completely depolarizing supermap, we show that the entropy functional for quantum superchannels is nondecreasing under a quantum super-superchannel that subpreserves the completely depolarizing supermap. Then we prove the subadditivity of the quantum superchannel entropy under tensor products. Further, we show that for a replacer superchannel $\Theta_{0}$, which maps every channel to a fixed channel $\mathcal{N}_{0}$, the entropy of $\Theta_{0}$ is upper bounded by the entropy of $\mathcal{N}_{0}$.
I.3 Structure of the paper
The structure of this paper is as follows. In Section II, we fix our notations and present the preliminaries and background necessary to derive the main results of this paper. In Section II, we discuss some properties of the space of CP maps, quantum channels, and superchannels; and define the notion of representing maps for superchannels. In Section III, we define the entropy of a quantum channel via an axiomatic approach and we also extend the definition of entropy for quantum channels with infinite-dimensional input of output space. We also prove several bounds on channel entropy gain for a wide class of channels including tele-covariant channels. In Section IV, we introduce the notion of the sufficiency of superchannels and we prove a stronger version of data processing inequality and entropy gain of quantum channels, with a remainder term which depends on recovery supermap. In Section V, we define the generalized divergence and entropy functional for superchannels and we discuss some properties of both the functionals. Finally, in Section VI, we give the concluding remarks of our findings. We also define a divergence functional for the higher order process, and provide a necessary condition for the functional to be a generalized divergence. Using this definition of generalized divergence, we define the entropy functional for higher-order processes.
II Preliminaries: Standard notations, definitions, and facts
In this section, we briefly mention some standard definitions and notations that are important for deriving and discussing our results.
The separable Hilbert space associated with a quantum system $A$ is denoted by $A$ and its dimension is denoted by $|A|:={\mathrm{d}im}(A)$. The Hilbert space of a joint system $AB$ is given by the tensor product $A\otimes B$ of the separable Hilbert spaces $A$ and $B$. Let $A\simeq B$ denote that the two Hilbert spaces $A$ and $B$ are isomorphic. The space of all linear operators acting on $A$ is denoted by ${\mathcal{L}}(A)$ and the identity operator on $A$ is denoted by $\mathbbm{1}_{A}$. The space $\mathcal{L}\left(A\right)$ is equipped with the Hilbert-Schmidt inner product $\langle\mathsf{P},\mathsf{Q}\rangle\equiv\mathrm{tr}\left({\mathsf{P}}^{{\dagger}}\mathsf{Q}\right)$ $\forall~{}\mathsf{P},\mathsf{Q}\in\mathcal{L}\left(A\right)$. Let $\mathcal{L}\left(A\right)_{+}\subset\mathcal{L}\left(A\right)$ denote the set of all positive semidefinite operators and $\mathcal{L}_{h}\left(A\right)$ denote the subset containing all the hermitian operators in $\mathcal{L}\left(A\right)$. There is a natural partial ordering in $\mathcal{L}_{h}\left(A\right)$ as follows: for operators $\{\mathsf{P},\mathsf{Q}\}\in\mathcal{L}_{h}\left(A\right)$, we say that $\mathsf{P}\geq\mathsf{Q}$ if and only if $(\mathsf{P}-\mathsf{Q})\in\mathcal{L}\left(A\right)_{+}$. The state of a quantum system is described by a positive semidefinite trace-class operator with unit trace, called density operator, and is denoted by $\rho$. Let $\mathcal{D}\left(A\right)$ and $\mathcal{D}(A\otimes B)$ denote the set of all density operators on $A$ and on $A\otimes B$, respectively. The rank-one elements of $\mathcal{D}\left(A\right)$ are called the pure states. With a slight abuse of notation, we may call a normalized vector |$\ket{\psi}\in A$ as a (pure) state whenever it is clear that we are referring to the corresponding rank-one density operator $\psi_{A}:=\outerproduct{\psi}{\psi}_{A}\in\mathcal{D}(A)$. For the sake of simplicity, we may denote an operator $\rho\in\mathcal{L}(A\otimes B)$ as $\rho_{AB}$. For an operator $\rho_{AB}\in\mathcal{L}(A\otimes B)$, $\rho_{A}\in\mathcal{L}(A)$ is called the marginal operator for the system $A$ (or reduced operator of the system $A$) and is defined as $\rho_{A}=\operatorname{tr}_{B}(\rho_{AB})$. Similarly, one defines $\rho_{B}=\operatorname{tr}_{A}(\rho_{AB})$. If $\rho_{AB}\in\mathcal{D}(A\otimes B)$, then the marginals $\rho_{A}\in\mathcal{D}(A)$ and $\rho_{B}\in\mathcal{D}(B)$ are called reduced states of $\rho_{AB}$. Let $\mathsf{FRank}\left(A\otimes B\right)\subset\mathcal{L}\left(A\otimes B\right)$ be a set of rank one operators $\Psi_{AB}\in\mathcal{L}\left(A\otimes B\right)$ such that $\mathrm{rank}\left(\Psi_{A}\right)=|A|$, where $A\simeq B$. The maximally entangled operator |$\outerproduct{\Psi^{+}}{\Psi^{+}}_{AB}:=\Psi^{+}_{AB}\in\mathcal{L}(A\otimes B)_{+}$ is defined via |$\ket{\Psi^{+}}_{AB}:=\sum_{i=1}^{d}\ket{i,i}_{AB}$, where $d=\min\{|A|,|B|\}$, and we call the normalized maximally entangled operator $d^{-1}\Psi^{+}_{AB}\in\mathcal{D}(A\otimes B)$ as a maximally entangled state of $AB$.
Let $\mathcal{L}\left(A,B\right)$ be the space of all linear maps from $\mathcal{L}\left({A}\right)$ to $\mathcal{L}\left({B}\right)$. We denote a linear map $\mathcal{N}\in\mathcal{L}\left(A,B\right)$ as $\mathcal{N}_{A\rightarrow B}$. We will use both notations $\mathcal{N}_{A\to B}$ and $\mathcal{N}:\mathcal{L}(A)\to\mathcal{L}(B)$, interchangeably for a linear map $\mathcal{N}$ whenever there is no ambiguity. A linear map $\mathcal{N}_{A\rightarrow B}$ is called positive if it maps elements of $\mathcal{L}\left({A}\right)_{+}$ to the elements of $\mathcal{L}\left({B}\right)_{+}$, and completely positive if $\mathrm{id}_{R}\otimes\mathcal{N}_{A\rightarrow B}$ (where $\mathrm{i}d_{R}$ denotes the identity superoperator acting on a reference system $R$) is a positive map for arbitrary size of the reference system $R$. A map $\mathcal{N}_{A\rightarrow B}$ is called trace preserving if for all $\mathsf{X}\in{\mathcal{L}}({A})$, $\operatorname{tr}\left(\mathcal{N}_{A\rightarrow B}(\mathsf{X})\right)=\operatorname{tr}\left(\mathsf{X}\right)$, and it is called trace nonincreasing if for all $\mathsf{X}\in{\mathcal{L}}(A)$, $\operatorname{tr}\left(\mathcal{N}_{A\rightarrow B}(\mathsf{X})\right)\leq\operatorname{tr}(\mathsf{X})$. A quantum channel $\mathcal{N}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left({B}\right)$ is a linear, completely positive and trace preserving (CPTP) map which describes a physically meaningful evolution of a quantum system. A linear map $\mathcal{N}_{A\rightarrow B}$ is called sub-unital if $\mathcal{N}_{A\rightarrow B}\left(\mathbbm{1}_{{A}}\right)\leq\mathbbm{1}_{{B}}$, unital if $\mathcal{N}_{A\rightarrow B}\left(\mathbbm{1}_{{A}}\right)=\mathbbm{1}_{{B}}$, and super-unital if $\mathcal{N}_{A\rightarrow B}\left(\mathbbm{1}_{{A}}\right)>\mathbbm{1}_{{B}}$. The adjoint $\mathcal{N}^{*}_{B\rightarrow A}$ of a linear map $\mathcal{N}_{A\rightarrow B}$ is the unique linear map that satisfies the following equation:
$\langle\mathsf{P},\mathcal{N}_{A\rightarrow B}\left(\mathsf{Q}\right)\rangle=\langle\mathcal{N}_{B\rightarrow A}^{*}\left(\mathsf{P}\right),\mathsf{Q}\rangle,$ (7)
for all $\mathsf{Q}\in\mathcal{L}\left({A}\right)$ and $\mathsf{P}\in\mathcal{L}\left({B}\right)$. The adjoint map of a quantum channel is a completely positive and unital map. For a trace nonincreasing map, the adjoint map is sub-unital while for a trace nondecreasing map, the adjoint map is super-unital. The trace norm, Hilbert-Schmidt norm, and the operator norm for an operator $\mathsf{X}\in\mathcal{L}(A)$ are defined as $\norm{\mathsf{X}}_{1}:=\mathrm{tr}\left(\sqrt{\mathsf{X}^{\dagger}\mathsf{X}}\right)$, $\norm{\mathsf{X}}_{2}:=\sqrt{\mathrm{tr}\left(\mathsf{X}^{\dagger}\mathsf{X}\right)}$, and $\norm{\mathsf{X}}_{\infty}:=\sup_{x\in A,x\neq 0}\frac{\norm{\mathsf{X}(x)}}{\norm{x}}$, where $\norm{\cdot}$ denotes the norm on $A$ induced by the inner product on $A$.
Definition 1 (Tele-covariant channels Hol (02)).
(see also (DBWH, 21, Sec. V.C.)) Let $G$ be a finite group and let $\{U_{g}\}_{g}$ and $\{V_{g}\}_{g}$ be unitary representations of $G$ in the Hilbert spaces $A$ and $B$, respectively. A quantum channel $\mathcal{N}_{A\rightarrow B}$ is called covariant with respect to $G$ if for all $\rho\in\mathcal{D}(A)$,
$\mathcal{N}_{A\rightarrow B}\circ\mathcal{U}_{g}=\mathcal{V}_{g}\circ\mathcal{N}_{A\rightarrow B}\hskip 5.69046pt\forall g\in G,$ (8)
where $\mathcal{U}_{g}(\cdot):=U_{g}(\cdot)U_{g}^{\ast}$ and $\mathcal{V}_{g}(\cdot):=V_{g}(\cdot)V_{g}^{\ast}$. The quantum channel $\mathcal{N}_{A\rightarrow B}$ is called tele-covariant if it is covariant with respect to $G$ with unitary representation that satisfies $|G|^{-1}\sum_{g}U_{g}\left(\mathsf{X}\right)U^{*}_{g}=\frac{\operatorname{tr}(\mathsf{X})}{|A|}\mathbbm{1}_{A}$ for all $\mathsf{X}\in\mathcal{L}(A)$.
von Neumann entropy.—The von Neumann entropy (or the entropy) of a quantum state $\rho_{A}\in{\mathcal{D}}(A)$ is a non-negative quantity defined as
$S(\rho_{A})=-\operatorname{tr}({\rho_{A}\log{\rho_{A}}}),$ (9)
where the logarithm is taken with respect to base two.
Generalized divergences.— A map $\mathbf{D}:\mathcal{L}\left(A\right)_{+}\times\mathcal{L}\left(A\right)_{+}\rightarrow\mathbb{R}$ is called generalized state divergence if it satisfies data processing inequality, i.e.,
$\mathbf{D}(\rho\;\|\;\sigma)\geq\mathbf{D}(\mathcal{N}\left(\rho\right)\;\|\;\mathcal{N}\left(\sigma\right)),$ (10)
where $\rho,\sigma\in\mathcal{L}\left(A\right)_{+}$ and $\mathcal{N}:\mathcal{L}(A)_{+}\rightarrow\mathcal{L}(A)_{+}$ is any quantum channel. The quantum relative entropy is one example of generalized divergence. Given a density operator $\rho\in{\mathcal{D}}(A)$ and a positive semidefinite operator $\sigma\in{\mathcal{L}}(A)_{+}$, the relative entropy is defined as Ume (62):
$D(\rho\;\|\;\sigma):=\begin{cases}\operatorname{tr}(\rho(\log{\rho}-\log{\sigma}))&\text{if}\ \operatorname{supp}(\rho)\subseteq\operatorname{supp}(\sigma),\\ +\infty&\text{otherwise}.\end{cases}$ (11)
The CJKS isomorphism.— Let $A$ and $B$ be two finite-dimensional Hilbert spaces, and let $\mathcal{N}_{A\rightarrow B}$ be a linear map. Let $\{e_{ij}\}_{ij}$ be a complete set of matrix units for $\mathcal{L}\left({A}\right)$ and $\{f_{ij}\}_{ij}$ be a complete set of matrix units for $\mathcal{L}\left({B}\right)$. Then the CJKS matrix Stø (13); Jam (74); Cho (75); Kra (83); Sud (85), which we shall refer to as the Choi operator, for $\mathcal{N}_{A\rightarrow B}$ is defined as the operator $\widehat{\mathsf{C}}_{\mathcal{N}}=\sum_{{i,j}=1}^{|A|}e_{ij}\otimes\mathcal{N}_{A\rightarrow B}\left(e_{ij}\right)$, which belongs to $\mathcal{L}\left({A}\otimes{B}\right)$. Let us further define $\mathsf{C}_{\mathcal{N}}=\widehat{\mathsf{C}}_{\mathcal{N}}/|A|$. If $\mathcal{N}_{A\rightarrow B}$ is a quantum channel then $\mathsf{C}_{\mathcal{N}}$ is a quantum state and we will call it the Choi state of the channel $\mathcal{N}_{A\rightarrow B}$. Consider a map $J_{A,B}:\mathcal{L}(A,B)\rightarrow\mathcal{L}(A\otimes B)$ defined by $J_{A,B}\left(\mathcal{N}_{A\rightarrow B}\right)=\widehat{\mathsf{C}}_{\mathcal{N}}$. One can check that this map is linear and injective. To show that it is also surjective, let us consider an element $\mathsf{X}\in\mathcal{L}(A\otimes B)$ which can be written as $\mathsf{X}=\sum_{i,j,k,l}x_{ijkl}e_{ij}\otimes f_{kl}=\sum_{i,j}e_{ij}\otimes\sum_{k,l}x_{ijkl}f_{kl}.$ Now, $\Gamma_{\mathsf{X}}(e_{ij})=\sum_{k,l}x_{ijkl}f_{kl}$, defines a map $\Gamma_{\mathsf{X}}:\mathcal{L}({A})\rightarrow\mathcal{L}(B)$ by linear extension. So, we can write
$\mathsf{X}=\sum_{{i,j}=1}^{|A|}e_{ij}\otimes\Gamma_{\mathsf{X}}(e_{ij}),$ (12)
which implies that the map $J_{A,B}$ is surjective as well. This map is known as Choi-Jamiołkowski isomorphism. The action of $\Gamma_{\mathsf{X}}$ on an element $\mathsf{Q}\in\mathcal{L}(A)$ can be written as, $\Gamma_{\mathsf{X}}\left(\mathsf{Q}\right)=\mathrm{tr}_{A}\left[\left(\mathsf{Q}^{t}\otimes\mathbbm{1}_{B}\right)\mathsf{X}\right]$, where $t$ represents transposition with respect to the usual canonical basis of $A$ and ${\mathrm{t}r}_{A}$ is partial trace with respect to system $A$. The Choi-Jamiołkowski isomorphism leads to the concept of a channel-state duality. To understand the channel-state duality we need to have a look at the CJKS theorem on completely positive maps.
CJKS theorem on completely positive maps.— The Choi operator $\widehat{\mathsf{C}}_{\mathcal{N}}=\sum_{{i,j}=1}^{|A|}$ $e_{ij}\otimes\mathcal{N}_{A\rightarrow B}(e_{ij})$in $\mathcal{L}(A\otimes B)$ is positive semidefinite if and only if the linear map $\mathcal{N}_{A\rightarrow B}$ is completely positive.
Quantum superchannels.— A linear map $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ is called a supermap which maps elements of $\mathcal{L}\left(A,B\right)$ to elements of $\mathcal{L}\left(C,D\right)$. A supermap $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ is called completely CP-preserving if the map $\mathrm{id}\otimes\Theta:\mathcal{L}\left(A^{\prime},B^{\prime}\right)\otimes\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(A^{\prime},B^{\prime}\right)\otimes\mathcal{L}\left(C,D\right)$ is CP-preserving for for all choices of the Hilbert spaces $A^{\prime}$ and $B^{\prime}$, where $\mathrm{id}$ denotes the identity supermap. A supermap is called TP-preserving if it maps a trace preserving map into a trace preserving map. A quantum superchannel is defined as a completely CP-preserving and TP-preserving supermap. It is clear from the definition that a quantum superchannel maps a quantum channel to a quantum channel.
Let $\mathcal{N}_{A\rightarrow B}$ and $\mathcal{M}_{C\rightarrow D}$ be any two quantum channels related via a superchannel $\Theta$ as $\Theta\left(\mathcal{N}_{A\rightarrow B}\right)=\mathcal{M}_{C\rightarrow D}$. The action of $\Theta$ on $\mathcal{N}_{A\rightarrow B}$ can be physically realized as follows CDP08b ; Gou19a :
$\Theta\left(\mathcal{N}_{A\rightarrow B}\right)=\mathcal{V}_{BR\rightarrow D}\circ\left(\mathcal{N}_{A\rightarrow B}\otimes\mathrm{id}_{R}\right)\circ\mathcal{U}_{C\rightarrow AR},$ (13)
where $\mathcal{V}\in\mathcal{L}\left({BR},D\right)$ is a postprocessing channel, $\mathcal{U}\in\mathcal{L}\left({C},AR\right)$ is a preprocessing channel, and $R$ is some reference system.
Space of quantum channels and superchannels.— Here we discuss the properties of the space of quantum channels and superchannels. We first introduce a notion of inner product on $\mathcal{L}(A,B)$ as the following lemma.
Lemma 1 (Gou19a ).
Let $\{b^{A}_{j}\}_{j}$ be an orthonormal basis for $\mathcal{L}\left({A}\right)$. Then, the space $\mathcal{L}(A,B)$ of linear maps is equipped with the following inner product:
$\mathcal{N}_{A\rightarrow B},\mathcal{M}_{A\rightarrow B}\right>=\sum_{j}\left<\mathcal{N}(b^{A}_{j}),\mathcal{M}(b^{A}_{j})\right>,\forall\hskip 5.69046pt\mathcal{N}_{A\rightarrow B},\mathcal{M}_{A\rightarrow B}\in\mathcal{L}(A,B),$ (14)
where the inner product on the right-hand side is the usual Hilbert-Schmidt inner product. This inner product is independent of the choice of basis.
The above inner product was defined in Gou19a , and it was asserted that the inner product is independent of choice of basis. In Appendix B.1, we provide a simple proof of this statement. For a given pair of bases $\{b^{A}_{i}\}_{i}$ and $\{b^{B}_{j}\}_{j}$, for $\mathcal{L}(A)$ and $\mathcal{L}(B)$, respectively, we can define an orthonormal basis $\{\mathcal{E}^{A\rightarrow B}_{ij}\}$ for the space $\mathcal{L}(A,B)$ as follows
$\mathcal{E}^{A\rightarrow B}_{ij}\left(\mathsf{K}\right)=\left<b^{A}_{i},\mathsf{K}\right>b^{B}_{j}=\mathrm{tr}\left(({b^{A}_{i})}^{\dagger}\mathsf{K}\right)b^{B}_{j},$ (15)
where $\mathsf{K}\in\mathcal{L}\left(A\right)$. It can be verified that the above-defined basis is orthonormal, i.e, $\mathcal{E}^{A\rightarrow B}_{ij},\mathcal{E}^{A\rightarrow B}_{kl}\right>=\delta_{ik}\delta_{jl}$. The adjoint $\Theta^{*}:\mathcal{L}\left({C},D\right)\rightarrow\mathcal{L}\left({A},B\right)$ of a linear map $\Theta:\mathcal{L}\left({A},B\right)\rightarrow\mathcal{L}\left({C},D\right)$ is a linear map that for all $\mathcal{N}_{A\rightarrow B}$ and $\mathcal{N^{\prime}}_{C\rightarrow D}$ satisfies
$\Theta(\mathcal{N}),\mathcal{N^{\prime}}\right>=\left<\mathcal{N},\Theta^{*}(\mathcal{N^{\prime}})\right>,$ (16)
where the inner products in the left and right hand side are in $\mathcal{L}\left({C},D\right)$ and $\mathcal{L}\left({A},B\right)$, respectively.
In Gou19a , it was shown that the Choi-type theorem (CJKS isomorphism) also holds for supermaps. Here we provide an alternative proof that is suitable for our context.
Choi type theorem for supermaps.— Let $\{e^{Q}_{ij}\}_{ij}$ be a canonical basis for the space $\mathcal{L}\left({Q}\right)$, where $Q\in\{A,B,C,D\}$. In the bra-ket notation, we can write $e^{Q}_{ij}=\outerproduct{i_{Q}}{j_{Q}}$ |, where {|$\{\ket{i_{Q}}\}_{i}$ is an orthonormal basis for $Q$. The orthonormal basis for $\mathcal{L}\left({A},B\right)$ is the following,
$\mathcal{E}^{A\rightarrow B}_{ijkl}\left(\mathsf{K}\right)=\left<e^{A}_{ij},\mathsf{K}\right>e^{B}_{kl}=\mathrm{tr}\left({e_{ij}^{A}}^{\dagger}\mathsf{K}\right)e^{B}_{kl}.$ (17)
Let $\mathscr{L}\left[\mathcal{L}\left({A},B\right),\mathcal{L}\left({C},D\right)\right]$ denote the set of all the linear maps which map elements of $\mathcal{L}\left({A},B\right)$ to $\mathcal{L}\left(C,D\right)$. The spaces $\mathscr{L}\left[\mathcal{L}\left({A},B\right),\mathcal{L}\left({C},D\right)\right]$ and $\mathcal{L}\left({A},B\right)\otimes\mathcal{L}\left({C},D\right)$ are isomorphic as vector spaces under the following identification
$\Theta\leftrightarrow\Lambda_{\Theta}:=\sum_{ijkl}\mathcal{E}^{A\rightarrow B}_{ijkl}\otimes\Theta\left(\mathcal{E}^{A\rightarrow B}_{ijkl}\right),$ (18)
where $\Theta\in\mathscr{L}\left[\mathcal{L}\left({A},B\right),\mathcal{L}\left({C},D\right)\right]$ and $\Lambda_{\Theta}\in\mathcal{L}\left({A},B\right)\otimes\mathcal{L}\left({C},D\right)$. It should be noted that the object $\Lambda_{\Theta}$ has a mathematical structure similar to Choi operators. The lemma below supports this similarity further.
Lemma 2 (Gou19a ).
A supermap $\Theta\in\mathscr{L}\left[\mathcal{L}\left({A},B\right),\mathcal{L}\left({C},D\right)\right]$ is completely CP-preserving if and only if $\Lambda_{\Theta}=\sum_{ijkl}\mathcal{E}^{A\rightarrow B}_{ijkl}\otimes\Theta\left(\mathcal{E}^{A\rightarrow B}_{ijkl}\right)\in\mathcal{L}\left({A},B\right)\otimes\mathcal{L}\left({C},D\right)$ is CP.
The above lemma in its present form is not mentioned in Gou19a , but we can arrive in the present form by using (Gou19a, , Theorem 1). See Appendix B.2 for an alternate, detailed proof of the above lemma.
Representing Map.— Let $\Theta\in\mathscr{L}\left[\mathcal{L}\left({A},B\right),\mathcal{L}\left({C},D\right)\right]$ be a supermap. If this map is completely CP-preserving and TP-preserving then it is called a quantum superchannel. The supermap $\Theta\in\mathscr{L}\left[\mathcal{L}\left({A},B\right),\mathcal{L}\left({C},D\right)\right]$ induces a map $\mathsf{T}:\mathcal{L}\left({A\otimes B}\right)\rightarrow\mathcal{L}\left({C\otimes D}\right)$, at the level of Choi operators, and is given by
$\mathsf{T}(\mathsf{X})=\widehat{\mathsf{C}}_{\Theta\left(\Gamma_{\mathsf{X}}\right)},$ (19)
where $\mathsf{X}\in\mathcal{L}\left({A\otimes B}\right)$ is the Choi operator of $\Gamma_{\mathsf{X}}$ and is defined in Eq. (12), and $\widehat{\mathsf{C}}_{\Theta\left(\Gamma_{\mathsf{X}}\right)}$ is Choi operator for map $\Theta\left(\Gamma_{\mathsf{X}}\right)$. Following CDP08b , we will call this map the representing map for the supermap $\Theta$. The spaces $\mathscr{L}\left[\mathcal{L}\left({A},B\right),\mathcal{L}\left({A},B\right)\right]$ and $\mathcal{L}\left(A\otimes B,A\otimes B\right)$ are algebras with respect to composition and are isomorphic under the identification: $\Theta\leftrightarrow\mathsf{T}$ CDP08b ; Gou19a ; SSSS (23). The following lemma provides the necessary and sufficient conditions for the representing map $\mathsf{T}$ to be completely positive.
Lemma 3 (Gou19a ; CDP08b ).
A linear supermap $\Theta$ is completely CP-preserving if and only if its representing map $\mathsf{T}$ is CP.
We have already shown a supermap $\Theta\in\mathscr{L}\left[\mathcal{L}\left({A},B\right),\mathcal{L}\left({C},D\right)\right]$ is completely CP-preserving if and only if $\Lambda_{\Theta}=\sum_{ijkl}\mathcal{E}^{A\rightarrow B}_{ijkl}\otimes\Theta\left(\mathcal{E}^{A\rightarrow B}_{ijkl}\right)\in\mathcal{L}\left(A,B\right)\otimes\mathcal{L}\left(C,D\right)$ is CP, where $\Lambda_{\Theta}$ can be viewed as a map from $\mathcal{L}(A\otimes C)$ to $\mathcal{L}(B\otimes D)$, and $\{\mathcal{E}^{{A}\rightarrow B}_{\delta s\beta q}\}_{\delta s\beta q}$ is a basis for $\mathcal{L}(A,B)$. Hence, we can determine its complete positivity by checking the positivity of its Choi operator $\widehat{\mathsf{C}}_{\Lambda_{\Theta}}$. For detailed proof, see Appendix B.3.
Note that the representing map $\mathsf{T}:\mathcal{L}\left(A\otimes{B}\right)\rightarrow\mathcal{L}\left({C}\otimes{D}\right)$ preserves the trace for those $\mathsf{X}\in\mathcal{L}\left(A\otimes{B}\right)$ for which $\Gamma_{\mathsf{X}}$ is trace preserving but in general $\mathsf{T}$ could neither be a trace nonincreasing map nor a trace increasing map. However, we can transform the map $\mathsf{T}$ into a map $\mathsf{T}{{}^{\prime}}:\mathcal{L}\left(A\otimes{B}\right)\rightarrow\mathcal{L}\left(C\otimes{D}\right)$ in such a way that $\mathsf{T}{{}^{\prime}}$ is trace preserving and $\mathsf{T}(\mathsf{X})=\mathsf{T}{{}^{\prime}}(\mathsf{X})$ holds whenever $\Gamma_{\mathsf{X}}$ is trace preserving. The explicit form of the map $\mathsf{T}{{}^{\prime}}$ is given by the following expression:
$\mathsf{T}{{}^{\prime}}(\mathsf{X})=\mathsf{T}(\mathsf{X})+\left[\mathrm{tr}(\mathsf{X})-\mathrm{tr}(\mathsf{T}(\mathsf{X}))\right]\sigma_{0},$ (20)
where $\sigma_{0}\in\mathcal{L}(C\otimes D)$. In particular, $\mathsf{T}{{}^{\prime}}\left(\widehat{\mathsf{C}}_{\mathcal{N}}\right)=\mathsf{T}\left(\widehat{\mathsf{C}}_{\mathcal{N}}\right)$ holds for all quantum channels $\mathcal{N}$, if $\mathsf{T}$ does not change the space of operators, or $|A|=|C|$. Observe that $\mathsf{T}^{\prime}$ is CP if $\mathsf{T}$ is CP and trace nonincreasing map and we choose $\sigma_{0}$ to be a fixed state. For the case when $\mathsf{T}$ is neither trace nonincreasing map nor trace increasing map, we have derived necessary and sufficient conditions for $\mathsf{T}^{\prime}$ to be a CP map in terms of Choi operator of $\mathsf{T}$ in Appendix B.4. Let us consider a class $\mathsf{SCT}$ of superchannels such that for any superchannel $\Theta\in\mathsf{SCT}$, the representing map $\mathsf{T}$ of $\Theta$ satisfies our necessary and sufficient criteria such that $\mathsf{T}^{\prime}$ is CP (see Appendix B.4). In section IV , we will use this class of superchannels to establish several results for the channel recovery problem.
A given representing map $\mathsf{T}:\mathcal{L}\left(A\otimes{B}\right)\rightarrow\mathcal{L}\left(C\otimes{D}\right)$ induces a map $\Lambda:\mathcal{L}(A,B)\rightarrow\mathcal{L}(C,D)$ as follows,
$\Lambda(\mathcal{N})(\mathsf{X})=\mathrm{tr}_{C}\left[(\mathsf{X}^{t}\otimes\mathbbm{1})\mathsf{T}\left(\widehat{\mathsf{C}}_{\mathcal{N}}\right)\right],$ (21)
where $\mathsf{X}^{t}$ denote the transpose of operator $\mathsf{X}$, $\mathcal{N}\in\mathcal{L}(A,B)$, and $\widehat{\mathsf{C}}_{\mathcal{N}}$ is the Choi operator for $\mathcal{N}$. Before proceeding further we first find the adjoints of the maps $\mathsf{T}$ and $\mathsf{T}{{}^{\prime}}$.
Lemma 4 (Gou19a ).
The adjoint $\mathsf{T}^{*}:\mathcal{L}(C\otimes D)\rightarrow\mathcal{L}(A\otimes B)$ of the representing map $\mathsf{T}$ of a supermap $\Theta:\mathcal{L}(A,B)\rightarrow\mathcal{L}(C,D)$ is given by
$\mathsf{T}^{*}\left(\mathsf{Y}\right)=\widehat{\mathsf{C}}_{\Theta^{*}(\Gamma_{\mathsf{Y}})},$ (22)
where $\Theta^{*}$ is the adjoint of $\Theta$, $\mathsf{Y}\in\mathcal{L}\left({C\otimes D}\right)$ is the Choi operator of $\Gamma_{\mathsf{Y}}$, and $\widehat{\mathsf{C}}_{\Theta^{*}(\Gamma_{\mathsf{Y}})}$ is the Choi operator for ${\Theta^{*}(\Gamma_{\mathsf{Y}})}$.
This result, which tells that the identification $\mathsf{T}\leftrightarrow\Theta$ is a -isomorphism, was proved in Gou19a . We provide an alternate proof of this result in Appendix B.5.
Remark 1.
With the help of representing map, we notice that a supermap $\Theta$ is completely CP-preserving if and only if its adjoint $\Theta^{*}$ is completely CP-preserving.
Lemma 5.
The adjoint $\mathsf{T}^{{}^{\prime}*}:\mathcal{L}\left({C}\otimes{D}\right)\rightarrow\mathcal{L}\left({A}\otimes{B}\right)$ of the map $\mathsf{T}{{}^{\prime}}$, which is defined via Eq. (20), is given by
$\mathsf{T}^{{}^{\prime}*}(\mathsf{Y})=\mathsf{T}^{*}(\mathsf{Y})+\mathrm{tr}\left(\mathsf{Y}^{\dagger}\sigma_{0}\right)\left(\mathbbm{1}_{AB}-\mathsf{T}^{*}(\mathbbm{1}_{CD})\right),$ (23)
where $\sigma_{0}\in\mathcal{D}(C\otimes D)$ is a fixed state.
Proof.
Let us define a map $\mathsf{T}_{0}:\mathcal{L}\left({A}\otimes{B}\right)\rightarrow\mathcal{L}\left({C}\otimes{D}\right)$ such that its action on any $\mathsf{X}\in\mathcal{L}(A\otimes B)$ is given by $\mathsf{T}_{0}(\mathsf{X})=\left[\mathrm{tr}(\mathsf{X})-\mathrm{tr}(\mathsf{T}(\mathsf{X}))\right]\sigma_{0}$. Now, from the definition of the adjoint, for all $Y\in\mathcal{L}\left({C}\otimes{D}\right)$, we have
[…]
=$\mathrm{tr}\left(\mathsf{Y}^{\dagger}\sigma_{0}\right)\left[\mathbbm{1}_{AB}-\mathsf{T}^{*}(\mathbbm{1}_{CD})\right],\mathsf{X}\right>.$ (24)
So, we have $\mathsf{T}^{*}_{0}\left(\mathsf{Y}\right)=\mathrm{tr}\left(\mathsf{Y}^{\dagger}\sigma_{0}\right)[\mathbbm{1}_{AB}-\mathsf{T}^{*}(\mathbbm{1}_{CD})]$. As the adjoint is linear, we have the desired result. ∎
We now introduce the completely depolarizing map, which will play an important role in expressing our main results.
Definition 2.
The action of the completely depolarizing map $\mathcal{R}_{A\rightarrow B}$ is defined as
$\mathcal{R}_{A\rightarrow B}(\mathsf{X})=\mathrm{tr}(\mathsf{X})\mathbbm{1}_{B}\hskip 5.69046pt\forall~{}~{}\mathsf{X}\in\mathcal{L}(A).$ (25)
Lemma 6.
The representing map $\mathsf{T}:\mathcal{L}\left({A}\otimes{B}\right)\rightarrow\mathcal{L}\left({C}\otimes{D}\right)$ of a supermap $\Theta$ is trace preserving if and only if $\Theta^{*}:\mathcal{L}(C,D)\rightarrow\mathcal{L}(A,B)$ is completely depolarizing preserving ($\mathcal{R}$-preserving), i.e., $\Theta^{*}(\mathcal{R}_{C\rightarrow D})=\mathcal{R}_{A\rightarrow B}$.
Proof.
Let $\mathcal{R}_{A\rightarrow B}$ and $\mathcal{R}_{C\rightarrow D}$ be the completely depolarizing maps of $\mathcal{L}\left(A,B\right)$ and $\mathcal{L}\left({C,D}\right)$, respectively. Then we have the following relations.
[…]
This completes the proof of the lemma. ∎
Remark 2.
A similar proof technique can be applied to prove that the adjoint $\mathsf{T}^{*}$ of the representing map $T$ of a supermap $\Theta$ is trace preserving if and only if $\Theta$ is $\mathcal{R}$-preserving.
Representing map with respect to rank one operators with full rank reduced states.—Consider a rank one operator $\Psi_{RA}\in\mathsf{FRank}(R\otimes A)$ and an orthonormal basis {|$\{\ket{j}\}_{j}$ of $R\simeq A$. Such an operator can be written as $\Psi_{RA}=\outerproduct{\Psi}{\Psi}_{RA}$ with |$\ket{\Psi}_{RA}=(\mathsf{A}_{\Psi}\otimes\mathbbm{1}_{A})\sum_{i=1}^{|A|}\ket{i,i}_{RA}$, where $\mathsf{A}_{\Psi}\in\mathcal{L}(R)$ is defined via $\bra{i}\mathsf{A}_{\Psi}\ket{j}:=\langle i,j\ket{\Psi}_{RA}$. As $\mathrm{rank}(\mathsf{A}_{\Psi})=\mathrm{rank}(\mathsf{A}_{\Psi}\mathsf{A}_{\Psi}^{*})$, we have the reduced state $\Psi_{R}=\mathsf{A}_{\Psi}\mathsf{A}_{\Psi}^{*}$ of $\Psi_{RA}$ invertible if and only if $\mathsf{A}_{\Psi}$ is invertible. For such an operator, we define the Choi operator for a linear map $\mathcal{N}:\mathcal{L}(A)\rightarrow\mathcal{L}(B)$ as the following:
$\widehat{\mathsf{C}}_{\mathcal{N}}^{\Psi}:=\left(\mathrm{id}_{R\rightarrow A}\otimes\mathcal{N}_{A\rightarrow B}\right)\left(\Psi_{RA}\right).$ (26)
Now the Choi operator $\widehat{\mathsf{C}}_{\mathcal{N}}^{\Psi}\in\mathcal{L}(A\otimes B)$ and the Choi operator $\widehat{\mathsf{C}}_{\mathcal{N}}\in\mathcal{L}(A\otimes B)$ defined with respect to a maximally entangled operator $\Psi^{+}_{AB}$ are related as follows.
$\widehat{\mathsf{C}}_{\mathcal{N}}^{\Psi}$ $:=\left(\mathrm{id}_{R\rightarrow A}\otimes\mathcal{N}_{A\rightarrow B}\right)\left(\Psi_{RA}\right)$
$=(\mathsf{A}_{\Psi}\otimes\mathbbm{1}_{B})\widehat{\mathsf{C}}_{\mathcal{N}}(A^{*}_{\Psi}\otimes\mathbbm{1}_{B})$
$=T_{\Psi}\left(\widehat{\mathsf{C}}_{\mathcal{N}}\right),$ (27)
where $T_{\Psi}:\mathcal{L}\left({A}\otimes{B}\right)\rightarrow\mathcal{L}\left({A}\otimes{B}\right)$ is given by $T_{\Psi}(\mathsf{X}_{AB})=(\mathsf{A}_{\Psi}\otimes\mathbbm{1}_{B})\mathsf{X}_{AB}(\mathsf{A}^{*}_{\Psi}\otimes\mathbbm{1}_{B}),$ for all $\mathsf{X}_{AB}\in\mathcal{L}({A}\otimes{B})$. As $\mathsf{A}_{\Psi}$ is invertible, we have $T_{\Psi}$ invertible, and looking at the above form it is clear that $T_{\Psi}$ is completely positive, hence the Choi-Jamiołkowski isomorphism holds if we use $\Psi_{RA}\in\mathsf{FRank}(R\otimes A)$ instead of maximally entangled operator $\Psi^{+}_{RA}\in\mathsf{FRank}(R\otimes A)$.
Remark 3.
Both $\Psi^{+}_{RA}$ and $\Psi_{RA}$ belong in $\mathsf{FRank}(R\otimes A)$ and we have $\operatorname{tr}\left(\Psi^{+}_{RA}\right)=|A|=\operatorname{tr}\left(\Psi_{RA}\right)$. Thus, Eq. (II) is valid for $\mathsf{C}^{\Psi}_{\mathcal{N}}=\widehat{\mathsf{C}}^{\Psi}_{\mathcal{N}}/|A|$ and $\mathsf{C}_{\mathcal{N}}=\widehat{\mathsf{C}}_{\mathcal{N}}/|A|$, i.e., $\mathsf{C}_{\mathcal{N}}^{\Psi}=T_{\Psi}\left(\mathsf{C}_{\mathcal{N}}\right)$. If $\mathcal{N}$ is a quantum channel then both $\mathsf{C}_{\mathcal{N}}^{\Psi}$ and $\mathsf{C}_{\mathcal{N}}$ become the Choi states.
Note that $T_{\Psi}$ induces a supermap $\mathrm{\Theta}_{\Psi}:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(A,B\right)$ given by
$\mathrm{\Theta}_{\Psi}(\mathcal{N})(\mathsf{X})=\operatorname{tr}_{A}\left((\mathsf{X}^{t}\otimes\mathbbm{1}_{B})T_{\Psi}\left(\widehat{\mathsf{C}}_{\mathcal{N}}\right)\right),$ (28)
for all $\mathcal{N}\in\mathcal{L}\left({A},{B}\right)$ and $\mathsf{X}\in\mathcal{L}({A})$, where $\mathsf{X}^{t}$ represents the transpose of $\mathsf{X}$ with respect to the usual canonical basis. Similarly, we can define $T_{\Phi}$, $\mathrm{\Theta}_{\Phi}$, and $\widehat{\mathsf{C}}_{\mathcal{M}}^{\Phi}$ for another rank one operator $\Phi_{RC}\in\mathsf{FRank}(R\otimes C)$, and a map $\mathcal{M}:\mathcal{L}(C)\rightarrow\mathcal{L}(D)$.
We now define a representing map $\mathfrak{T}:\mathcal{L}({A}\otimes{B})\rightarrow\mathcal{L}({C}\otimes{D})$ for a supermap $\mathrm{\Theta}:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ with respect to the Choi operators $\widehat{\mathsf{C}}_{\mathcal{N}}^{\Psi}$ and $\widehat{\mathsf{C}}_{\mathcal{M}}^{\Phi}$ by the following:
$\mathfrak{T}\left(\widehat{\mathsf{C}}_{\mathcal{N}}^{\Psi}\right)=\widehat{\mathsf{C}}_{\mathrm{\Theta}(\mathcal{N})}^{\Phi}.$ (29)
Now we establish the relation between the representing map $\mathsf{T}$ of the supermap $\mathrm{\Theta}$ defined with respect to the Choi operators $\widehat{\mathsf{C}}_{\mathcal{N}}$ and $\widehat{\mathsf{C}}_{\mathcal{M}}$, and the representing map $\mathfrak{T}$ of the supermap $\mathrm{\Theta}$ defined with respect to the Choi operators $\widehat{\mathsf{C}}_{\mathcal{N}}^{\Psi}$ and $\widehat{\mathsf{C}}_{\mathcal{M}}^{\Phi}$ as follows. We have $\mathsf{T}\left(\widehat{\mathsf{C}}_{\mathcal{N}}\right)=\widehat{\mathsf{C}}_{\mathrm{\Theta}(\mathcal{N})}\Leftrightarrow\mathsf{T}\circ T_{\Psi}^{-1}\left(\widehat{\mathsf{C}}_{\mathcal{N}}^{\Psi}\right)=T_{\Phi}^{-1}\left(\widehat{\mathsf{C}}_{\mathrm{\Theta}(\mathcal{N})}^{\Phi}\right)$, which implies $T_{\Phi}\circ\mathsf{T}\circ T_{\Psi}^{-1}\left(\widehat{\mathsf{C}}_{\mathcal{N}}^{\Psi}\right)=\widehat{\mathsf{C}}_{\mathrm{\Theta}(\mathcal{N})}^{\Phi}$. By comparing this with Eq. (29), we have
$\mathfrak{T}=T_{\Phi}\circ\mathsf{T}\circ T_{\Psi}^{-1}.$ (30)
Noticing that $T^{-1}_{\Psi}$ and $T_{\Phi}$ are both completely positive, we have that $\mathfrak{T}$ is completely positive if and only if $\mathsf{T}$ is completely positive. From Lemma 3, $\mathsf{T}$ is completely positive if and only if $\mathrm{\Theta}$ is completely CP-preserving. Thus, we conclude that $\mathfrak{T}$ is completely positive if and only if $\mathrm{\Theta}$ is completely CP-preserving.
The adjoint map $\mathfrak{T}^{*}$.— We already know that $\mathsf{T}^{*}\left(\widehat{\mathsf{C}}_{\mathcal{M}}\right)=\widehat{\mathsf{C}}_{\mathrm{\Theta}^{*}(\mathcal{M})}$. As the map $\mathsf{T}\leftrightarrow\mathrm{\Theta}$ is product preserving, we have $T^{-1}\left(\widehat{\mathsf{C}}_{\mathcal{M}}\right)=\widehat{\mathsf{C}}_{\mathrm{\Theta}^{-1}(\mathcal{M})}$. From this relation we derive the action of $\mathfrak{T}^{*}:\mathcal{L}({C}\otimes{D})\rightarrow\mathcal{L}({A}\otimes{B})$ on $\widehat{\mathsf{C}}_{\mathcal{M}}^{\Phi}$ as
$\mathfrak{T}^{*}\left(\widehat{\mathsf{C}}_{\mathcal{M}}^{\Phi}\right)=(T_{\Psi}^{-1})^{*}\circ\mathsf{T}^{*}\circ(T_{\Phi})^{*}\circ T_{\Phi}\left(\widehat{\mathsf{C}}_{\mathcal{M}}\right)=\widehat{\mathsf{C}}^{\Psi}_{\mathrm{\Theta}_{\Psi}^{-1}\circ\mathrm{\Theta}^{-1*}_{\Psi}\circ\mathrm{\Theta}^{*}\circ\mathrm{\Theta}^{*}_{\Phi}\circ\mathrm{\Theta}_{{\Phi}(\mathcal{M})}}.$ (31)
By defining $\mathrm{\Theta}^{*}_{\Phi}\circ\mathrm{\Theta}_{\Phi}:=\hat{\mathrm{\Theta}}_{\Phi}$ and $\mathrm{\Theta}^{*}_{\Psi}\circ\mathrm{\Theta}_{\Psi}:=\hat{\mathrm{\Theta}}_{\Psi}$ we can rewrite above equation as
$\mathfrak{T}^{*}\left(\widehat{\mathsf{C}}_{\mathcal{M}}^{\Phi}\right)$ $=\widehat{\mathsf{C}}^{\Psi}_{{\hat{\mathrm{\Theta}}_{\Psi}}^{-1}\circ\mathrm{\Theta}^{*}\circ\hat{\mathrm{\Theta}}_{{\Phi}(\mathcal{M})}}.$ (32)
Representing map for super-supermaps.— Let us denote by $\mathring{\mathbf{\Lambda}}:\mathscr{L}\left[\mathcal{L}\left(A,B\right),\mathcal{L}\left(C,D\right)\right]\rightarrow\mathscr{L}\left[\mathcal{L}\left(A^{\prime},B^{\prime}\right),\mathcal{L}\left(C^{\prime},D^{\prime}\right)\right]$ a higher order process transforming supermaps to supermaps. We call it a super-supermap. Now, we can define its representing supermap $\mathbb{T}_{\mathring{\mathbf{\Lambda}}}:\mathcal{L}\left(A,B\right)\otimes\mathcal{L}\left(C,D\right)\rightarrow\mathcal{L}\left(A^{\prime},B^{\prime}\right)\otimes\mathcal{L}\left(C^{\prime},D^{\prime}\right)$ by
$\mathbb{T}_{\mathring{\mathbf{\Lambda}}}(\Lambda_{\Theta})=\Lambda_{\mathring{\mathbf{\Lambda}}(\Theta)}.$ (33)
where the supermap $\Theta\in\mathscr{L}\left[\mathcal{L}\left(A,B\right),\mathcal{L}\left(C,D\right)\right]$ and $\Lambda_{\Theta}=\sum_{ijkl}\mathcal{E}^{A\rightarrow B}_{ijkl}\otimes\Theta\left(\mathcal{E}^{A\rightarrow B}_{ijkl}\right)$. For a given supermap $\mathbb{T}_{\mathring{\mathbf{\Lambda}}}$, one can find its corresponding super-supermap $\mathring{\mathbf{\Lambda}}$ as
$\mathring{\mathbf{\Lambda}}[\Theta](\mathcal{N})=\operatorname{tr}_{A^{\prime}}[(\mkern 2.0mu\overline{\mkern-2.0mu\mathcal{N}\mkern-2.0mu}\mkern 2.0mu^{*}\otimes\operatorname{id})\mathbb{T}_{\mathring{\mathbf{\Lambda}}}(\Lambda_{\Theta})],$ (34)
where $\mkern 2.0mu\overline{\mkern-2.0mu\mathcal{N}\mkern-2.0mu}\mkern 2.0mu$ represents the complex conjugate of the map $\mathcal{N}$ defined as $\mkern 2.0mu\overline{\mkern-2.0mu\mathcal{N}\mkern-2.0mu}\mkern 2.0mu:=\sum_{ijkl}\bar{\alpha}_{ijkl}\mathcal{E}^{A^{\prime}\rightarrow B^{\prime}}_{ijkl}$ when $\mathcal{N}$ is written as $\mathcal{N}:=\sum_{ijkl}\alpha_{ijkl}\mathcal{E}^{A^{\prime}\rightarrow B^{\prime}}_{ijkl}$ and represents the adjoint as defined in Eq. (7). The trace of a map $\mathcal{N}:\mathcal{L}(A^{\prime})\rightarrow\mathcal{L}(A^{\prime})$ is defined by $\operatorname{tr}[\mathcal{N}]=\sum_{ij}\left<e^{A^{\prime}}_{ij},\mathcal{N}\left(e^{A^{\prime}}_{ij}\right)\right>$ with $e^{A^{\prime}}_{ij}$ being the canonical basis of the space $A^{\prime}$. Using the form of Eq. (13) for $\mathbb{T}_{\mathring{\mathbf{\Lambda}}}$ in Eq. (34), we expect a representation similar to Eq. (13) to hold for a super-superchannel $\mathring{\mathbf{\Lambda}}$.
III Entropy change of quantum channels
Distinguishing two different quantum states is one of the most basic tasks in the quantum information theory as multitude of problems in the quantum information theory can be recast as state distinguishability problems Hel (76). In the asymptotic limit of quantum state discrimination the minimum error in distinguishing two quantum states is expressed neatly by quantum Chernoff bound ACMnT+ (07), which is very closely related to the relative entropy between the states Ved (02). This provides an operational meaning to the relative entropy between two quantum states. On the other hand, relative entropy plays a pivotal role in proving a plethora of results such as the strong subaddivity of the von Neumann entropy NC (00) and recoverability of quantum states affected by quantum noise OP (93). In fact, the quantum relative entropy can be treated as a parent quantity for analysis as many other information theoretic quantities such as the von Neumann entropy, the quantum mutual information NC (00), and the free energy BaHO+ (13) can be re-expressed in terms of relative entropy. Our interest here is in the von Neumann entropy. The von Neumann entropy is a measure of randomness or uncertainty present in a quantum system and has a variety of uses in quantum information theory, e.g., it is crucial in the theory of entanglement Sch (96), quantum communications Hol (98); YO (93), and quantum cryptography CBTW (17). It is a natural generalization of Shannon entropy Sha (48). We can write von Neumann entropy in terms of the quantum relative entropy as follows
[…]
If the quantum system is finite-dimensional, the above formula can also be written as
$S(A)_{\rho}=\log|A|-D\left(\rho_{A}\;\middle\|\;\mathbbm{\widetilde{1}}_{A}\right),$ (36)
where $\mathbbm{\widetilde{1}}_{A}:=\mathbbm{1}_{A}/|A|$ is the maximally mixed state of the system $A$.
While quantum states describe a quantum system at any instant completely, its dynamics is described by the quantum channels. It is insightful to note that quantum states themselves can be thought of as quantum channels with one dimensional inputs and quantum measurements can be thought of as quantum channels with one dimensional outputs CDP08b . When viewed from an operational perspective, a quantum state is a preparation process while a measurement describes post-processing. In this sense, quantum channels can be thought of as a universal paradigm describing both static and dynamic processes of quantum systems.
To this end, one can naturally ask: Is there a similar notion of entropy (or relative entropy) that can be applied to quantum channels? One way to answer this question is to first find a meaningful notion of relative entropy between quantum channels, then, similar to quantum states, relative entropy between quantum channels can be used to define the entropy of quantum channels. This procedure was followed in GW (21) to define the notion of entropy for a quantum channel. The notion of relative entropy between two quantum channels also has an operational meaning analogous to quantum states, namely, it quantifies the difficulty in discriminating two quantum channels GW (21). It was advocated in GW (21) that the entropy functional for a quantum channel should satisfy some physically motivated axioms (see Appendix C.4). In GW (21), Eq. (36) is considered as the starting point for defining entropy of quantum channels. For a quantum channel $\mathcal{N}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left(B\right)$, they have defined its entropy $S(\mathcal{N})$ as
$S(\mathcal{N}):=\log|B|-D\left[\mathcal{N}\;\middle\|\;\widetilde{\mathcal{R}}\right],$ (37)
where ~$\widetilde{\mathcal{R}}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left(B\right)$ is the completely depolarizing channel defined by ~$\widetilde{\mathcal{R}}(\mathsf{X})=\mathrm{tr}(\mathsf{X})\widetilde{\mathbbm{1}}_{B}$ and $D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]$ is the relative entropy between two quantum channels $\mathcal{N}$ and $\mathcal{M}$ defined by GW (21)
$D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]:=\underset{\rho_{RA}\in\mathcal{D}(R\otimes A)}{\mathrm{sup}}~{}D\left(\mathrm{id}_{R}\otimes\mathcal{N}(\rho_{RA})\;\middle\|\;\mathrm{id}_{R}\otimes\mathcal{M}(\rho_{RA})\right).$ (38)
It was taken as an axiom that entropy should increase under the action of random unitary superchannels. A stronger statement was also proved: if a superchannel $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ maps the completely depolarizing channel to the completely depolarizing channel, then the entropy of a quantum channel never decreases under the action of such quantum superchannels, provided $|B|=|D|$ and $|A|=|C|$. The entropy, Eq. (37), never decreases under the action of a random unitary superchannel as a consequence of this stronger statement.
Recall that under a positive trace preserving sub-unital map $\mathcal{N}$, the entropy of a quantum state never decreases i.e., $S(\mathcal{N}(\rho))\geq S(\rho)$. We want to investigate if a similar statement holds for the entropy of quantum channels as well. We note that the generalization of the identity matrix is the completely depolarizing map (see Definition 2) and hence the generalization of sub-unitality for a supermap $\mathrm{\Theta}$ is captured by $\mathrm{\Theta}(\mathcal{R})\leq\mathcal{R}$, where the ordering between maps is defined in Proposition 7. Here, we ask whether the entropy of a quantum channel is nondecreasing under those quantum superchannels $\mathrm{\Theta}$ for which $\mathrm{\Theta}(\mathcal{R}_{A\rightarrow B})\leq\mathcal{R}_{C\rightarrow D}$ holds. To answer this question, first we rewrite the definition of entropy of a quantum channel which is in a similar spirit to Eq. (35) (see Definition 5 ), but to do that we first need the notion of relative entropy between a quantum channel and a completely positive (CP) map. This is because, in the settings of quantum channels, the generalization of identity is the completely depolarising map. In Proposition 8, we have shown that if $\Theta(\mathcal{R}_{A\rightarrow B})\leq\mathcal{R}_{C\rightarrow D}$, then the entropy of channels is non-decreasing under the action of this superchannel, i.e., $S\left[\Theta(\mathcal{N}_{A\rightarrow B})\right]\geq S\left[\mathcal{N}_{A\rightarrow B}\right]$. We also observe that a necessary condition for the relation $S\left[\Theta(\mathcal{N}_{A\rightarrow B})\right]\geq S\left[\mathcal{N}_{A\rightarrow B}\right]$ to hold is $|B|\leq|D|$.
III.1 Data processing inequality
The generalized channel divergence between two given quantum channels $\mathcal{N}_{A\rightarrow B}$ and $\mathcal{M}_{A\rightarrow B}$ was defined in CMW (16); LKDW (18); Yua (19). In the following, we analogously define the relative entropy between a quantum channel and a CP map.
Definition 3 (CMW (16); LKDW (18)).
The relative entropy between a quantum channel $\mathcal{N}_{A\to B}$ and a completely positive map $\mathcal{M}_{A\to B}$ is defined as
$D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]=\underset{\rho_{RA}\in\mathcal{D}(R\otimes A)}{\mathrm{sup}}~{}D(\mathrm{id}_{R}\otimes\mathcal{N}(\rho_{RA})\;\|\;\mathrm{id}_{R}\otimes\mathcal{M}(\rho_{RA})),$ (39)
where $R$ is a reference system of arbitrary dimension. Due to purification, data processing inequality, and Schmidt decomposition, it suffices to optimize over the set of all pure bipartite states $\Psi_{RA}$, i.e.,
$D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]=\underset{\Psi_{RA}\in\mathcal{D}(R\otimes A)}{\mathrm{sup}}~{}D(\mathrm{id}_{R}\otimes\mathcal{N}(\Psi_{RA})\;\|\;\mathrm{id}_{R}\otimes\mathcal{M}(\Psi_{RA})).$ (40)
We now investigate properties of the relative entropy between a quantum channel $\mathcal{N}_{A\rightarrow A^{\prime}}$ and a completely positive map $\mathcal{M}_{A\rightarrow A^{\prime}}$. (i) The relative entropy $D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]$ is non-negative if the state that realizes the supremum, say $\widetilde{\Psi}_{RA}$, is such that $\mathrm{tr}\left(\mathrm{id}_{R}\otimes\mathcal{M}\left(\widetilde{\Psi}_{RA}\right)\right)\leq 1$. (ii) The relative entropy $D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]$ is monotonic under the action of a quantum superchannel, i.e., follows data processing type inequality, which we prove as the following proposition.
Proposition 1.
The relative entropy between a quantum channel $\mathcal{N}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left({A^{\prime}}\right)$ and a completely positive map $\mathcal{M}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left({A^{\prime}}\right)$ is nonincreasing under the action of a quantum superchannel $\Theta:\mathcal{L}\left({A},{A^{\prime}}\right)\rightarrow\mathcal{L}\left({C},{C^{\prime}}\right)$, i.e.,
$D\left[\Theta\left(\mathcal{N}\right)\;\middle\|\;\Theta\left(\mathcal{M}\right)\right]\leq D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right].$ (41)
We have provided detailed proof of the above proposition in Appendix C.5. Recall that the relative entropy between a state and a positive operator is monotonic under the action of a quantum channel. The above proposition tells us that the same properties also hold if we replace the state with a channel, the channel with a superchannel, and the positive operator with a CP linear map. Note that the relative entropy between two quantum states (or two positive semidefinite operators) is monotonic under a larger class of operations, namely, under all positive and trace preserving maps (compared to a restricted class of CPTP maps) MHR (17). What happens to the monotonicty of relative entropy between two quantum channels if we restrict to process the quantum channels under the larger class of CP-preserving and trace preserving linear maps (cf. Proposition 1)? We answer this question in the affirmative and thereby generalize the monotonicity result of MHR (17) to the higher order transformations. More specifically, if we restrict ourselves to a particular class of channels such that for any pair of channels $\mathcal{N}$ and $\mathcal{M}$, the supremum in the relative entropy functional occurs at some state $\Psi$ such that its reduced state is full rank then the monotonicity of relative entropy also holds for CP-preserving and TP-preserving supermaps, which we prove as our next proposition.
Proposition 2.
The relative entropy between quantum channels $\mathcal{N}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left({A^{\prime}}\right)$ and $\mathcal{M}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left({A^{\prime}}\right)$ is nonincreasing under the action of a CP-preserving and TP-preserving supermap $\Theta:\mathcal{L}\left({A},{A^{\prime}}\right)\rightarrow\mathcal{L}\left({C},{C^{\prime}}\right)$, i.e.,
$D\left[\Theta\left(\mathcal{N}\right)\;\middle\|\;\Theta\left(\mathcal{M}\right)\right]\leq D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right],$ (42)
if the pure states that realize the supremum in the relative entropy functional $D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]$ and $D\left[\Theta\left(\mathcal{N}\right)\;\middle\|\;\Theta\left(\mathcal{M}\right)\right]$ have the full rank marginals, and the representing map of $\Theta$ is a trace non increasing map.
The detailed proof of the above proposition, which is one of our main results, is provided in Appendix C.2. We further provide an explicit condition on supermap $\Theta$ in Eq. (118) such that the monotonicity holds for the above-mentioned class of channels even if we take supermap $\Theta$ to be a CP-preserving map but not TP-preserving. It will be interesting to see if the result holds for all quantum channels. We also extend some properties that hold for the relative entropy between a state and a positive operator to the relative entropy between a channel and a CP map, for example, if we have $\mathsf{P}\leq\mathsf{Q}$, where $\mathsf{P}$ and $\mathsf{Q}$ are positive operators, then for any state $\rho$, we also have $D\left(\rho\;\middle\|\;\mathsf{P}\right)\geq D\left(\rho\;\middle\|\;\mathsf{Q}\right)$. We prove the channel version of this inequality as the following proposition.
Proposition 3.
Consider a quantum channel $\mathcal{N}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left({A^{\prime}}\right)$ and two completely positive maps $\mathcal{M}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left({A^{\prime}}\right)$ and ~$\widetilde{\mathcal{M}}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left({A^{\prime}}\right)$ with ordering $\widetilde{\mathcal{M}}\geq\mathcal{M}$, then we have
$D\left[\mathcal{N}\;\middle\|\;\widetilde{\mathcal{M}}\right]\leq D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right].$ (43)
Proof.
Since $\widetilde{\mathcal{M}}$ and $\mathcal{M}$ are completely positive maps with ordering $\widetilde{\mathcal{M}}\geq\mathcal{M}$, then the following ordering will also hold: $\Delta_{\mathcal{M},\widetilde{\mathcal{M}}}(\epsilon):=\log\left\{\left(\mathrm{id}_{R}\otimes\widetilde{\mathcal{M}}\right)\left(\Psi_{RA}\right)+\epsilon\mathbbm{1}_{RA^{\prime}}\right\}-\log\left\{\left(\mathrm{id}_{R}\otimes\mathcal{M}\right)\left(\Psi_{RA}\right)+\epsilon\mathbbm{1}_{RA^{\prime}}\right\}\geq 0$ for any $\epsilon>0$. Now, we have
00 0 $\leq\lim_{\epsilon\rightarrow 0+}\operatorname{tr}\left(\left(\mathrm{id}_{R}\otimes\mathcal{N}\right)\left(\Psi_{RA}\right)~{}\Delta_{\mathcal{M},\widetilde{\mathcal{M}}}(\epsilon)\right)$
$=\lim_{\epsilon\rightarrow 0+}D\left(\left(\mathrm{id}_{R}\otimes\mathcal{N}\right)\left(\Psi_{RA}\right)\;\middle\|\;\left(\mathrm{id}_{R}\otimes\mathcal{M}\right)\left(\Psi_{RA}\right)+\epsilon\mathbbm{1}_{RA^{\prime}}\right)$
$-D\left(\left(\mathrm{id}_{R}\otimes\mathcal{N}\right)\left(\Psi_{RA}\right)\;\middle\|\;\left(\mathrm{id}_{R}\otimes\widetilde{\mathcal{M}}\right)\left(\Psi_{RA}\right)+\epsilon\mathbbm{1}_{RA^{\prime}}\right)$
$=D\left(\left(\mathrm{id}_{R}\otimes\mathcal{N}\right)\left(\Psi_{RA}\right)\;\middle\|\;\left(\mathrm{id}_{R}\otimes\mathcal{M}\right)\left(\Psi_{RA}\right)\right)-D\left(\left(\mathrm{id}_{R}\otimes\mathcal{N}\right)\left(\Psi_{RA}\right)\;\middle\|\;\left(\mathrm{id}_{R}\otimes\widetilde{\mathcal{M}}\right)\left(\Psi_{RA}\right)\right).$
After taking the supremum over the set of pure states $\Psi_{RA}$, we obtain our desired inequality, i.e., $D\left[\mathcal{N}\;\middle\|\;\widetilde{\mathcal{M}}\right]\leq D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]$. ∎
Proposition 4.
Consider a pair of quantum channels $\mathcal{N}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left({A^{\prime}}\right)$ and $\widetilde{\mathcal{N}}:\mathcal{L}\left({B}\right)\rightarrow\mathcal{L}\left({B^{\prime}}\right)$, and another pair of completely positive maps $\mathcal{M}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left({A^{\prime}}\right)$ and ~$\widetilde{\mathcal{M}}:\mathcal{L}\left({B}\right)\rightarrow\mathcal{L}\left({B^{\prime}}\right)$, then we have
$D\left[\mathcal{N}\otimes\widetilde{\mathcal{N}}\;\middle\|\;\mathcal{M}\otimes\widetilde{\mathcal{M}}\right]\geq D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]+D\left[\widetilde{\mathcal{N}}\;\middle\|\;\widetilde{\mathcal{M}}\right].$ (44)
The proof of this Proposition is provided in Appendix C.3.
III.2 Entropy of quantum channels and its properties
In this section, we study the entropy of a quantum channel and its behavior under quantum superchannels. We start by first introducing a class of supermaps called $\mathcal{R}$-subpreserving supermaps as we will later show that the entropy of a quantum channel is nondecreasing under the action of $\mathcal{R}$-subpreserving superchannels.
Definition 4.
Let $\mathcal{R}_{A\rightarrow B}$ and $\mathcal{R}_{C\rightarrow D}$ be two completely depolarizing maps, i.e., $\mathcal{R}_{A\rightarrow B}(\mathsf{X})=\operatorname{tr}(\mathsf{X})\mathbbm{1}_{B}$ and $\mathcal{R}_{A\rightarrow B}(\mathsf{Y})=\operatorname{tr}(\mathsf{Y})\mathbbm{1}_{D}$ for all $\mathsf{X}\in\mathcal{L}(A)$ and $\mathsf{Y}\in\mathcal{L}(C)$. Then, a supermap $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ is called $\mathcal{R}$-subpreserving if
$\Theta(\mathcal{R}_{A\rightarrow B})\leq\mathcal{R}_{C\rightarrow D},$ (45)
where the above ordering means that the map $\mathcal{R}_{C\rightarrow D}-\Theta(\mathcal{R}_{A\rightarrow B})$ is completely positive.
The entropy of a quantum channel is formally defined in terms of relative entropy between the quantum channel and the completely depolarizing map as follows.
Definition 5.
Consider a quantum channel $\mathcal{N}_{A\rightarrow B}$, and the completely depolarizing map $\mathcal{R}_{A\rightarrow B}$. The entropy of $\mathcal{N}_{A\rightarrow B}$ is defined as (cf. Eq. (37))
$S\left[\mathcal{N}_{A\rightarrow B}\right]:=-D\left[\mathcal{N}_{A\rightarrow B}\;\middle\|\;\mathcal{R}_{A\rightarrow B}\right],$ (46)
where $D\left[\mathcal{N}_{A\rightarrow B}\;\middle\|\;\mathcal{R}_{A\rightarrow B}\right]$ is the relative entropy between a quantum channel $\mathcal{N}_{A\rightarrow B}$ and the completely depolarizing map $\mathcal{R}_{A\rightarrow B}$.
Note that since the relative entropy between two quantum channels quantifies the minimum error in discriminating the two channels  GW (21), the definition of entropy (Definition 5) is naturally endowed with the same operational meaning. That is, the entropy of a quantum channel quantifies how well it can be distinguished from the completely randomizing channel. We now show that the definition of the entropy functional has some nice properties: (1) it is nondecreasing under random unitary superchannels, (2) it is additive, and (3) it satisfies normalization condition. First, we prove that the entropy functional of a quantum channel is additive.
Proposition 5.
The entropy functional of quantum channels is additive under tensor product of channels i.e,
$S\left[{\mathcal{N}}_{A_{1}\rightarrow B_{1}}\otimes{\mathcal{M}}_{A_{2}\rightarrow B_{2}}\right]=S\left[{\mathcal{N}}_{A_{1}\rightarrow B_{1}}\right]+S\left[{\mathcal{M}}_{A_{2}\rightarrow B_{2}}\right].$ (47)
To show that the entropy of quantum channels is additive it is sufficient to prove that the following equality also holds
$D\left[\mathcal{N}_{A_{1}\rightarrow B_{1}}\otimes{\mathcal{M}}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{1}\rightarrow B_{1}}\otimes{\mathcal{R}}_{A_{2}\rightarrow B_{2}}\right]=D\left[\mathcal{N}_{A_{1}\rightarrow B_{1}}\;\middle\|\;\mathcal{R}_{A_{1}\rightarrow B_{1}}\right]+D\left[\mathcal{M}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{2}\rightarrow B_{2}}\right].$ (48)
The complete proof is presented in Appendix C.5. Now we prove that the entropy functional is maximum for the completely depolarizing channel and zero for a replacement channel which outputs a pure state.
Proposition 6.
Let $\mathcal{N}_{A\rightarrow B}$ be a replacer channel, i.e., for all $\rho\in\mathcal{D}(A)$, it satisfies $\mathcal{N}_{A\rightarrow B}(\rho)=\sigma_{0}$ for a fixed $\sigma_{0}\in\mathcal{D}(B)$. Then $S(\mathcal{N}_{A\rightarrow B})=0$ when $\sigma_{0}$ is a pure state and $S(\mathcal{N}_{A\rightarrow B})$ is maximum i.e., $S(\mathcal{N}_{A\rightarrow B})=\log|B|$ when $\sigma_{0}=\frac{1}{|B|}\mathbbm{1}_{B}$, i.e., when $\mathcal{N}_{A\rightarrow B}$ is the completely depolarizing channel.
Proof.
Let $\Psi_{RA}$ be an arbitrary pure state. It can be shown that $\mathrm{id}_{R}\otimes\mathcal{N}_{A\rightarrow B}(\Psi_{RA})=\Psi_{R}\otimes\sigma_{0}$ and $\mathrm{id}_{R}\otimes\mathcal{R}_{A\rightarrow B}(\Psi_{RA})=\Psi_{R}\otimes\mathbbm{1}_{B}$. Now, we have
[…]
Another crucial property that the entropy functional of a quantum channel has to satisfy is that it should be nondecreasing under the action of a random unitary superchannel. The definition used in Gou19b satisfies this property as a corollary of the fact that the channel entropy always increases under those quantum superchannel which preserves the completely depolarizing channel. Here we generalize this result and show that the entropy functional is nondecreasing under completely depolarizing map subpreserving superchannels. Random isometry superchannels are one of the examples of such class of superchannels and we prove this as the following proposition.
Definition 6.
A random isometry superchannel $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ is defined as:
$\mathrm{\Theta}(\mathcal{N}_{A\rightarrow B})=\sum_{i}p_{i}\mathcal{V}_{B\rightarrow D}^{i}\circ\mathcal{N}_{A\rightarrow B}\circ\mathcal{U}_{C\rightarrow A}^{i},$ (49)
where $\{p_{i}\}_{i}$ is a probability distribution, and $\mathcal{V}_{B\rightarrow D}^{i}$, and $\mathcal{U}_{C\rightarrow A}^{i}$ are isometry channels defined as $\mathcal{V}_{B\rightarrow D}^{i}(X)=v^{i}Xv^{i\dagger}$ and $\mathcal{U}_{C\rightarrow A}^{i}(Y)=u^{i}Yu^{i\dagger}$ respectively. Here, $u^{i}:C\rightarrow A$ and $v^{i}:B\rightarrow D$ are isometries.
Proposition 7.
A random isometry superchannel $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ is an $\mathcal{R}$-subpreserving superchannel.
Proof.
Consider a random isometry superchannel $\Theta$ given by Eq. (49). Consider a positive semidefinite operator $\mathsf{X}\in\mathcal{L}\left({C}\right)$, we then have $\left[\Theta(\mathcal{R}_{A\rightarrow B})\right]\left(X\right)=\sum_{i}p_{i}\mathcal{V}_{B\rightarrow D}^{i}\circ\mathcal{R}_{A\rightarrow B}\circ\mathcal{U}_{C\rightarrow A}^{i}\left(\mathsf{X}\right)$. Using the definitions of $\mathcal{U}^{i}_{C\rightarrow A}$, and $\mathcal{R}_{A\rightarrow B}$, and $\mathcal{V}^{i}_{B\rightarrow D}$, we get
$\left[\Theta(\mathcal{R}_{A\rightarrow B})\right]\left(X\right)=\sum_{i}p_{i}v^{i}v^{i\dagger}\mathbbm{1}_{D}\mathrm{tr}\left(\mathsf{X}\right)=\sum_{i}p_{i}v^{i}v^{i\dagger}\mathcal{R}_{C\rightarrow D}\left(\mathsf{X}\right).$ (50)
Let us define $V:D\rightarrow D$ by $V=\sum_{i}p_{i}v^{i}v^{i\dagger}$. So we have the following relation
$\left[\Theta(\mathcal{R}_{A\rightarrow B})\right](\mathsf{X})=V\mathcal{R}_{C\rightarrow D}(\mathsf{X})=\mathcal{R}_{C\rightarrow D}\left(\mathsf{X}\right)V.$ (51)
Notice that $v^{i}v^{i\dagger}$ are projectors from $D$ onto the range of $v^{i}$ which implies that $\mathbbm{1}_{D}-V\geq 0$. Whenever $\mathsf{X}$ is positive semidefinite, $\mathcal{R}_{C\rightarrow D}(\mathsf{X})$ is positive semidefinite. As $\mathbbm{1}_{D}-V$ and $\mathcal{R}_{C\rightarrow D}(\mathsf{X})$ commutes, their product is also positive semidefinite i.e. $(\mathbbm{1}_{D}-V)\mathcal{R}_{C\rightarrow D}(X)\geq 0$ which implies that the map $(\mathbbm{1}_{D}-V)\mathcal{R}_{C\rightarrow D}$ is positive. It can be shown that $(\mathbbm{1}_{D}-V)\mathcal{R}_{C\rightarrow D}$ is completely positive as well. Hence, we have the following ordering, $\mathbbm{1}_{D}\mathcal{R}_{C\rightarrow D}\geq V\mathcal{R}_{C\rightarrow D}$, which implies
$\Theta(\mathcal{R}_{A\rightarrow B})\leq\mathcal{R}_{C\rightarrow D}.$ (52)
Remark 4.
Recall that for an $\mathcal{R}$-subpreserving supermap $\Theta$, we have $\Theta(\mathcal{R}_{A\rightarrow B})\leq\mathcal{R}_{C\rightarrow D}$. If we take $\Theta$ to be a random unitary superchannel, then we get $\Theta(\mathcal{R}_{A\rightarrow B})=\mathcal{R}_{C\rightarrow D}$ i.e., $\Theta$ becomes $\mathcal{R}$-preserving superchannel. Using Eq. (13), we observe that a necessary condition for a superchannel $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ to be an $\mathcal{R}$-subpreserving superchannel is $|B|\leq|D|$. Whereas, a necessary condition for a superchannel $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ to be an $\mathcal{R}$-preserving superchannel is $|B|=|D|$.
Proposition 8.
Let $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ be an $\mathcal{R}$-subpreserving superchannel, then the entropy of a quantum channel $\mathcal{N}_{A\to B}$ is nondecreasing under the action of $\Theta$, i.e.,
$S\left[\Theta(\mathcal{N}_{A\rightarrow B})\right]\geq S\left[\mathcal{N}_{A\rightarrow B}\right].$ (53)
Proof.
Let us assume that for a superchannel $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ the inequality $\Theta(\mathcal{R}_{A\rightarrow B})\leq\mathcal{R}_{C\rightarrow D}$ holds, where the maps $\mathcal{R}_{A\rightarrow B}\in\mathcal{L}\left(A,B\right)$ and $\mathcal{R}_{C\rightarrow D}\in\mathcal{L}\left(C,D\right)$ are completely depolarising maps. Then, for any positive operator $\rho_{RC}\in\mathcal{L}\left({R\otimes C}\right)$, where $R$ is a reference system, we have $(\mathrm{id}_{R}\otimes\Theta(\mathcal{R}_{A\rightarrow B}))(\rho_{RC})\leq(\mathrm{id}_{R}\otimes\mathcal{R}_{C\rightarrow D})(\rho_{RC})$. Now, for any $\epsilon>0$, adding $\epsilon\mathbbm{1}$ on both sides so as to make both sides strictly positive operators, we also have the following inequality: $\log\{\mathrm{id}_{R}\otimes\mathrm{\Theta}(\mathcal{R}_{A\rightarrow B})(\rho_{RC}+\epsilon\mathbbm{1})\}\leq\log\{\mathrm{id}_{R}\otimes\mathcal{R}_{C\rightarrow D}(\rho_{RC}+\epsilon\mathbbm{1})\}$, which further implies
$-\operatorname{tr}\left[\mathrm{id}_{R}\otimes\mathrm{\Theta}(\mathcal{N}_{A\rightarrow B})(\rho_{RC})\log\left(\mathrm{id}_{R}\otimes\mathrm{\Theta}(\mathcal{R}_{A\rightarrow B})(\rho_{RC}+\epsilon\mathbbm{1})\right)\right]$
$\geq-\operatorname{tr}\left[\mathrm{id}_{R}\otimes\mathrm{\Theta}(\mathcal{N}_{A\rightarrow B})(\rho_{RC})\log\left(\mathrm{id}_{R}\otimes\mathcal{R}_{C\rightarrow D}(\rho_{RC}+\epsilon\mathbbm{1})\right)\right].$ (54)
Now we add $\mathrm{tr}\left[\left(\mathrm{id}_{R}\otimes\Theta\left(\mathcal{N}_{A\rightarrow B}\right)\right)\left(\rho_{RA}\right)\log\left(\mathrm{id}\otimes\Theta\left(\mathcal{N}_{A\rightarrow B}\right)\right)\left(\rho_{RA}\right)\right]$ on both sides of the above inequality and take the limit $\epsilon\rightarrow 0$ according to Eq. (139). After that we take the supremum over $\rho_{RC}$ to arrive at the following inequality.
$\sup_{\rho_{RC}}D\left[\left(\mathrm{id}_{R}\otimes\Theta\left(\mathcal{N}_{A\rightarrow B}\right)\right)\left(\rho_{RC}\right)\;\middle\|\;\left(\mathrm{id}_{R}\otimes\Theta\left(\mathcal{R}_{A\rightarrow B}\right)\right)\left(\rho_{RC}\right)\right]$
$\geq\sup_{\rho_{RC}}D\left[\left(\mathrm{id}_{R}\otimes\Theta\left(\mathcal{N}_{A\rightarrow B}\right)\right)\left(\rho_{RC}\right)\;\middle\|\;\left(\mathrm{id}_{R}\otimes\mathcal{R}_{C\rightarrow D}\right)\left(\rho_{RC}\right)\right].$
Observe that the left-hand side of the above inequality is nothing but the relative entropy between quantum channels $\Theta\left(\mathcal{N}_{A\rightarrow B}\right)$ and $\Theta\left(\mathcal{R}_{A\rightarrow B}\right)$, and the right-hand side is the relative entropy between quantum channels $\Theta\left(\mathcal{N}_{A\rightarrow B}\right)$ and $\mathcal{R}_{C\rightarrow D}$. Using the monotonicity of relative entropy of quantum channels under the action of quantum superchannels, we get $D\left[\mathcal{N}_{A\rightarrow B}\;\middle\|\;\mathcal{R}_{A\rightarrow B}\right]\geq D\left[\Theta\left(\mathcal{N}_{A\rightarrow B}\right)\;\middle\|\;\mathcal{R}_{C\rightarrow D}\right]$, which implies that
$S\left[\Theta\left(\mathcal{N}_{A\rightarrow B}\right)\right]\geq S\left[\mathcal{N}_{A\rightarrow B}\right].$ (55)
This completes the proof of the proposition. ∎
The Proposition 8 is an expression for the entropy gain of quantum channels under the action of $\mathcal{R}$-subpreserving superchannels. As a special case, this proposition yields an expression for entropy gain of quantum channels under the action of random isometry superchannels, generalizing the result of entropy gain of quantum states under noisy operations. The proof relies on the fact that the random isometry superchannels are sub-completely randomizing map-preserving superchannels (see Proposition 7). Although it is clear that the entropy of quantum channels is nondecreasing under the action of random isometry superchannels it does not provide any estimate of change of entropy in terms of some remainder terms. In the following, we study the entropy gain of quantum channels in detail and estimate a remainder term for entropy gain for a class of quantum channels where we assume that the state that realizes the supremum has full rank reduced state. Note that the tele-covariant channels are an example of such class of channels where the supremum occurs at a maximally entangled state for which the reduced state is proportional to the identity operator. More specifically, we have the following theorem, which is one of our main results.
Theorem 1.
Let $\Theta:\mathcal{L}\left(A,A^{\prime}\right)\rightarrow\mathcal{L}\left(C,C^{\prime}\right)$ be a superchannel. Let $\mathcal{N}\in\mathcal{L}\left(A,A^{\prime}\right)$ and $\Theta\left(\mathcal{N}\right)\in\mathcal{L}\left(C,C^{\prime}\right)$ be two quantum channels. Let $\Psi_{\widetilde{A}A}\in\mathsf{FRank}\left(\widetilde{A}\otimes{A}\right)\cap\mathcal{D}\left(\widetilde{A}\otimes{A}\right)$ and $\Phi_{\widetilde{C}C}\in\mathsf{FRank}\left(\widetilde{C}\otimes{C}\right)\cap\mathcal{D}\left(\widetilde{C}\otimes{C}\right)$ are the pure states that realize the supremum in the entropy functional $S\left[\mathcal{N}\right]$ and $S\left[\Theta\left(\mathcal{N}\right)\right]$, respectively. Then
$S\left[\Theta\left(\mathcal{N}\right)\right]-S\left[\mathcal{N}\right]\geq D\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\;\middle\|\;\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)_{\alpha}\right)+\Delta^{\prime}\left[\Psi_{\widetilde{A}},\Phi_{\widetilde{C}}\right],$ (56)
where $\mathsf{C}_{\mathcal{N}}^{\Psi}\in\mathcal{D}(\widetilde{A}\otimes A)$ is the Choi state for channel $\mathcal{N}$, $\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)_{\alpha}=\frac{1}{\alpha^{\alpha}}\left(\mathfrak{T}^{*}\circ\mathfrak{T}\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)\right)^{\alpha}$ and $\alpha=\norm{\mathfrak{T}^{*}(\mathbbm{1})}_{\infty}$. $\mathfrak{T}$ is representing map of $\Theta$ and $\mathfrak{T}^{*}$ is its adjoint map, defined respectively in Eqs. (31) and (32). $\Delta^{\prime}\left[\Psi_{\widetilde{A}},\Phi_{\widetilde{C}}\right]:=S\left(\Psi_{\widetilde{A}}\right)-S\left(\Phi_{\widetilde{{C}}}\right)$, $\widetilde{A}\simeq A$, and $\widetilde{C}\simeq C$.
The proof of the theorem is presented in Appendix C.7. In the theorem above, we can lower bound $\Delta^{\prime}\left[\Psi_{\widetilde{A}},\Phi_{\widetilde{C}}\right]$ more succinctly if it can be proven that both the marginal states $\Psi_{\widetilde{A}}$ and $\Phi_{\widetilde{C}}$ are connected with a unital CP map. In Appendix C.6, we prove several entropy inequalities for the change in state under CP maps, unital CP maps, and positive maps. To prove the existence of a positive map that maps $\Phi_{\widetilde{C}}$ to $\Psi_{\widetilde{A}}$, we first argue that for $\Psi_{\widetilde{A}}\geq 0$ and $\Phi_{\widetilde{C}}>0$ (in fact, in our context we have $\Psi_{\widetilde{A}}>0$ and $\Phi_{\widetilde{C}}>0$), there always exists a completely positive map $\mathcal{F}:\mathcal{L}\left(\widetilde{C}\right)\rightarrow\mathcal{L}\left(\widetilde{A}\right)$, defined as
$\mathcal{F}(Y)=\left(\Psi_{\widetilde{A}}^{1/2}\Phi_{\widetilde{C}}^{-1/2}\right)Y\left(\Psi_{\widetilde{A}}^{1/2}\Phi_{\widetilde{C}}^{-1/2}\right)^{\dagger}~{}~{}\forall Y\in\mathcal{L}\left(\widetilde{C}\right)$
such that $\mathcal{F}\left(\Phi_{\widetilde{C}}\right)=\Psi_{\widetilde{A}}$. Since a completely positive map is also a positive map, we can use Theorem 3 of Appendix C.6 to arrive at the following inequality, $\Delta^{\prime}\left[\Psi_{\widetilde{A}},\Phi_{\widetilde{C}}\right]\geq D\left(\Psi_{\widetilde{C}}\;\middle\|\;\left(\Psi_{\widetilde{C}}\right)_{\alpha}\right)$, where $\alpha:=\norm{\mathcal{F}^{*}(\mathbbm{1}_{\widetilde{A}})}_{\infty}$ and $\left(\Psi_{\widetilde{C}}\right)_{\alpha}:=\left(\frac{1}{\alpha}\right)^{\alpha}\left(\mathcal{F}^{*}\circ\mathcal{F}\left(\Psi_{\widetilde{C}}\right)\right)^{\alpha}$. Now, we can modify the lower bound on channel entropy gain as
$S\left[\Theta\left(\mathcal{N}\right)\right]-S\left[\mathcal{N}\right]\geq D\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\;\middle\|\;\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)_{\alpha}\right)+D\left(\Psi_{\widetilde{C}}\;\middle\|\;\left(\Psi_{\widetilde{C}}\right)_{\alpha}\right).$ (57)
Deriving a non-negative lower bound on the entropy difference under the action of a superchannel as above holds significant implications for identifying a set of superchannels that are channel entropy nondecreasing. A trivial example of a channel entropy increasing superchannel is a superchannel that maps any input channel to a completely depolarising channel. If we consider superchannel $\Theta$ and channel $\mathcal{N}$ such that both the terms in the RHS of Eq. (57) are greater than or equal to 00, i.e., $D\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\;\middle\|\;\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)_{\alpha}\right)\geq 0$, and $D\left(\Psi_{\widetilde{C}}\;\middle\|\;\left(\Psi_{\widetilde{C}}\right)_{\alpha}\right)\geq 0$, we can easily conclude that superchannel $\Theta$ is entropy nondecreasing superchannel. As an example, in the following, we compute the lower bounds in Eq. (57) for superchannels that map tele-covariant channels to tele-covariant channels.
Example.
For a tele-covariant quantum channel (see Definition 1), the supremum in the entropy functional occurs at a maximally entangled state Gou19a . The entropy of a tele-covariant channel $\mathcal{N}_{A\to B}$ is then given as
$S\left[\mathcal{N}\right]$ $=S\left(\mathrm{id}_{\widetilde{A}}\otimes\mathcal{N}\left(|A|^{-1}\Psi^{+}_{\widetilde{A}A}\right)\right)-S\left(|A|^{-1}\mathbbm{1}_{A}\right)$
$=S\left(\mathsf{C}_{\mathcal{N}}\right)-\log|A|,$ (58)
where $|A|^{-1}\Psi^{+}_{\widetilde{A}A}$ is a maximally entangled state and $\mathsf{C}_{\mathcal{N}}$ is the Choi state for channel $\mathcal{N}$. Let $\mathrm{cov}\left(\mathcal{L}(A,B)\right)$ and $\mathrm{cov}\left(\mathcal{L}(C,D)\right)$ denote the set of all tele-covariant channels in $\mathcal{L}(A,B)$ and $\mathcal{L}\left(C,D\right)$, respectively. Now, consider a tele-covariance preserving superchannel $\Theta:\mathrm{cov}\left(\mathcal{L}(A,B)\right)\rightarrow\mathrm{cov}\left(\mathcal{L}(C,D)\right)$, i.e., a superchannel that maps a tele-covariant channel to a tele-covariant channel. The entropy of the tele-covariant channel $\Theta(\mathcal{N})$ is the following,
$S\left[\Theta\left(\mathcal{N}\right)\right]$ $=S\left(\left(\mathrm{id}_{\widetilde{C}}\otimes\Theta\left(\mathcal{N}\right)\right)\left(|C|^{-1}\Psi^{+}_{\widetilde{C}C}\right)\right)-S\left(|C|^{-1}\mathbbm{1}_{C}\right)$
where $|C|^{-1}\Psi^{+}_{\widetilde{C}C}$ is a maximally entangled state. Using Theorem 1, we can write the gain in the channel entropy as
$).\displaystyle S\left[\Theta\left(\mathcal{N}\right)\right]-S\left[\mathcal{N}\right]\geq D\left(\mathsf{C}_{\mathcal{N}}\;\middle\|\;\left(\mathsf{C}_{\mathcal{N}}\right)_{\alpha}\right)+\log{\frac{|A|}{|C|}}.$ (60)
Observe that $\Delta^{\prime}\left[|A|^{-1}\Psi^{+}_{\widetilde{A}},|C|^{-1}\Psi^{+}_{\widetilde{C}}\right]$ is already a constant in this case and equal to $\log\widetilde{d}$, where $\widetilde{d}=|A|/|C|$. Further, we have $\left(\mathsf{C}_{\mathcal{N}}\right)_{\alpha}=\left(\frac{1}{\alpha}\right)^{\alpha}\left(\mathsf{C}_{\mathrm{\Theta}^{*}\circ\mathrm{\Theta}(\mathcal{N})}\right)^{\alpha}$, where $\alpha=\norm{\mathsf{C}_{\Theta^{*}(\mathcal{R})}}_{\infty}$. Notice that if superchannel $\Theta$ is such that $|A|\geq|C|$, then we always have $\log\widetilde{d}\geq 0$. Note that if $\operatorname{tr}\left(\left(\mathsf{C}_{\mathcal{N}}\right)_{\alpha}\right)\leq 1$, then we have $D\left(\mathsf{C}_{\mathcal{N}}\;\middle\|\;\left(\mathsf{C}_{\mathcal{N}}\right)_{\alpha}\right)\geq 0$.
III.3 Entropy of continuous-variable quantum channels
In this section, we show that our definition of the entropy of discrete-variable quantum channels is easily extended to continuous-variable quantum channels, which is analogous to the definition of the entropy for quantum states in terms of relative entropy.
Let $A$ be a finite- or infinite-dimensional, separable Hilbert space and $B$ be an infinite-dimensional, separable Hilbert space. Let $\widehat{H}_{B}$ be a hermitian operator representing the Hamiltonian associated with the quantum system $B$; the lowest eigenvalue of $\widehat{H}_{B}$ is bounded from below by 00. Then, a thermal operator $\widehat{\tau}_{B}^{\beta}$ for $B$ with an inverse temperature $\beta\in[0,\infty]$ is given as
$\widehat{\tau}_{B}^{\beta}=\operatorname{e}^{-\beta\widehat{H}_{B}}.$ (61)
A thermal operator $\widehat{\tau}_{B}^{\beta}$ is bounded and trace-class for all $\beta\in(0,\infty)$. We note that for $\dim(B)=\infty$ and as $\beta\to 0^{+}$, we get
$\lim_{\beta\to 0^{+}}\widehat{\tau}_{B}^{\beta}=\mathbbm{1}_{B}.$ (62)
Let us consider a completely thermalizing map $\mathcal{R}^{\beta}_{A\to B}:\mathcal{L}({A})_{+}\to\mathcal{L}({B})_{+}$, whose action on all trace-class operator $\mathsf{X}_{A}\in\mathcal{L}({A})_{+}$ is given by
$\mathcal{R}^{\beta}_{A\to B}(\mathsf{X}_{A})=\operatorname{tr}[\mathsf{X}_{A}]\widehat{\tau}_{B}^{\beta}.$ (63)
We can now define $\beta$-parameterized generalized entropy functional $\mathbf{S}_{\beta}$ of an arbitrary quantum channel $\mathcal{N}_{A\to B}$, where $\dim(B)=\infty$, for $\beta\in(0,\infty)$ as
$\mathbf{S}_{\beta}[\mathcal{N}]:=-\mathbf{D}\left[\mathcal{N}_{A\to B}\;\middle\|\;\mathcal{R}^{\beta}_{A\to B}\right]=-\sup_{\rho_{RA}\in\mathcal{D}(R\otimes A)}\mathbf{D}\left(\operatorname{id}_{R}\otimes\mathcal{N}_{A\to B}(\rho_{RA})\;\middle\|\;\operatorname{id}_{R}\otimes\mathcal{R}^{\beta}_{A\to B}(\rho_{RA})\right),$ (64)
where $R$ is of arbitrary dimension and $\mathbf{D}(\cdot\|\cdot)$ is a generalized divergence functional. It easily follows that
$\mathbf{D}\left(\operatorname{id}_{R}\otimes\mathcal{N}_{A\to B}(\rho_{RA})\;\middle\|\;\operatorname{id}_{R}\otimes\mathcal{R}^{\beta}_{A\to B}(\rho_{RA})\right)=\mathbf{D}\left(\mathcal{N}_{A\to B}(\rho_{RA})\;\middle\|\;\rho_{R}\otimes\widehat{\tau}^{\beta}_{B}\right).$
Definition 7.
The generalized entropy functional $\mathbf{S}[\mathcal{N}]$ of an arbitrary quantum channel $\mathcal{N}_{A\to B}$, where $|B|=\infty$, is defined as
$\mathbf{S}[\mathcal{N}]:=\lim_{\beta\to 0^{+}}\mathbf{S}_{\beta}[\mathcal{N}]=-\lim_{\beta\to 0^{+}}\mathbf{D}\left[\mathcal{N}\;\middle\|\;\mathcal{R}^{\beta}\right].$ (65)
Similarly, if the input states to the channel are accessible only from a set $\mathsf{Access}(R\otimes A)\subset\mathcal{D}(R\otimes A)$ of states because of some physical constraints, then we define the resource-constrained generalized entropy functional $\mathbf{S}_{\mathsf{Access}}[\mathcal{N}]$ of an arbitrary quantum channel $\mathcal{N}_{A\to B}$, where inputs to $\mathbf{D}[\mathcal{N}\|\mathcal{R}^{\beta}]$ are only from $\mathsf{Access}(R\otimes A)$ (cf. Win (16); GW (21)), as
$\mathbf{S}_{\mathsf{Access}}[\mathcal{N}]:=\lim_{\beta\to 0^{+}}\mathbf{S}_{{}_{\mathsf{Access}},\beta}[\mathcal{N}]=-\lim_{\beta\to 0^{+}}\mathbf{D}_{\mathsf{Access}}[\mathcal{N}~{}\|~{}\mathcal{R}^{\beta}],$ (66)
where
$\mathbf{D}_{\mathsf{Access}}[\mathcal{N}_{A\to B}~{}\|~{}\mathcal{R}^{\beta}_{A\to B}]=\sup_{\rho_{RA}\in\mathsf{Access}(R\otimes A)}\mathbf{D}\left(\operatorname{id}_{R}\otimes\mathcal{N}_{A\to B}(\rho_{RA})\;\middle\|\;\operatorname{id}_{R}\otimes\mathcal{R}^{\beta}_{A\to B}(\rho_{RA})\right).$ (67)
IV Sufficiency for the recoverability of quantum channels
The data processing inequality states that the relative entropy between two states $\rho$ and $\sigma$ is nonincreasing under CPTP maps, i.e., $D\left(\rho\;\middle\|\;\sigma\right)\geq D\left(\mathcal{N}\left(\rho\right)\;\middle\|\;\mathcal{N}\left(\sigma\right)\right),$ where $\mathcal{N}\in\mathcal{L}(A,B)$ is some quantum channel. The necessary and sufficient condition for the saturation of this inequality is related to the sufficiency (or reversibility) of quantum channel $\mathcal{N}$ Pet (86); MP (04); Pet (88). A quantum channel is called sufficient for a set of states $\{\rho_{i}\}_{i}$ if there exists another CPTP map $\mathcal{P}$ called the recovery map, such that for each $\rho_{i}$, we have $\left(\mathcal{P}\circ\mathcal{N}\right)(\rho_{i})=\rho_{i}$. Now, it can be immediately checked that the quantum channel $\mathcal{N}$ preserves the relative entropy between $\rho$ and $\sigma$ if there exists a recovery map $\mathcal{P}$ for the set $\{\rho,\sigma\}$. Conversely, if quantum channel $\mathcal{N}$ does not change the relative entropy between $\rho$ and $\sigma$, then its action on the set $\{\rho,\sigma\}$ is reversible and the recovery map, called the Petz recovery map $\mathcal{P}_{\sigma,\mathcal{N}}$, is known in terms of $\sigma$ and $\mathcal{N}$ Pet (86, 88). The Petz recovery map $\mathcal{P}_{\sigma,\mathcal{N}}:\mathcal{L}(B)\rightarrow\mathcal{L}(A)$ is a linear, completely positive, and trace nonincreasing map defined as
$\mathcal{P}_{\sigma,\mathcal{N}}\left(\mathsf{X}\right)\equiv\sigma^{\frac{1}{2}}\mathcal{N}^{*}\left(\left(\mathcal{N}\left(\sigma\right)\right)^{\frac{1}{2}}\mathsf{X}\left(\mathcal{N}\left(\sigma\right)\right)^{\frac{1}{2}}\right)\sigma^{\frac{1}{2}},$ (68)
where $\mathcal{N}^{*}$ is the adjoint map of $\mathcal{N}$. In recent years, some refinement of data processing inequality have also been proposed in terms of a recovery map Wil (15); JRS+ (18); BLW (15); STH (16). In  JRS+ (18), it was shown that there exists a universal recovery map (channel) $\mathcal{P}^{\mathrm{R}}$ such that the monotonicity of the quantum relative entropy between two arbitrary quantum states $\rho$ and $\sigma$ under the action of an arbitrary quantum channel $\mathcal{N}\in\mathcal{L}(A,B)$ can be refined as
$D\left(\rho\;\middle\|\;\sigma\right)-D(\mathcal{N}\left(\rho\right)\;\|\;\mathcal{N}\left(\sigma\right))\geq-\log F\left(\rho,\left(\mathcal{P}^{\mathrm{R}}\circ\mathcal{N}\right)(\rho)\right),$ (69)
where $F(\rho,\sigma):=\norm{\sqrt{\rho}\sqrt{\sigma}}^{2}_{1}$ is the fidelity between $\rho,\sigma\in\mathcal{D}(A)$. Here, the recovery channel $\mathcal{P}^{\mathrm{R}}$ is given by JRS+ (18) (see also Wil (13))
[…]
where $\xi\in\mathcal{D}\left(A\right)$, $\mathsf{X}\in\mathcal{L}(B)$, $\Pi_{\mathcal{N}(\sigma)}$ is the projection onto the support of $\mathcal{N}(\sigma)$, and $\alpha_{t}\equiv\pi\left(\cosh(\pi t)+1\right)^{-1}/2$ can be interpreted as probability density function on $t\in\mathbb{R}$. And $\mathcal{P}^{t}_{\sigma,\mathcal{N}}\equiv\mathcal{U}_{\sigma,-t}\circ\mathcal{P}_{\sigma,\mathcal{N}}\circ\mathcal{U}_{\mathcal{N}(\sigma),t}$, where $\mathcal{U}_{\sigma,t}$ is a partial isometric map defined as $\mathcal{U}_{\sigma,t}(\mathsf{X})=\sigma^{it}\mathsf{X}\sigma^{-it}$. The fidelity between any two arbitrary states $\rho,\sigma$ is always bounded as $F(\rho,\sigma)\in[0,1]$, with $F(\rho,\sigma)=1$ iff $\rho=\sigma$ and $F(\rho,\sigma)=0$ iff $\rho\perp\sigma$. The term $\operatorname{tr}\left(\left(\mathbbm{1}_{A}-\Pi_{\mathcal{N}(\sigma)}\right)\mathsf{X}\right)\xi$ present in the above equation is there to ensure $\mathcal{P}^{\mathrm{R}}$ is trace preserving. It can be easily checked that the recovery channel $\mathcal{P}^{\mathrm{R}}$ perfectly recovers $\sigma$ from $\mathcal{N}(\sigma)$, i.e., $\left(\mathcal{P}^{\mathrm{R}}\circ\mathcal{N}\right)(\sigma)=\sigma$, and also if $\rho$ and $(\mathcal{P}^{\mathrm{R}}\circ\mathcal{N})(\rho)$ are close then $-\log F\left(\rho,(\mathcal{P}^{\mathrm{R}}\circ\mathcal{N})(\rho)\right)$ will be a small number. So, the above refinement says that it is possible to perform a recovery operation in which one can perfectly recover one state and approximately recover the other states if the change in the quantum relative entropy between the two quantum states before and after the action of the quantum channel is small Wil (15); JRS+ (18).
Motivated by these, one can ask the following question: is it possible to reverse the action of a quantum superchannel $\Theta$ on some given pair of channels $\{\mathcal{N},\mathcal{M}\}$? To answer this question, we first define the notion of sufficiency of superchannels, in the similar spirit of the sufficiency of the quantum channels, as follows. A superchannel $\Theta$ is called sufficient for a set of channels $\{\mathcal{N}_{i}\}_{i}$ if there exists another superchannel $\Theta^{\mathrm{R}}$, such that for each $\mathcal{N}_{i}$, we have $\left(\Theta^{\mathrm{R}}\circ\Theta\right)\left(\mathcal{N}_{i}\right)=\mathcal{N}_{i}$. Note that the relative entropy between two quantum channels $\mathcal{N},\mathcal{M}\in\mathcal{L}\left(A,A^{\prime}\right)$ is also nonincreasing under the action of a quantum superchannel $\Theta:\mathcal{L}\left(A,A^{\prime}\right)\rightarrow\mathcal{L}\left(C,C^{\prime}\right)$, i.e., $D\left[\Theta\left(\mathcal{N}\right)\;\middle\|\;\Theta\left(\mathcal{M}\right)\right]\leq D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]$ (see Proposition 1, and also CMW (16); LKDW (18); Yua (19)). If the action of the superchannel $\Theta$ on the set of channels $\{\mathcal{N},\mathcal{M}\}$ can be reversed, i.e., if there exists another superchannel $\Theta^{\mathrm{R}}:\mathcal{L}\left(C,C^{\prime}\right)\rightarrow\mathcal{L}\left(A,A^{\prime}\right)$ such that $\left(\Theta^{\mathrm{R}}\circ\Theta\right)\left(\mathcal{N}\right)=\mathcal{N}$, and $\left(\Theta^{\mathrm{R}}\circ\Theta\right)\left(\mathcal{M}\right)=\mathcal{M}$, then one can immediately show that the superchannel $\Theta$ does not change the relative entropy between channels $\mathcal{N}$ and $\mathcal{M}$. Here we prove that the converse is also true, i.e., if $D\left[\Theta\left(\mathcal{N}\right)\;\middle\|\;\Theta\left(\mathcal{M}\right)\right]=D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]$, then there exists a supermap $\Theta^{\mathrm{R}}:\mathcal{L}\left(C,C^{\prime}\right)\rightarrow\mathcal{L}\left(A,A^{\prime}\right)$ such that $\left(\Theta^{\mathrm{R}}\circ\Theta\right)\left(\mathcal{N}\right)=\mathcal{N}$, and $\left(\Theta^{\mathrm{R}}\circ\Theta\right)\left(\mathcal{M}\right)=\mathcal{M}$, for a particular class of channels. First, we prove a refinement of the data processing inequality for the relative entropy of quantum channels as the following theorem.
Theorem 2.
Let $\Theta:\mathcal{L}\left(A,A^{\prime}\right)\rightarrow\mathcal{L}\left(C,C^{\prime}\right)$ be a superchannel in the class $\mathsf{SCT}$ (see Sec. II) and $\mathcal{N}_{A\rightarrow A^{\prime}}$, and $\mathcal{M}_{A\rightarrow A^{\prime}}$ be two quantum channels. Let $\Psi_{\widetilde{A}A}\in\mathsf{FRank}\left(\widetilde{A}\otimes{A}\right)\cap\mathcal{D}\left(\widetilde{A}\otimes{A}\right)$ and $\Phi_{\widetilde{C}C}\in\mathsf{FRank}\left(\widetilde{C}\otimes{C}\right)\cap\mathcal{D}\left(\widetilde{C}\otimes{C}\right)$ be the pure states that realize the supremum in the relative entropy functional $D\left[\Theta\left(\mathcal{N}\right)\;\middle\|\;\Theta\left(\mathcal{M}\right)\right]$ and $D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]$, respectively. Then we have
$D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]-D\left[\Theta\left(\mathcal{N}\right)\;\middle\|\;\Theta\left(\mathcal{M}\right)\right]\geq-\log F\left(\mathsf{C}^{\Psi}_{\mathcal{N}},\left(\mathcal{P}^{\mathrm{R}}\circ\mathfrak{T}^{\prime}\right)\left(\mathsf{C}^{\Psi}_{\mathcal{N}}\right)\right),$ (71)
where $F(\rho,\sigma):=\norm{\sqrt{\rho}\sqrt{\sigma}}^{2}_{1}$ is the fidelity of states $\rho$ and $\sigma$, $\mathsf{C}_{\mathcal{N}}^{\Psi}$ is the Choi state for channel $\mathcal{N}_{A\rightarrow A^{\prime}}$ with respect to $\Psi$ and $\mathcal{P}^{\mathrm{R}}$ is the recovery map defined by Eq. (70). $\mathfrak{T}^{\prime}(\mathsf{X}):=\mathfrak{T}(\mathsf{X})+\left[\mathrm{tr}(\mathsf{X})-\mathrm{tr}(\mathfrak{T}(\mathsf{X}))\right]\sigma_{0}$, where $\sigma_{0}\in\mathcal{L}(C\otimes C^{\prime})$ is chosen in such a way that $\mathfrak{T}^{\prime}$ is a CPTP map (see Appendix B.4) and $\mathfrak{T}:\mathcal{L}\left(A\otimes A^{\prime}\right)\rightarrow\mathcal{L}\left(C\otimes C^{\prime}\right)$ is the representing map for superchannel $\Theta$ (see Remark 3 and Eq. (29)).
The detailed proof of this theorem is provided in Appendix  D while we sketch the proof of this theorem in the following. The basic idea behind the proof is the fact that a supermap induces a linear map, called the representing map, at the level of Choi matrices. The representing map $\mathfrak{T}$ of the superchannel $\Theta$ is in general a CP map that becomes trace preserving if the adjoint of $\Theta$ satisfies a simple condition (see Eq. (118) of Appendix III). This, in turn, implies that $\Theta^{*}$ is a completely depolarising map preserving (see Lemma 6) if $\Psi$ and $\Phi$ are maximally entangled states. Even if this trace preserving condition of the representing map $\mathfrak{T}$ is not satisfied, since we have $\Theta\in\mathsf{SCT}$, we can always replace the representing map $\mathfrak{T}$ with the following CPTP map: $\mathfrak{T}^{\prime}(\mathsf{X}):=\mathfrak{T}(\mathsf{X})+\left[\mathrm{tr}(\mathsf{X})-\mathrm{tr}(\mathfrak{T}(\mathsf{X}))\right]\sigma_{0}$ for all $\mathsf{X}\in\mathcal{L}(A\otimes A^{\prime})$. Here $\sigma_{0}\in\mathcal{L}(C\otimes C^{\prime})$ is chosen such that $\mathfrak{T}^{\prime}$ is a CPTP map (see Appendix B.4) and we further have $\mathfrak{T}^{\prime}(\mathsf{C}^{\Psi}_{\mathcal{N}})=\mathfrak{T}(\mathsf{C}^{\Psi}_{\mathcal{N}})$ and $\mathfrak{T}^{\prime}(\mathsf{C}^{\Psi}_{\mathcal{M}})=\mathfrak{T}(\mathsf{C}^{\Psi}_{\mathcal{M}})$. The recovery map $\mathcal{P}^{\mathrm{R}}$ is defined in such a way that it always recovers $\mathsf{C}^{\Psi}_{\mathcal{M}}$ from $\mathfrak{T}^{\prime}(\mathsf{C}^{\Psi}_{\mathcal{M}})$.
Remark 5.
The Theorem 2 provides the necessary and sufficient condition for the sufficiency of quantum superchannels. We can define a supermap $\Theta^{\mathrm{R}}:\mathcal{L}\left(C,C^{\prime}\right)\rightarrow\mathcal{L}\left(A,A^{\prime}\right)$, which we call the recovery supermap, corresponding to the recovery map $\mathcal{P}^{\mathrm{R}}$, such that it always recovers $\mathcal{M}$ from $\Theta\left(\mathcal{M}\right)$. The action of such a recovery supermap on $\widetilde{\mathcal{N}}$ is given as follows.
$\Theta^{\mathrm{R}}\left(\widetilde{\mathcal{N}}\right)(\mathsf{A})=\mathrm{tr}_{C}\left((\mathsf{A}^{t}\otimes\mathbbm{1})T^{-1}_{\Psi}\mathcal{P}^{\mathrm{R}}\left(\mathsf{C}^{\Phi}_{\widetilde{\mathcal{N}}}\right)\right).$ (72)
It can be verified that if $\left(\mathcal{P}^{\mathrm{R}}\circ\mathfrak{T}\right)\left(\mathsf{C}^{\Psi}_{\mathcal{N}}\right)=\mathsf{C}^{\Psi}_{\mathcal{N}}$ then $\Theta^{\mathrm{R}}$ also recovers $\mathcal{N}$ from $\Theta\left(\mathcal{N}\right)$. So, the above refinement says that if the data processing inequality saturates, i.e., $D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]=D\left[\Theta\left(\mathcal{N}\right)\;\middle\|\;\Theta\left(\mathcal{M}\right)\right]$, then there exists a recovery supermap $\Theta^{\mathrm{R}}$ such that
$\Theta^{\mathrm{R}}\circ\Theta\left(\mathcal{N}\right)=\mathcal{N}~{}~{}\text{and}~{}~{}\Theta^{\mathrm{R}}\circ\Theta\left(\mathcal{M}\right)=\mathcal{M}.$ (73)
An immediate consequence of the above theorem is the following proposition.
Proposition 9.
Let $\mathcal{N},\mathcal{M}\in\mathcal{L}\left(A,A^{\prime}\right)$ be two tele-covariant quantum channels (with respect to the same unitary representation of a group) (DW19a, , Definition 4), and let $\Theta:\mathcal{L}\left(A,A^{\prime}\right)\rightarrow\mathcal{L}\left(C,C^{\prime}\right)$ be a tele-covariance preserving superchannel and $\Theta\in\mathsf{SCT}$. Then we have
$D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]-D\left[\Theta\left(\mathcal{N}\right)\;\middle\|\;\Theta\left(\mathcal{M}\right)\right]\geq-\log F\left(\mathsf{C}_{\mathcal{N}},\left(\mathcal{P}^{\mathrm{R}}\circ\mathfrak{T}^{\prime}\right)\left(\mathsf{C}_{\mathcal{N}}\right)\right),$ (74)
where $\mathsf{C}_{\mathcal{N}}$ is the Choi state for channel $\mathcal{N}_{A\rightarrow A^{\prime}}$, and $\mathcal{P}^{\mathrm{R}}$ and $\mathfrak{T}^{\prime}$ are defined in Theorem 2.
The proof of the proposition follows directly from the Theorem 2 by noting the following: when we take channels $\mathcal{N}$ and $\mathcal{M}$ to be covariant with respect to the same unitary representation of a group and the superchannel $\Theta$ to be a tele-covariant channel preserving superchannel, $\mathsf{C}^{\Psi}_{\mathcal{N}}$ becomes $\mathsf{C}_{\mathcal{N}}$ and the recovery supermap $\Theta^{\mathrm{R}}$ changes accordingly.
Next, we estimate the entropy gain of a tele-covariant channel under the action of tele-covariant channel preserving superchannels as an application of Theorem 2. In particular, we have the following lemma.
Lemma 7.
Let $\mathcal{N}\in\mathcal{L}\left(A,A^{\prime}\right)$ be a tele-covariant channel, $\Theta:\mathcal{L}\left(A,A^{\prime}\right)\rightarrow\mathcal{L}\left(C,C^{\prime}\right)$ be a tele-covariant channel preserving superchannel and $\Theta\in\mathsf{SCT}$. Then change in the entropy of quantum channel $\mathcal{N}$ is bounded as
$S\left[\Theta\left(\mathcal{N}\right)\right]-S\left[\mathcal{N}\right]\geq D\left(\mathsf{C}_{\mathcal{N}}\;\middle\|\;\left(\widetilde{\mathcal{P}}^{\mathrm{R}}\circ\mathsf{T}^{\prime}\right)\mathsf{C}_{\mathcal{N}}\right),$ (75)
where $\mathsf{C}_{\mathcal{N}}$ is the Choi state for channel $\mathcal{N}$ and $\mathsf{T}\in\mathcal{L}\left(A\otimes A^{\prime},C\otimes C^{\prime}\right)$ is the representing map for superchannel $\Theta$ (see Remark 3 and Eq. (29)). $\mathsf{T}^{\prime}(\mathsf{X}):=\mathsf{T}(\mathsf{X})+\left[\mathrm{tr}(\mathsf{X})-\mathrm{tr}(\mathsf{T}(\mathsf{X}))\right]\sigma_{0}$, where $\sigma_{0}\in\mathcal{L}(C\otimes C^{\prime})$ is chosen such that $\mathsf{T}^{\prime}$ is a CPTP map (see Appendix B.4). The recovery map $\widetilde{\mathcal{P}}^{\mathrm{R}}\left(\mathsf{X}\right)$ (see (BDW, 16, Eq. (36))) is given by $\widetilde{\mathcal{P}}^{\mathrm{R}}\left(\mathsf{X}\right)={\mathsf{T}^{\prime}}^{*}(\mathsf{X})+\operatorname{tr}\left(\left(\mathrm{id}-{\mathsf{T}^{\prime}}^{*}\right)\mathsf{X}\right)\xi$, where $\xi\in\mathcal{L}(C\otimes C^{\prime})$, $\mathsf{X}\in\mathcal{L}(C\otimes C^{\prime})$.
Proof.
Since $\mathcal{N}$ is a tele-covariant channel and $\Theta$ is a tele-covariant channel preserving superchannel, we have $S\left[\mathcal{N}\right]=S(\mathsf{C}_{\mathcal{N}})-\log\left(|A|\right)$ and $S\left[\Theta\left(\mathcal{N}\right)\right]=S\left(\mathsf{C}_{\Theta(\mathcal{N})}\right)-\log\left(|C|\right)$, where $\mathsf{C}_{\mathcal{N}}$ and $\mathsf{C}_{\Theta(\mathcal{N})}$ are the Choi states of $\mathcal{N}$ and $\Theta(\mathcal{N})$, respectively. Then the entropy difference becomes $S\left[\Theta\left(\mathcal{N}\right)\right]-S\left[\mathcal{N}\right]=S\left(\mathsf{C}_{\Theta(\mathcal{N})}\right)-S(\mathsf{C}_{\mathcal{N}})+\log\widetilde{d}$, where $\widetilde{d}:=|A|/|C|$. Also, $\mathsf{C}_{\mathcal{N}}$ and $\mathsf{C}_{\Theta(\mathcal{N})}$ are connected by the representing map $\mathsf{T}$ which is CP but not trace preserving, in general. So we replace it with a channel $\mathsf{T}^{\prime}$ such that $\mathsf{T}{{}^{\prime}}(\mathsf{C}_{\Psi})=\mathsf{T}(\mathsf{C}_{\Psi})$ holds when $|A|=|C|$ (see Eq. (20)). The entropy difference then is lower bounded by (see (BDW, 16, Eq. (37)), also DKSW (18))
$S\left[\Theta\left(\mathcal{N}\right)\right]-S\left[\mathcal{N}\right]\geq D\left(\mathsf{C}_{\mathcal{N}}\;\middle\|\;\left(\widetilde{\mathcal{P}}^{\mathrm{R}}\circ\mathsf{T}^{\prime}\right)\left(\mathsf{C}_{\mathcal{N}}\right)\right),$ (76)
where each term in the above inequality is defined in the statement of the lemma. This completes the proof. ∎
The Lemma 7 tells that if the entropy difference before and after applying the superchannel is considerably small, then one can conclude that states $\mathsf{C}_{\mathcal{N}}$ and $\left(\widetilde{\mathcal{P}}^{\mathrm{R}}\circ\mathsf{T}^{\prime}\right)(\mathsf{C}_{\mathcal{N}})$ are very close. That is by performing the recovery operation $\widetilde{\mathcal{P}}^{\mathrm{R}}$, one can approximately recover the state $\mathsf{C}_{\mathcal{N}}$, or equivalently it means that there exists a recovery supermap $\Theta^{\mathrm{R}}$, defined as $\Theta^{\mathrm{R}}(\widetilde{\mathcal{N})}(\mathsf{A})=\absolutevalue{A}\mathrm{tr}_{C}\left((\mathsf{A}^{t}\otimes\mathbbm{1})\widetilde{\mathcal{P}}^{\mathrm{R}}(\mathsf{C}_{\widetilde{\mathcal{N}}})\right)$ that can approximately recover $\mathcal{N}$ from $\Theta(\mathcal{N})$. On more application side, noting that quantum channels form an indispensable part in any quantum network (recall that states, measurements and all physical transformations can be understood as linear maps in the quantum domain), it is of significant importance to protect quantum channels from unwanted noise (some superchannel). Our inequalities provide a necessary and sufficient condition for this task in terms of entropy and relative entropy of channels.
V Generalized divergence and entropy of superchannels
In quantum information theory, the information is usually encoded in the state of the quantum system and then any quantum information processing task may be understood simply as a collection of parallel and sequential transformations (linear CPTP maps) acting on the quantum states. However, one can encode quantum information in quantum channels too and then any quantum information processing task may be understood as a collection of parallel and sequential quantum channels and quantum superchannels acting on quantum states and quantum channels, respectively. In this light, one may think quantum channels as first order transformations and quantum superchannels as second order transformations. In fact, one can recursively do this to obtain arbitrary finite order processes mapping a lower order process to another process of the same lower order. Such a paradigm has attracted considerable attention recently, see e.g. CDP08a ; CDP (09); BP (19); Jen (21); HO (22).
In Section III, we generalized the definition of entropy to quantum channels, where we argued that similar to the entropy of quantum states, the entropy of quantum channels can also be defined in terms of relative entropy between the given quantum channel and the completely depolarising map (see Definition 46). Here, we first extend the definition of generalized divergences to superchannels using the generalized divergence between two quantum channels. Recall that a map $\mathbf{D}:\mathcal{L}\left(A\right)_{+}\times\mathcal{L}\left(A\right)_{+}\rightarrow\mathbb{R}$ is called a generalized state divergence if it satisfies data processing inequality: $\mathbf{D}\left(\rho\;\middle\|\;\sigma\right)\geq\mathbf{D}\left(\mathcal{N}(\rho)\;\middle\|\;\mathcal{N}(\sigma)\right)$, where $\rho,\sigma\in\mathcal{L}\left(A\right)_{+}$ and $\mathcal{N}:\mathcal{L}(A,A^{\prime})$ is a quantum channel. Given a generalized state divergence $\mathbf{D}$, we define the generalized divergence between a quantum channel $\mathcal{N}\in\mathcal{L}\left(A,A^{\prime}\right)$ and a CP map $\mathcal{M}\in\mathcal{L}\left(A,A^{\prime}\right)$ as
$\mathbf{D}\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]=\sup_{\Psi_{RA}\in\mathcal{D}(R\otimes A)}\mathbf{D}\left(\left(\mathrm{id}_{R}\otimes\mathcal{N}\right)(\Psi_{RA})\;\middle\|\;\left(\mathrm{id}_{R}\otimes\mathcal{M}\right)(\Psi_{RA})\right),$ (77)
where $R$ is a reference system of arbitrary dimension. Now, given a generalized divergence $\mathbf{D}$ between channels (cf. CMW (16); LKDW (18); Yua (19)), we define a divergence functional between a quantum superchannel $\Theta$ and a supermap $\Gamma$, denoted by $\mathbf{D}^{(2)}$, as follows.
Definition 8.
Given a generalized divergence $\mathbf{D}\left[\cdot\;\middle\|\;\cdot\right]$ between a quantum channel and a CP map (as in Eq. (77)), a divergence between a quantum superchannel $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ and a completely CP-preserving supermap $\Gamma:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ is defined as
$\mathbf{D}^{(2)}\left[\Theta\;\middle\|\;\Gamma\right]=\sup_{\mathcal{N}\in\mathcal{L}(R\otimes A,R\otimes B)}\mathbf{D}\left[\left(\mathrm{id}\otimes\Theta\right)\left(\mathcal{N}\right)\;\middle\|\;\left(\mathrm{id}\otimes\Gamma\right)\left(\mathcal{N}\right)\right],$ (78)
where the supremum is taken over the whole space $\mathcal{L}(R\otimes A,R\otimes B)$ of bipartite channels, $\mathrm{id}$ is the identity quantum superchannel, and $R$ is an arbitrary Hilbert space.
We note that generalized divergence between superchannels was also defined in Hir (23), but our definition differs notably and can be recursively generalized to higher-order processes. Let us first consider a class $\mathsf{SSChan}$ of super-superchannels, i.e., a class of higher-order processes that map superchannels in $\mathscr{L}\left[\mathcal{L}\left(A,B\right),\mathcal{L}\left(C,D\right)\right]$ to superchannels in $\mathscr{L}\left[\mathcal{L}\left(A^{\prime},B^{\prime}\right),\mathcal{L}\left(C^{\prime},D^{\prime}\right)\right]$ such that for any $\mathring{\mathbf{\Lambda}}\in\mathsf{SSChan}$ we can write
$\mathring{\mathbf{\Lambda}}\left(\Theta\right)=\Lambda_{\text{post}}\circ(\Theta\otimes\operatorname{id})\circ\Lambda_{\text{pre}}~{}~{}~{}\forall~{}\Theta\in\mathscr{L}\left[\mathcal{L}\left(A,B\right),\mathcal{L}\left(C,D\right)\right],$ (79)
where $\Lambda_{\text{pre}}\in\mathscr{L}\left[\mathcal{L}\left(A^{\prime},B^{\prime}\right),\mathcal{L}\left(A\otimes R,B\otimes R\right)\right]$ and $\Lambda_{\text{post}}\in\mathscr{L}\left[\mathcal{L}\left(C\otimes R,D\otimes R\right),\mathcal{L}\left(C^{\prime},D^{\prime}\right)\right]$ are preprocessing and postprocessing superchannels, respectively, similar to the preprocessing and postprocessing quantum channels of Eq. (13). $R$ is the reference system on which $\operatorname{id}$ acts. Looking at the mathematical similarity to quantum channels, we expect that the dilation equality, Eq. (79), will hold for all quantum super-superchannels. But, for the rest of this section, we restrict ourselves to the class $\mathsf{SSChan}$ of super-superchannels. For any pair of superchannels $\Theta,\Gamma\in\mathscr{L}\left[\mathcal{L}\left(A,B\right),\mathcal{L}\left(C,D\right)\right]$, our definition of the divergence $\mathbf{D}^{(2)}$ for superchannels satisfies the monotonicity relation $\mathbf{D}^{(2)}\left[\Theta\;\middle\|\;\Gamma\right]\geq\mathbf{D}^{(2)}\left[\mathring{\mathbf{\Lambda}}\left(\Theta\right)\;\middle\|\;\mathring{\mathbf{\Lambda}}\left(\Gamma\right)\right]$ for all $\mathring{\mathbf{\Lambda}}\in\mathsf{SSChan}$. The proof goes as follows. For $\mathring{\mathbf{\Lambda}}\in\mathsf{SSChan}$, the superchannels $\mathring{\mathbf{\Lambda}}(\Theta)$ and $\mathring{\mathbf{\Lambda}}(\Gamma)$ can be written in terms of concatenation of superchannels given by Eq. (79). Using our definition of divergence between two superchannels, we can write
$\mathbf{D}^{(2)}\left[\mathring{\mathbf{\Lambda}}\left(\Theta\right)\;\middle\|\;\mathring{\mathbf{\Lambda}}\left(\Gamma\right)\right]=\sup_{\mathcal{N}\in\mathcal{L}(R^{\prime}\otimes A^{\prime},R^{\prime}\otimes B^{\prime})}\mathbf{D}\left[\left(\mathrm{id}_{R^{\prime}}\otimes\mathring{\mathbf{\Lambda}}\left(\Theta\right)\right)\left(\mathcal{N}\right)\;\middle\|\;\left(\mathrm{id}_{R^{\prime}}\otimes\mathring{\mathbf{\Lambda}}\left(\Gamma\right)\right)\left(\mathcal{N}\right)\right],$ (80)
where $R^{\prime}$ is the Hilbert space of a reference system. Using Eq. (79), we can write $\left(\operatorname{id}_{R^{\prime}}\otimes\mathring{\mathbf{\Lambda}}\left(\Theta\right)\right)\left(\mathcal{N}\right)=\left(\operatorname{id}_{R^{\prime}}\otimes\Lambda_{\text{post}}\right)\left(\operatorname{id}_{R^{\prime}}\otimes\left(\Theta\otimes\mathrm{id}_{R}\right)\right)\widetilde{\mathcal{N}}$, where $\widetilde{\mathcal{N}}:=\left(\operatorname{id}_{R^{\prime}}\otimes\Lambda_{\text{pre}}\right)(\mathcal{N})$. Similarly, we have $\left(\operatorname{id}_{R^{\prime}}\otimes\mathring{\mathbf{\Lambda}}\left(\Gamma\right)\right)\left(\mathcal{N}\right)=\left(\operatorname{id}_{R^{\prime}}\otimes\Lambda_{\text{post}}\right)\left(\operatorname{id}_{R^{\prime}}\otimes\left(\Gamma\otimes\mathrm{id}_{R}\right)\right)\widetilde{\mathcal{N}}$. Using these in the right-hand side of Eq. (80), we obtain
$\mathbf{D}^{(2)}\left[\mathring{\mathbf{\Lambda}}\left(\Theta\right)\;\middle\|\;\mathring{\mathbf{\Lambda}}\left(\Gamma\right)\right]$
$\leq\sup_{\widetilde{\mathcal{N}}\in\mathcal{L}(R^{\prime}\otimes A^{\prime}\otimes R,~{}R^{\prime}\otimes B^{\prime}\otimes R)}\mathbf{D}\left[\left(\operatorname{id}_{R^{\prime}}\otimes\Lambda_{\text{post}}\right)\left(\operatorname{id}_{R^{\prime}}\otimes\left(\Theta\otimes\mathrm{id}_{R}\right)\right)\left(\widetilde{\mathcal{N}}\right)\;\middle\|\;\left(\operatorname{id}_{R^{\prime}}\otimes\Lambda_{\text{post}}\right)\left(\operatorname{id}_{R^{\prime}}\otimes\left(\Gamma\otimes\mathrm{id}_{R}\right)\right)\left(\widetilde{\mathcal{N}}\right)\right]$
$\leq\sup_{\widetilde{\mathcal{N}}\in\mathcal{L}(R^{\prime}\otimes A^{\prime}\otimes R,~{}R^{\prime}\otimes B^{\prime}\otimes R)}\mathbf{D}\left[\left(\operatorname{id}_{R^{\prime}}\otimes\left(\Theta\otimes\mathrm{id}_{R}\right)\right)\left(\widetilde{\mathcal{N}}\right)\;\middle\|\;\left(\operatorname{id}_{R^{\prime}}\otimes\left(\Gamma\otimes\mathrm{id}_{R}\right)\right)\left(\widetilde{\mathcal{N}}\right)\right]=\mathbf{D}^{(2)}\left[\Theta\;\middle\|\;\Gamma\right].$
where the first inequality follows from the fact that the set of channels consisting $\widetilde{\mathcal{N}}$ is a larger set compared to the set consisting of $\mathcal{N}$. The second inequality follows by using the fact that $\left(\operatorname{id}_{R^{\prime}}\otimes\Lambda_{\text{post}}\right)$ is a valid quantum super-superchannel and the generalized channel divergence $\mathbf{D}\left[\cdot\|\cdot\right]$ is monotonic under quantum super-superchannels. Thus, we have $\mathbf{D}^{(2)}\left[\Theta\;\middle\|\;\Gamma\right]\geq\mathbf{D}^{(2)}\left[\mathring{\mathbf{\Lambda}}\left(\Theta\right)\;\middle\|\;\mathring{\mathbf{\Lambda}}\left(\Gamma\right)\right]$.
In the following, we define a notion of generalized entropy for quantum superchannels based on the notion of the generalized divergence. Recall that generalized divergences $\mathbf{D}\left(\cdot\;\middle\|\;\cdot\right)$ are functionals that take a pair of operators, say $A,B\geq 0$, as input and give a real number as output. One can define generalized entropies by taking $B$ to be the identity operator. For example, by taking $\mathbf{D}\left(\cdot\;\middle\|\;\cdot\right)$ as relative entropy, and $B=\mathbbm{1}$, we get von Neumann entropy, which we have already discussed in the previous section. Also, by taking $\mathbf{D}\left(\cdot\;\middle\|\;\cdot\right)$ as sandwiched Renyi divergence WWY (14); MLDS+ (13), and again $B=\mathbbm{1}$, one can get $\alpha$-Renyi entropy. To define the entropy of quantum channels, we started with the relative entropy which is a generalized divergence, and replaced states with channels and identity with the completely depolarising map and optimized overall input space of the channel. Similarly, we define the entropy of superchannels by taking the generalized divergence of channels as the starting point and replacing channels with superchannels and replacing the completely depolarising map with the supermap that serves its generalization in the space of supermaps, and then optimizing over all input space of the superchannel. Recall that the completely depolarizing map was defined as $\mathcal{R}\left(\mathsf{X}\right)=\operatorname{tr}(\mathsf{X})\mathbbm{1}$, where $\mathsf{X}$ was an arbitrary operator belonging to the input space. We now generalize this definition of the completely depolarizing map for the space of supermaps as follows. Let us denote by $\mathcal{R}^{(2)}$ a completely depolarising supermap.
Definition 9.
The action of a completely depolarizing supermap $\mathcal{R}^{(2)}:\mathcal{L}\left(A,A^{\prime}\right)\rightarrow\mathcal{L}\left(C,C^{\prime}\right)$ is defined as
$\mathcal{R}^{(2)}(\mathsf{X})=\operatorname{tr}\left(\widehat{\mathsf{C}}_{\mathsf{X}}\right)\mathcal{R},$ (81)
where $\mathsf{X}$ is an arbitrary map in the input space of $\mathcal{R}^{(2)}$, $\mathcal{R}$ is the completely depolarising map, and $\widehat{\mathsf{C}}_{\mathsf{X}}$ is the Choi operator for $\mathsf{X}$.
The map $\mathcal{R}^{(2)}$ is a completely CP-preserving supermap. With this definition, we are now in a position to define the entropy of higher-order processes.
Definition 10.
We define the generalized entropy of a superchannel $\Theta:\mathcal{L}\left(A,A^{\prime}\right)\rightarrow\mathcal{L}\left(C,C^{\prime}\right)$ as
$\mathbf{S}\left[\Theta\right]:=-\mathbf{D}^{(2)}\left[\Theta\;\middle\|\;\mathcal{R}^{(2)}\right],$ (82)
where $\mathbf{D}^{(2)}\left[\Theta\;\middle\|\;\mathcal{R}^{(2)}\right]$ is a generalized divergence between superchannel $\Theta$ and the completely depolarizing supermap $\mathcal{R}^{(2)}$. If we fix generalized divergence in Eq. (83) to be quantum relative entropy, then we get the entropy $S[\Theta]$ of the superchannel $\Theta$, i.e.,
${S}\left[\Theta\right]:=-D^{(2)}\left[\Theta\;\middle\|\;\mathcal{R}^{(2)}\right].$ (83)
In the remaining section, we work only with Eq. (83), i.e., the quantum relative entropy. We show that the entropy of superchannels, given by Eq. (83), satisfies the following properties: (1) $S\left[\Theta\right]$ is nondecreasing under $\mathcal{R}^{(2)}$-subpreserving super-superchannels with a mild assumption (see Proposition 10), (2) $S\left[\Theta\right]$ is subadditive under tensor product of superchannels (see Proposition 11), (3) For a replacer superchannel $\Theta_{0}$, which maps every channel to a fix channel $\mathcal{N}_{0}$, the entropy of $\Theta_{0}$ is upper bounded by the entropy of $\alpha\mathcal{N}_{0}$, where $A$ is the input space of $\mathcal{N}_{0}$ and $\alpha=1/|A|$ (see Proposition 12). We first show that the entropy of superchannels is nondecreasing under $\mathcal{R}^{(2)}$-subpreserving super-superchannels as the following proposition.
Proposition 10.
Let $\mathring{\mathbf{\Lambda}}:\mathscr{L}\left[\mathcal{L}\left(A,B\right),\mathcal{L}\left(C,D\right)\right]\rightarrow\mathscr{L}\left[\mathcal{L}\left(A^{\prime},B^{\prime}\right),\mathcal{L}\left(C^{\prime},D^{\prime}\right)\right]$ be a quantum super-superchannel mapping quantum superchannels to quantum superchannels such that it satisfies decomposability as in Eq. (79). Let $\mathcal{R}^{(2)}:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ and ~$\widetilde{\mathcal{R}}^{(2)}:\mathcal{L}\left(A^{\prime},B^{\prime}\right)\rightarrow\mathcal{L}\left(C^{\prime},D^{\prime}\right)$ be the generalizations of completely depolarising map to the space of superchannels such that $\mathring{\mathbf{\Lambda}}\left[\mathcal{R}^{(2)}\right]\leq\widetilde{\mathcal{R}}^{(2)}$ (i.e., ~$\widetilde{\mathcal{R}}^{(2)}-\mathring{\mathbf{\Lambda}}\left[\mathcal{R}^{(2)}\right]$ is completely CP-preserving). Then for any $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$, we have $S\left[\mathring{\mathbf{\Lambda}}\left[\Theta\right]\right]\geq S\left[\Theta\right]$.
Proof.
Since the super-superchannel $\mathring{\mathbf{\Lambda}}$ satisfies the decomposability condition given by Eq. (79), the data processing inequality holds for the pair $\Theta$ and $\mathcal{R}^{(2)}$ under the super-superchannel $\mathring{\mathbf{\Lambda}}$. Thus, we have
$\sup_{\mathcal{\mathcal{N}}\in\mathcal{L}\left(R^{\prime}\otimes A,R^{\prime}\otimes B\right)}D\left[\left(\mathrm{id}_{R^{\prime}}\otimes\Theta\right)\left(\mathcal{N}\right)\;\middle\|\;\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{R}^{(2)}\right)\left(\mathcal{N}\right)\right]$
$\geq\sup_{\widetilde{\mathcal{\mathcal{N}}}\in\mathcal{L}\left(R^{\prime}\otimes A^{\prime},R^{\prime}\otimes B^{\prime}\right)}D\left[\left(\mathrm{id}_{R^{\prime}}\otimes\mathring{\mathbf{\Lambda}}\left[\Theta\right]\right)\left(\widetilde{\mathcal{N}}\right)\;\middle\|\;\left(\mathrm{id}_{R^{\prime}}\otimes\mathring{\mathbf{\Lambda}}\left[\mathcal{R}^{(2)}\right]\right)\left(\widetilde{\mathcal{N}}\right)\right]$
$\geq\sup_{\widetilde{\mathcal{\mathcal{N}}}\in\mathcal{L}\left(R^{\prime}\otimes A^{\prime},R^{\prime}\otimes B^{\prime}\right)}D\left[\left(\mathrm{id}_{R^{\prime}}\otimes\mathring{\mathbf{\Lambda}}\left[\Theta\right]\right)\left(\widetilde{\mathcal{N}}\right)\;\middle\|\;\left(\mathrm{id}_{R^{\prime}}\otimes\widetilde{\mathcal{R}}^{(2)}\right)\left(\widetilde{\mathcal{N}}\right)\right],$
where we have used Proposition 3 in the second line (by identifying $\left(\mathrm{id}_{R^{\prime}}\otimes\mathring{\mathbf{\Lambda}}\left[\mathcal{R}^{(2)}\right]\right)\left(\widetilde{\mathcal{N}}\right)$ with $\mathcal{M}$ and $\left(\mathrm{id}_{R^{\prime}}\otimes\widetilde{\mathcal{R}}^{(2)}\right)\left(\widetilde{\mathcal{N}}\right)$ with $\widetilde{\mathcal{M}}$ so that $\mathcal{M}\leq\widetilde{\mathcal{M}}$ holds). The second line also uses the following property of supremum. Suppose that $f,g:\mathtt{A}\rightarrow\mathbb{R}$ and $f\leq g$, if $g$ is bounded from above then $\sup_{\mathtt{A}}f\leq\sup_{\mathtt{A}}g$. Now, using the definition of entropy of superchannels, we obtain $S\left[\mathring{\mathbf{\Lambda}}\left[\Theta\right]\right]\leq S\left[\Theta\right]$. ∎
Remark 6.
We have proved the Proposition 10 only for the class of super-superchannels that are decomposable as in Eq. (79). However, the proposition only requires the data processing inequality for the pair $\Theta$ and $\mathcal{R}^{(2)}$ under the super-superchannel $\mathring{\mathbf{\Lambda}}$. It will be interesting to show that the data processing inequality holds true for the pair $\Theta$ and $\mathcal{R}^{(2)}$ under all the super-superchannels $\mathring{\mathbf{\Lambda}}$ and with that we can remove the condition of decomposability from the Proposition 10. We believe that all super-superchannels are decomposable in the form of Eq. (79), however, a proof or a counter-proof is left open for now.
Proposition 11.
The entropy functional of quantum superchannels is subadditive under tensor product of superchannels, i.e,
$S\left[\Theta\otimes\Gamma\right]\leq S\left[\Theta\right]+S\left[\Gamma\right],$ (84)
where $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ and $\Gamma:\mathcal{L}\left(A^{\prime},B^{\prime}\right)\rightarrow\mathcal{L}\left(C^{\prime},D^{\prime}\right)$ are any two superchannels.
Proof.
Let $\mathcal{R}^{(2)}:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ and ~$\widetilde{\mathcal{R}}^{(2)}:\mathcal{L}\left(A^{\prime},B^{\prime}\right)\rightarrow\mathcal{L}\left(C^{\prime},D^{\prime}\right)$ be the generalizations of completely depolarising map to the space of superchannels (see Eq. (81)). From the definition of the generalized divergence of superchannels, we have
$D^{(2)}\left[\Theta\otimes\Gamma\;\middle\|\;\mathcal{R}^{(2)}\otimes\widetilde{\mathcal{R}}^{(2)}\right]=\sup_{\mathcal{N}\in\mathcal{L}(E\otimes A\otimes A^{\prime},E\otimes B\otimes B^{\prime})}D\left[\left(\mathrm{id}_{E}\otimes\Theta\otimes\Gamma\right)\left(\mathcal{N}\right)\;\middle\|\;\left(\mathrm{id}_{E}\otimes\mathcal{R}^{(2)}\otimes\widetilde{\mathcal{R}}^{(2)}\right)\left(\mathcal{N}\right)\right],$
where we can take the ancilla $E=R_{1}\otimes R_{2}$ to be isomorphic to $A_{1}\otimes A_{2}$ with the reference systems $R_{1}$ and $R_{2}$ being isomorphic to $A_{1}$ and $A_{2}$, respectively. Now, using the facts that $R_{1}\otimes R_{2}\otimes A_{1}\otimes A_{2}$ is isomorphic to $R_{1}\otimes A_{1}\otimes R_{2}\otimes A_{2}$, and supremum over a set is always greater than or equal to the supremum over its subsets, we obtain
$\sup_{\mathcal{N}\in\mathcal{L}(E\otimes A\otimes A^{\prime},E\otimes B\otimes B^{\prime})}D\left[\left(\mathrm{id}_{E}\otimes\Theta\otimes\Gamma\right)\left(\mathcal{N}\right)\;\middle\|\;\left(\mathrm{id}_{E}\otimes\mathcal{R}^{(2)}\otimes\widetilde{\mathcal{R}}^{(2)}\right)\left(\mathcal{N}\right)\right]$
$\geq\sup_{\begin{subarray}{c}\mathcal{N}_{1}\in\mathcal{L}(R_{1}\otimes A,R_{1}\otimes B)\\ \mathcal{N}_{2}\in\mathcal{L}(R_{2}\otimes A^{\prime},R_{2}\otimes B^{\prime})\end{subarray}}D\left[\left(\mathrm{id}_{R_{1}}\otimes\Theta\right)\left(\mathcal{N}_{1}\right)\otimes\left(\mathrm{id}_{R_{2}}\otimes\Gamma\right)\left(\mathcal{N}_{2}\right)\;\middle\|\;\left(\mathrm{id}_{R_{1}}\otimes\mathcal{R}^{(2)}\right)\left(\mathcal{N}_{1}\right)\otimes\left(\mathrm{id}_{R_{2}}\otimes\widetilde{\mathcal{R}}^{(2)}\right)\left(\mathcal{N}_{2}\right)\right]$
$\geq\sup_{\mathcal{N}_{1}\in\mathcal{L}(R_{1}\otimes A,R_{1}\otimes B)}D\left[\left(\mathrm{id}_{R_{1}}\otimes\Theta\right)\left(\mathcal{N}_{1}\right)\;\middle\|\;\left(\mathrm{id}_{R_{1}}\otimes\mathcal{R}^{(2)}\right)\left(\mathcal{N}_{1}\right)\right]$
$\hskip 14.22636pt+\sup_{\mathcal{N}_{2}\in\mathcal{L}(R_{2}\otimes A^{\prime},R_{2}\otimes B^{\prime})}D\left[\left(\mathrm{id}_{R_{2}}\otimes\Gamma\right)\left(\mathcal{N}_{2}\right)\;\middle\|\;\left(\mathrm{id}_{R_{2}}\otimes\widetilde{\mathcal{R}}^{(2)}\right)\left(\mathcal{N}_{2}\right)\right]$
$=D^{(2)}\left[\Theta\;\middle\|\;\mathcal{R}^{(2)}\right]+D^{(2)}\left[\Gamma\;\middle\|\;\widetilde{\mathcal{R}}^{(2)}\right],$
where the second inequality follows from Proposition 4. Thus, we have
$-D^{(2)}\left[\Theta\otimes\Gamma\;\middle\|\;\mathcal{R}^{(2)}\otimes\widetilde{\mathcal{R}}^{(2)}\right]\leq-D^{(2)}\left[\Theta\;\middle\|\;\mathcal{R}^{(2)}\right]-D^{(2)}\left[\Gamma\;\middle\|\;\widetilde{\mathcal{R}}^{(2)}\right],$
which implies that $S\left[\Theta\otimes\Gamma\right]\leq S\left[\Theta\right]+S\left[\Gamma\right]$, completing the proof of the proposition. ∎
Proposition 12.
Let $\Theta_{0}:\mathcal{L}\left(A,A^{\prime}\right)\rightarrow\mathcal{L}\left(C,C^{\prime}\right)$ be a replacer superchannel defined as $\Theta_{0}\left(\mathsf{X}\right):=\alpha\operatorname{tr}\left(\widehat{\mathsf{C}}_{\mathsf{X}}\right)\mathcal{N}_{0}$, where $\mathsf{X}$ is an arbitrary map in the input space of $\Theta_{0}$, $\mathcal{N}_{0}\in\mathcal{L}(C,C^{\prime})$ is some fixed channel, $\widehat{\mathsf{C}}_{\mathsf{X}}$ is the Choi operator for $\mathsf{X}$, and $\alpha=|A|^{-1}$. Then
$S\left[\Theta_{0}\right]\leq S\left[\alpha\mathcal{N}_{0}\right].$
Proof.
Note that $\left(\mathrm{id}_{R^{\prime}}\otimes\Theta_{0}\right)(\mathcal{N})$ and $\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{R}^{(2)}\right)(\mathcal{N})$ can be written as $\mathcal{M}\otimes\alpha\mathcal{N}_{0}$ and $\mathcal{M}\otimes\mathcal{R}$, respectively for some map $\mathcal{M}$. Using the definitions of the replacer superchannel $\Theta_{0}$ and supermap $\mathcal{R}^{(2)}$, we can write
$\sup_{\mathcal{\mathcal{N}}\in\mathcal{L}(R^{\prime}\otimes A,R^{\prime}\otimes B)}D\left[\left(\mathrm{id}_{R^{\prime}}\otimes\Theta_{0}\right)(\mathcal{N})\;\middle\|\;\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{R}^{(2)}\right)(\mathcal{N})\right]$ $=\sup_{\mathcal{\mathcal{N}}\in\mathcal{L}(R^{\prime}\otimes A,R^{\prime}\otimes B)}D\left[\mathcal{M}\otimes\alpha\mathcal{N}_{0}\;\middle\|\;\mathcal{M}\otimes\mathcal{R}\right]$
$\geq D\left[\alpha\mathcal{N}_{0}\;\middle\|\;\mathcal{R}\right],$ (85)
where the inequality follows from Proposition 4. This completes the proof of the proposition. ∎
VI Discussion
Entropic functionals such as the von Neumann entropy and the relative entropy have played a pivotal role in the development of quantum technologies as they are the widely used functionals to gauge the performance and efficacy of quantum information processing and computing protocols. With the recent advances in quantum information theory, there is a prevailing consensus that the quantum channels constitute the fundamental objects whose control and manipulation yield the most general quantum information processing and computing protocols. Therefore, it is of utmost importance to associate a notion of entropic functionals to the network of quantum channels.
In this work, we start by discussing the notions of entropies and the relative entropies of quantum channels and generalize them to include the higher order processes. Using the results we derive in this work on the entropy gain of a quantum state under positive maps, which generalize some of the existing results in the literature, we provide a compact formula for the change in entropy of a quantum channel belonging to a wide class of quantum channels under the action of superchannels that preserves this class. Tele-covariant channels belong to this wide class of channels and for them we compute the entropy change under tele-covariance preserving superchannels as an example of our results. Further, we define the notion of the sufficiency of superchannels and obtain the necessary and sufficient conditions for sufficiency in terms of entropic inequalities. We then go on to the define the notions of generalized divergence and entropy for quantum super-superchannels and analyze their properties.
Following the similar ideas that we use to define generalized divergence and entropy of quantum superchannels, we can also define the generalized divergence and entropy of higher order processes. Given a generalized divergence between two $n^{\text{th}}$ order processes, we define the generalized divergence between two $(n+1)^{\text{th}}$ order processes as follows. Let $\mathcal{V}^{(n)}$ denote the vector space of $n^{\text{th}}$ order maps. We define $(n+1)^{\text{th}}$ order processes as the $(n+1)^{\text{th}}$ order maps that are completely $n^{\text{th}}$ order process preserving. Here $n=0$ represents quantum states, $n=1$ represents quantum channels, $n=2$ represents quantum superchannels, and so on. We denote the set of $n^{\text{th}}$ order processes by $\mathcal{V}^{(n)}_{+}$. Clearly, we have, $\mathcal{V}^{(n)}_{+}\subseteq\mathcal{V}^{(n)}$.
Possible generalized divergence functional.— Given a generalized divergence $\mathbf{D}^{(n)}\left[\cdot\;\middle\|\;\cdot\right]$ of $n^{\text{th}}$ order process, let us define the functional $\mathbf{D}^{(n+1)}\left[\Theta^{(n+1)}\;\middle\|\;\Gamma^{(n+1)}\right]$ between two $(n+1)^{\mathrm{\text{th}}}$ order processes $\Theta^{(n+1)}$ and $\Gamma^{(n+1)}$, where $n\in\mathbb{Z}_{+}$, as follows:
$\mathbf{D}^{(n+1)}\left[\Theta^{(n+1)}\;\middle\|\;\Gamma^{(n+1)}\right]=\sup_{\Omega^{(n)}\in\left(\mathcal{W}^{(n)}\otimes\mathcal{V}^{(n)}\right)_{+}}\mathbf{D}^{(n)}\left[\left(\mathrm{id}\otimes\Theta^{(n+1)}\right)\left(\Omega^{(n)}\right)\;\middle\|\;\left(\mathrm{id}\otimes\Gamma^{(n+1)}\right)\left(\Omega^{(n)}\right)\right],$ (86)
where $\mathcal{W}^{(n)}$ denotes an ancillary space that contains all $n^{\text{th}}$ order processes of all possible finite (positive integers) dimensions and the supremum is taken with respect to the joint process $\left(\mathcal{W}^{(n)}\otimes\mathcal{V}^{(n)}\right)_{+}$. Here $\mathrm{id}$ is the identity of $(n+1)^{\text{th}}$ order process. Note that the generalized divergence for $n^{\text{th}}$ order process is a functional $\mathbf{D}^{(n)}:\mathcal{V}^{(n)}_{+}\times\mathcal{V}^{(n)}_{+}\rightarrow\mathbb{R}$, which satisfies the following monotonicity relation under $(n+1)^{\text{th}}$ order process:
$\mathbf{D}^{(n)}\left[\Theta^{(n)}\;\middle\|\;\Gamma^{(n)}\right]\geq\mathbf{D}^{(n)}\left[\Phi^{(n+1)}\left(\Theta^{(n)}\right)\;\middle\|\;\Phi^{(n+1)}\left(\Gamma^{(n)}\right)\right],$ (87)
for all $\Theta^{(n)},\Gamma^{(n)}\in\mathcal{V}^{(n)}_{+}$ and $\Phi^{(n+1)}\in\mathcal{V}^{(n+1)}_{+}$. We expect that the following dilation theorem similar to Eq. (13) will hold for any $n+1$ order process: $\Phi^{(n+1)}\left(\Theta^{(n)}\right)=\Gamma_{\text{post}}^{(n)}\circ\left(\Theta^{(n)}\otimes\operatorname{id}\right)\circ\Lambda^{(n)}_{\text{pre}}$, where $\Gamma^{(n)}$ and $\Lambda^{(n)}$ are two given $n^{\text{th}}$ order processes similar to the preprocessing and postprocessing quantum channels as in Eq. (13). It is clear that the right-hand side is also a $n^{\text{th}}$ order process since it is a concatenation of $n^{\text{th}}$ order processes but we leave the proof of converse statement open. In Appendix E, we show that for all $\Phi^{(n+1)}\in\mathcal{V}^{(n+1)}_{+}$ and $\Theta^{(n)},\Gamma^{(n)}\in\mathcal{V}^{(n)}_{+}$ such that the above dilation holds, the functional $\mathbf{D}^{(n+1)}\left[\cdot\;\middle\|\;\cdot\right]$ (Eq. (86)) becomes a valid generalized divergence.
Now, to define the generalized entropy, we need the generalization of completely depolarizing maps to higher order processes, which we do as follows. Let ${\mathcal{R}}^{(n)}$ denote the completely depolarising map of $n^{\text{th}}$ order process. Then, the action of $\mathcal{R}^{(n)}:\mathcal{V}^{(n-1)}\rightarrow\mathcal{V}^{(n-1)}$, where $\mathcal{V}^{(n-1)}$ denotes the space of all $(n-1)^{\text{th}}$ order processes, is defined as $\mathcal{R}^{(n)}(\mathsf{X})=\widetilde{f}\left(\mathsf{X}\right)\mathcal{R}^{(n-1)}$, where $\mathsf{X}$ is an arbitrary $(n-1)^{\text{th}}$ order process and $\mathcal{R}^{(n-1)}$ is a completely depolarising map of $(n-1)^{\text{th}}$ order process, and $\widetilde{f}(\mathsf{X})$ is a linear functional that one has to choose suitably such that the map $\mathcal{R}^{(n)}$ becomes a completely $(n-1)^{\text{th}}$ order process preserving map. With these notions, we define the entropy $\mathbf{S}^{(n)}\left[\Theta^{(n)}\right]$ of a higher-order process ${\Theta}^{(n)}:\mathcal{V}^{(n-1)}\rightarrow\mathcal{V}^{(n-1)}$ as
$\mathbf{S}^{(n)}\left[\Theta^{(n)}\right]:=-\mathbf{D}^{(n)}\left[\Theta^{(n)}\;\middle\|\;\mathcal{R}^{(n)}\right],$ (88)
where $\mathbf{D}^{(n)}\left[\Theta^{(n)}\;\middle\|\;\mathcal{R}^{(n)}\right]$ is a generalized divergence between $n^{\text{th}}$ order process $\Theta^{(n)}$ and the completely depolarizing map $\mathcal{R}^{(n)}$ of $n^{\text{th}}$ order process.
We leave it as an open question to study the properties of $\mathbf{S}^{(n)}\left[\Theta^{(n)}\right]$, and to investigate an operational meaning for the definition of the generalized divergence in the discrimination task of higher-order quantum processes. Another interesting direction would be to discuss the properties of conditional entropy, mutual information, and conditional mutual information of multipartite quantum channels that can be derived based on the definition of the entropy and relative entropy of quantum channels. We believe that our main results and meta results will find interesting applications in wide domains of quantum information theory, such as quantum error correction, quantum communication, quantum many-body systems, and other areas of quantum physics where entropy change and data processing inequalities play crucial role.
Acknowledgments
We are grateful to Maarten Wegewijs and Kaiyuan Ji for carefully reading our paper and pointing out an error to a previous version of Theorem 2. We thank Mark M Wilde for discussions. VP and Sohail thank the International Institute of Information Technology Hyderabad for the support and hospitality during their visits in the summer 2023 where part of this work was done. SD acknowledges support from the Science and Engineering Research Board, Department of Science and Technology (SERB-DST), Government of India under Grant No. SRG/2023/000217. US and SD thank the International Institute of Information Technology Hyderabad for the Faculty Seed Grant.
Appendix A Structure of the Appendix
In the appendices, we provide detailed proofs of the results (lemmas, propositions, and theorems) of the main text and offer some additional observations. The structure of the appendix is as follows. In Appendix B, we provide the proofs of statements (which are about some desirable properties of spaces of channels and superchannels) from Section II. In Appendix C, we have proof of statements from Section III. This section also contains several bounds on entropy gain under CP, CP unital, and positive maps (see Theorem 3, Lemma 10 and Proposition 13). In Appendix D, we provide the proof of Theorem 2, which is also one of our main results and it provides a refinement of the data processing inequality. In Appendix E, we provide the proof of monotonicity of $\mathbf{D}^{(n)}$.
Appendix B Proofs of statements in Section II
B.1 Proof of Lemma 1
Let $\{b^{A}_{j}\}_{j}$ denote an orthonormal basis for $\mathcal{L}\left({A}\right)$. Any operator $\mathsf{K}\in\mathcal{L}(A)$ can be expressed as a linear combination of the elements of $\{b^{A}_{j}\}_{j}$ as: $\mathsf{K}=\sum_{j}\left<b_{j}^{A},\mathsf{K}\right>b_{j}^{A}$. Using this equation, the inner product between linear maps $\mathcal{N}_{A\rightarrow B}$ and $\mathcal{M}_{A\rightarrow B}$ can be written as
[…]
$=\sum_{i}\left<\mathcal{N}\left(\mathsf{B}^{A}_{i}\right),\mathcal{M}(a^{A}_{i})\right>$
$=\sum_{i}\left<\mathcal{N}(a^{A}_{i}),\mathcal{M}(a^{A}_{i})\right>,$
where $\{a^{A}_{i}\}$ is some orthonormal basis in $\mathcal{L}\left({A}\right)$ and we have used the fact that $\mathsf{B}^{A}_{i}:=\sum_{j}\left<b^{A}_{j},a^{A}_{i}\right>b^{A}_{j}=a^{A}_{i}$ in the last equality. This concludes the proof.
B.2 Proof of Lemma 2
We will first show that "if part" holds. Note that $\sum_{ijkl}\mathcal{E}^{A\rightarrow B}_{ijkl}\otimes\mathcal{E}^{A\rightarrow B}_{ijkl}$ is a CP map. As we have assumed that the supermap $\Theta$ is completely CP-preserving, we have $\Lambda_{\Theta}=\left(\mathrm{id}\otimes\Theta\right)\sum_{ijkl}\mathcal{E}^{A\rightarrow B}_{ijkl}\otimes\mathcal{E}^{A\rightarrow B}_{ijkl}$ is CP.
We will now prove the nontrivial "only if" part. Given that $\Lambda_{\Theta}$ is CP, we need to prove that if $\Omega\in\mathcal{L}\left(E,E^{\prime}\right)\otimes\mathcal{L}(A,B)$ is CP (where $E$ and $E^{\prime}$ are reference systems), then $\left({\mathrm{id}}\otimes\Theta\right)\Omega$ is also CP. First, we prove the following lemma
Lemma 8.
A map $\Omega\in\mathcal{L}\left(E,E^{\prime}\right)\otimes\mathcal{L}(A,B)$ is CP if and only if it can be written as
$\Omega=\sum_{n}\sum_{\{\alpha,\beta,\gamma,\delta\}}\sum_{\{p,q,r,s\}}{\kappa}^{n}_{\alpha\beta\gamma\delta}\bar{{\kappa}}^{n}_{pqrs}\mathcal{E}^{{E}\rightarrow{E^{\prime}}}_{\gamma r\alpha p}\otimes\mathcal{E}^{A\rightarrow B}_{\delta s\beta q},$ (89)
where $\kappa^{n}_{\alpha\beta\gamma\delta}$ are complex numbers, $\bar{\kappa}^{n}_{pqrs}$ denotes the complex conjugate of $\kappa^{n}_{pqrs}$ (these complex numbers depend on the CP map $\Omega$), and $\left\{\mathcal{E}^{{E}\rightarrow{E^{\prime}}}_{\gamma r\alpha p}\right\}_{\gamma r\alpha p}$, $\left\{\mathcal{E}^{A\rightarrow B}_{\delta s\beta q}\right\}_{\delta s\beta q}$ are orthonormal basis for $\mathcal{L}\left(E,E^{\prime}\right)$ and $\mathcal{L}\left(A,B\right)$, respectively.
Proof.
Let $\left\{e^{Q}_{ij}\right\}_{ij}$ be a canonical basis for the space $\mathcal{L}\left({Q}\right)$, where $Q=\{A,B,E,E^{\prime}\}$. In the bra-ket notation we can write $e^{Q}_{ij}=\outerproduct{i_{Q}}{j_{Q}}$ |, where {|$\left\{\ket{i_{Q}}\right\}_{i}$ is an orthonormal basis for $Q$. A map $\Omega$ is CP if and only if it has a Choi-Kraus decomposition as follows
$\Omega\left(\mathsf{X}\right)=\sum_{n}K_{n}\mathsf{X}K^{\dagger}_{n},$ (90)
where $K_{n}:E\otimes A\rightarrow E^{\prime}\otimes B$ are linear operators, and $\mathsf{X}\in\mathcal{L}(E\otimes A)$. So, we can write $K_{n}$ in terms of basis vectors of $A,B,E$, and $E^{\prime}$ as follows
[…]
$=\sum_{n}\sum_{\{\alpha\beta\gamma\delta\}}\sum_{\{pqrs\}}\kappa^{n}_{\alpha\beta\gamma\delta}\bar{\kappa}^{n}_{pqrs}\left<e^{E}_{\gamma r}\otimes e^{A}_{\delta s},\mathsf{X}\right>e^{E^{\prime}}_{\alpha p}\otimes e^{B}_{\beta q}$
$=\sum_{n}\sum_{\{\alpha\beta\gamma\delta\}}\sum_{\{pqrs\}}\sum_{\{ijkl\}}\kappa^{n}_{\alpha\beta\gamma\delta}\bar{\kappa}^{n}_{pqrs}x_{ijkl}\left<e^{E}_{\gamma r}\otimes e^{A}_{\delta s},e^{E}_{ij}\otimes e^{A}_{kl}\right>e^{{E^{\prime}}}_{\alpha p}\otimes e^{B}_{\beta q}$
[…]
$=\sum_{n}\sum_{\{\alpha\beta\gamma\delta\}}\sum_{\{pqrs\}}\sum_{\{ijkl\}}\kappa^{n}_{\alpha\beta\gamma\delta}\bar{\kappa}^{n}_{pqrs}x_{ijkl}\mathcal{E}^{E\rightarrow{E^{\prime}}}_{\gamma r\alpha p}\left(e^{E}_{ij}\right)\otimes\mathcal{E}^{{A}\rightarrow B}_{\delta s\beta q}\left(e^{A}_{kl}\right)$
$=\sum_{n}\sum_{\{\alpha\beta\gamma\delta\}}\sum_{\{pqrs\}}\kappa^{n}_{\alpha\beta\gamma\delta}\bar{\kappa}^{n}_{pqrs}\mathcal{E}^{E\rightarrow{E^{\prime}}}_{\gamma r\alpha p}\otimes\mathcal{E}^{A\rightarrow B}_{\delta s\beta q}\left(\sum_{\{ijkl\}}x_{ijkl}e^{E}_{ij}\otimes e^{A}_{kl}\right)$
$=\sum_{n}\sum_{\{\alpha\beta\gamma\delta\}}\sum_{\{pqrs\}}\kappa^{n}_{\alpha\beta\gamma\delta}\bar{\kappa}^{n}_{pqrs}\mathcal{E}^{E\rightarrow{E^{\prime}}}_{\gamma r\alpha p}\otimes\mathcal{E}^{A\rightarrow B}_{\delta s\beta q}\left(\mathsf{X}\right).$
The second line simply follows from the definition of the Hilbert-Schmidt inner product and we have written back everything in a canonical basis. In the third line, we have used a decomposition of $\mathsf{X}=\sum_{\{ijkl\}}x_{ijkl}e^{E}_{ij}\otimes e^{A}_{kl}$ to rewrite $\mathsf{X}$ in terms of canonical basis, and $\{\mathcal{E}^{E\rightarrow{E^{\prime}}}_{\gamma r\alpha p}\}_{\gamma r\alpha p}$, $\{\mathcal{E}^{{A}\rightarrow B}_{\delta s\beta q}\}_{\delta s\beta q}$ are basis for spaces $\mathcal{L}\left(E,E^{\prime}\right)$ and $\mathcal{L}\left(A,B\right)$, respectively. The fourth and fifth line follows from the definition of basis in $\{\mathcal{E}^{E\rightarrow{E^{\prime}}}_{\gamma r\alpha p}\}$ and $\{\mathcal{E}^{{A}\rightarrow B}_{\delta s\beta q}\}$ (see Eq.(17)). This concludes the proof of the lemma.
Now given Eq (89) we define a supermap ${\Xi}_{\mathsf{X}}:\mathcal{L}(A,B)\rightarrow\mathcal{L}\left(E,E^{\prime}\right)$ by linear extension which is given by the following equation
$\Xi_{\mathsf{X}}\left(\mathcal{E}^{{A}\rightarrow{B}}_{\delta s\beta q}\right)=\sum_{n}\sum_{\{\alpha\gamma rp\}}\kappa^{n}_{\alpha\beta\gamma\delta}\bar{\kappa}^{n}_{pqrs}\mathcal{E}^{{E}\rightarrow{E^{\prime}}}_{\gamma r\alpha p},$ (92)
where $\kappa^{n}_{\alpha\beta\gamma\delta}$ are complex numbers defined in Eq. (91) and $\bar{\kappa}^{n}_{\alpha\beta\gamma\delta}$ is the complex conjugate of $\kappa^{n}_{\alpha\beta\gamma\delta}$. So, we can write $\Omega$ as follows
$\Omega=\left(\Xi_{\mathsf{X}}\otimes\mathrm{id}\right)\left(\sum_{\{\delta,s,\beta,q\}}\mathcal{E}^{{A}\rightarrow B}_{\delta s\beta q}\otimes\mathcal{E}^{{A}\rightarrow B}_{\delta s\beta q}\right).$ (93)
Applying $\left(\mathrm{id}\otimes\Theta\right)$ on both sides of the above equation, we obtain
$\left(\mathrm{id}\otimes\Theta\right)\Omega$ $=\left(\Xi_{\mathsf{X}}\otimes\mathrm{id}\right)\left(\sum_{\{\delta,s,\beta,q\}}\mathcal{E}^{{A}\rightarrow B}_{\delta s\beta q}\otimes\Theta\left(\mathcal{E}^{{A}\rightarrow B}_{\delta s\beta q}\right)\right)=(\Xi_{\mathsf{X}}\otimes\mathrm{id})\Lambda_{\Theta}.$ (94)
It was shown in Lemma 8 that a map $\Lambda_{\Theta}\in\mathcal{L}(A,B)\otimes\mathcal{L}(C,D)$ is CP if and only if it can be written as
$\Lambda_{\Theta}=\sum_{m}\sum_{\{ijkl\}}\sum_{\{abcd\}}b^{m}_{ijkl}\bar{b}^{m}_{abcd}\mathcal{E}^{A\rightarrow B}_{kcia}\otimes\mathcal{E}^{C\rightarrow D}_{ldjb}.$ (95)
The map $\Lambda_{\Theta}$ in Eq. (94) is CP, thus, the use of the above form for $\Lambda_{\Theta}$ in Eq. (94) leads to
$(\mathrm{id}\otimes\Theta)\Omega$ $=\sum_{m}\sum_{\{ijkl\}}\sum_{\{abcd\}}b^{m}_{ijkl}\bar{b}^{m}_{abcd}~{}\Xi_{\mathsf{X}}\left(\mathcal{E}^{A\rightarrow B}_{kcia}\right)\otimes\mathcal{E}^{C\rightarrow D}_{ldjb}$
$=\sum_{n}\sum_{m}\sum_{\{ijkl\}}\sum_{\{abcd\}}\sum_{\{tuvw\}}b^{m}_{ijkl}\bar{b}^{m}_{abcd}{\kappa}^{n}_{vitk}\bar{\kappa}^{n}_{wauc}\mathcal{E}^{{E}\rightarrow{E^{\prime}}}_{tuvw}\otimes\mathcal{E}^{{C}\rightarrow{D}}_{ldjb}$
$=\sum_{nm}\sum_{\{jlbd\}}\sum_{\{tuvw\}}\left(\sum_{ik}{\kappa}^{n}_{vitk}{b}^{m}_{ijkl}\right)\left(\sum_{ac}\bar{{\kappa}}^{n}_{wauc}{\bar{b}}^{m}_{abcd}\right)\mathcal{E}^{{E}\rightarrow{E^{\prime}}}_{tuvw}\otimes\mathcal{E}^{{C}\rightarrow{D}}_{ldjb},$ (96)
where the second line simply follows from the definition of $\Xi_{\mathsf{X}}$ (see Eq. (92)). In terms of canonical basis, the coefficients $\sum_{ik}{\kappa}^{n}_{vitk}{b}^{m}_{ijkl}$ and $\sum_{ac}\bar{\kappa}^{n}_{wauc}\bar{b}^{m}_{abcd}$ have the following form
$\sum_{ik}{\kappa}^{n}_{vitk}{b}^{m}_{ijkl}$ $=\mathrm{tr}\left[\left(\outerproduct{j_{{D}}}{v_{{E^{\prime}}}}\otimes\mathbbm{1}_{{B}}\right)K_{n}\left(\outerproduct{t_{{E}}}{l_{{C}}}\otimes\mathbbm{1}_{A}\right)B_{m}^{t}\right]$
$\sum_{ac}\bar{\kappa}^{n}_{wauc}\bar{b}^{m}_{abcd}$ $=\mathrm{tr}\left[\left(\outerproduct{b_{{D}}}{w_{{E^{\prime}}}}\otimes\mathbbm{1}_{B}\right)\bar{K}_{n}\left(\outerproduct{u_{{A}}}{d_{{C}}}\otimes\mathbbm{1}_{A}\right)B_{m}^{\dagger}\right],$
where $B_{m}:{A}\otimes C\rightarrow{B}\otimes{D}$ are linear operators, and $B_{m}^{t}$ is transpose of $B_{m}$. Now we define operators $F^{mn}:E\otimes{C}\rightarrow{E^{\prime}}\otimes{D}$ such that its matrix elements are given by the following
$F^{mn}_{vjtl}=\mathrm{tr}\left[\left(\outerproduct{j_{{D}}}{v_{{E^{\prime}}}}\otimes\mathbbm{1}_{{B}}\right)K_{n}\left(\outerproduct{t_{{E}}}{l_{{C}}}\otimes\mathbbm{1}_{A}\right){B_{m}}^{t}\right].$
Note that the complex conjugate of the matrix elements of $F^{mn}$ are given by
$\bar{F}^{mn}_{wbud}$ $=\mathrm{tr}\left[\left(\outerproduct{b_{{D}}}{w_{{E^{\prime}}}}\otimes\mathbbm{1}_{B}\right)\bar{K}_{n}\left(\outerproduct{u_{{A}}}{d_{{C}}}\otimes\mathbbm{1}_{A}\right)B_{m}^{\dagger}\right].$
In terms of these matrix elements, Eq. (96) can be rewritten as the following
$(\mathrm{id}\otimes\Theta)\Omega$ $=\sum_{n,m}\sum_{\{jlbd\}}\sum_{\{tuvw\}}F^{mn}_{vjtl}\bar{F}^{mn}_{wbud}~{}\mathcal{E}^{{E}\rightarrow{E^{\prime}}}_{tuvw}\otimes\mathcal{E}^{{C}\rightarrow{D}}_{ldjb}.$
Looking at the above form it can be concluded that $(\mathrm{id}\otimes\Theta)\Omega$ is CP. This completes the proof of the theorem.
B.3 Proof of Lemma 3
We already have shown a supermap $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ is completely CP-preserving if and only if $\Lambda_{\Theta}=\sum_{ijkl}\mathcal{E}^{A\rightarrow B}_{ijkl}\otimes\Theta\left(\mathcal{E}^{A\rightarrow B}_{ijkl}\right)\in\mathcal{L}\left(A,B\right)\otimes\mathcal{L}\left(C,D\right)$ is CP, where $\left\{\mathcal{E}^{{A}\rightarrow B}_{\delta s\beta q}\right\}_{\delta s\beta q}$ is a basis for $\mathcal{L}\left(A,B\right)$. We can determine the complete positivity of $\Lambda_{\Theta}$ by checking the positivity of its Choi operator $\widehat{\mathsf{C}}_{\Lambda_{\Theta}}$ as defined below
$\widehat{\mathsf{C}}_{\Lambda_{\Theta}}=\sum_{pqrs}\left(e^{{A}}_{pq}\otimes e^{C}_{rs}\right)\otimes\Lambda_{\Theta}\left(e^{A}_{pq}\otimes e^{C}_{rs}\right)\in\mathcal{L}(A\otimes C\otimes B\otimes D),$ (97)
where $\{e^{Q}_{ij}\}_{ij}$ denotes a canonical basis for the space $\mathcal{L}\left(Q\right)$ and $Q=\{A,B,C,D\}$. Using the form of $\Lambda_{\Theta}$ from Eq. (18) in the above equation, we have
$\widehat{\mathsf{C}}_{\Lambda_{\Theta}}=\sum_{rs}\sum_{ijkl}\left(e^{A}_{ij}\otimes e^{C}_{rs}\right)\otimes\left(e^{B}_{kl}\otimes\Theta(\mathcal{E}^{A\rightarrow B}_{ijkl})e^{C}_{rs}\right).$ (98)
Now, the Choi operator of the representing map $\mathsf{T}$ of $\Theta$ is the following
$\widehat{\mathsf{C}}_{\mathsf{T}}$ $=\sum_{ijkl}\left(e^{A}_{ij}\otimes e^{B}_{kl}\right)\otimes\mathsf{T}\left(e^{A}_{ij}\otimes e^{B}_{kl}\right)=\sum_{ijkl}\widehat{\mathsf{C}}_{\mathcal{E}^{A\rightarrow B}_{ijkl}}\otimes\mathsf{T}\left(\widehat{\mathsf{C}}_{\mathcal{E}^{A\rightarrow B}_{ijkl}}\right)=\sum_{ijkl}\widehat{\mathsf{C}}_{\mathcal{E}^{A\rightarrow B}_{ijkl}}\otimes\widehat{\mathsf{C}}_{\Theta\left({\mathcal{E}^{A\rightarrow B}_{ijkl}}\right)},$
where we have used the fact that $e^{A}_{ij}\otimes e^{{B}}_{kl}$ is the Choi operator $\mathsf{C}_{\mathcal{E}^{{A}\rightarrow{B}}_{ijkl}}$ for $\mathcal{E}^{{A}\rightarrow{B}}_{ijkl}$ in the second equality, and the last equality directly follows from Eq. (19). Interchanging the spaces $B$ and $C$ in the Eq. (98), we have
$\widehat{\mathsf{C}}_{\Lambda_{\Theta}}$ $:=\sum_{rs}\sum_{ijkl}\left(e^{A}_{ij}\otimes e^{B}_{kl}\right)\otimes\left(e^{C}_{rs}\otimes\Theta(\mathcal{E}^{A\rightarrow B}_{ijkl})e^{C}_{rs}\right)$
$=\sum_{ijkl}\left(e^{A}_{ij}\otimes e^{B}_{kl}\right)\otimes\sum_{rs}\left(e^{C}_{rs}\otimes\Theta(\mathcal{E}^{A\rightarrow B}_{ijkl})e^{C}_{rs}\right)=\sum_{ijkl}\widehat{\mathsf{C}}_{\mathcal{E}^{A\rightarrow B}_{ijkl}}\otimes\widehat{\mathsf{C}}_{\Theta\left({\mathcal{E}^{A\rightarrow B}_{ijkl}}\right)}=\widehat{\mathsf{C}}_{\mathsf{T}}.$
Now, the transformation $\widehat{\mathsf{C}}_{\Lambda_{\Theta}}\leftrightarrow\widehat{\mathsf{C}}_{\mathsf{T}}$ is an -isomorphism. So, $\widehat{\mathsf{C}}_{\mathsf{T}}$ is positive semidefinite if and only if $\widehat{\mathsf{C}}_{\Lambda_{\Theta}}$ is positive semidefinite. Now, $\widehat{\mathsf{C}}_{\Lambda_{\Theta}}$ is positive semidefinite if and only if $\Lambda_{\Theta}$ is CP, and $\Lambda_{\Theta}$ is CP if and only if $\Theta$ is completely CP-preserving. As $\mathsf{T}$ is CP if and only if $\widehat{\mathsf{C}}_{\mathsf{T}}$ is positive semidefinite, we have $\mathsf{T}$ is CP if and only if $\Theta$ is completely CP-preserving. This completes the proof.
B.4 Quantum channels from representing maps
Let us define a map $\mathsf{T}^{\prime}:\mathcal{L}\left(A\otimes{B}\right)\rightarrow\mathcal{L}\left({C}\otimes{D}\right)$ as follows
$\mathsf{T}{{}^{\prime}}(\mathsf{X})=\mathsf{T}(\mathsf{X})+\mathsf{T}_{0}(\mathsf{X}),$ (99)
where $\mathsf{T}_{0}(\mathsf{X}):=\left[\mathrm{tr}(\mathsf{X})-\mathrm{tr}(\mathsf{T}(\mathsf{X}))\right]\sigma_{0}$, $\sigma_{0}\in\mathcal{L}(C\otimes D)$ and $\mathsf{T}:\mathcal{L}\left(A\otimes{B}\right)\rightarrow\mathcal{L}\left({C}\otimes{D}\right)$ is the representing map of the superchannel $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$. Note that the representing map $\mathsf{T}$ is CP but it could be trace non-increasing or trace increasing or neither trace increasing nor trace decreasing. However, if we take $\sigma_{0}$ to be an operator with unit trace then $\mathsf{T}^{\prime}$ becomes a trace preserving map but then $\mathsf{T}^{\prime}$ may not be a CP map. In the following, we wish to find out if $\mathsf{T}^{\prime}$ is a CP map given that $\mathsf{T}$ is CP and $\operatorname{tr}(\sigma_{0})=1$. In the case where $\mathsf{T}$ is trace non-increasing, the condition $\sigma_{0}\geq 0$ in Eq. (99) guarantees that the map $\mathsf{T}^{\prime}$ is CP. In other cases, we derive the necessary and sufficient conditions for $\mathsf{T}^{\prime}$ to be CP as follows.
Necessary and Sufficient Conditions.—The Choi operator of the map $\mathsf{T}_{0}$ in Eq. (99) is given by
$\widehat{\mathsf{C}}_{\mathsf{T}_{0}}=\sum_{ij}e_{ij}\otimes\mathsf{T}_{0}(e_{ij})=\sum_{ij}e_{ij}\otimes\left[\delta_{ij}-\mathrm{tr}(\mathsf{T}(e_{ij}))\right]\sigma_{0}=\left[\mathbbm{1}_{AB}-\mathsf{T}^{*}(\mathbbm{1}_{CD})\right]^{t}\otimes\sigma_{0},$ (100)
where $t$ denotes the transpose operation in the basis $\{e_{ij}\}$. Then Eq. (99) in terms of Choi operators implies
$\widehat{\mathsf{C}}_{\mathsf{T}^{\prime}}=\widehat{\mathsf{C}}_{\mathsf{T}}+\left[\mathbbm{1}_{AB}-\mathsf{T}^{*}(\mathbbm{1}_{CD})\right]^{t}\otimes\sigma_{0},$ (101)
where $\widehat{\mathsf{C}}_{\mathsf{T}^{\prime}}$ and $\widehat{\mathsf{C}}_{\mathsf{T}}$ are the Choi operators for the maps $\mathsf{T}^{\prime}$ and $\mathsf{T}$, respectively. We have freedom over $\sigma_{0}$ in the above equation and we need to choose it such that $\widehat{\mathsf{C}}_{\mathsf{T}^{\prime}}$ becomes positive semidefinite. Note that $\left[\mathbbm{1}_{AB}-\mathsf{T}^{*}(\mathbbm{1}_{CD})\right]^{t}$ is a hermitian operator and let its spectral decomposition be given by $\left[\mathbbm{1}_{AB}-\mathsf{T}^{*}(\mathbbm{1}_{CD})\right]^{t}=\sum_{m:\lambda_{m}>0}\lambda_{m}\Pi_{m}-\sum_{m:\lambda_{m}<0}|\lambda_{m}|\Pi_{m}+\sum_{m:\lambda_{m}=0}\lambda_{m}\Pi_{m}$, where $\{\Pi_{m}\}_{m:\lambda_{m}>0},~{}\{\Pi_{m}\}_{m:\lambda_{m}<0}$ and $\{\Pi_{m}\}_{m:\lambda_{m}=0}$ are orthogonal projectors. Further, define $\Pi_{+}:=\sum_{m:\lambda_{m}>0}\Pi_{m}$ and $\Pi_{-}:=\sum_{m:\lambda_{m}<0}|\lambda_{m}|\Pi_{m}$ as the projectors onto the support of positive and negative eigenvalues of the hermitian operator $\left[\mathbbm{1}_{AB}-\mathsf{T}^{*}(\mathbbm{1}_{CD})\right]^{t}$, respectively. We denote by $\Pi_{0}$ the projector onto the eigenspace corresponding to the zero eigenvalue of $\left[\mathbbm{1}_{AB}-\mathsf{T}^{*}(\mathbbm{1}_{CD})\right]^{t}$ One can now prove easily that $\widehat{\mathsf{C}}_{\mathsf{T}^{\prime}}$ is positive semidefinite if and only if the operator $\sigma_{0}\in\mathcal{L}(C\otimes D)$ satisfies
$\left(\Pi_{+}-\Pi_{-}\right)\otimes\sigma_{0}\leq\left(\left[|(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}-\mathbbm{1}_{AB}\right|^{-\frac{1}{2}}+\Pi_{0}]\otimes\mathbbm{1}_{CD}\right)\widehat{\mathsf{C}}_{\mathsf{T}}\left([\left|(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}-\mathbbm{1}_{AB}\right|^{-\frac{1}{2}}+\Pi_{0}]\otimes\mathbbm{1}_{CD}\right).$ (102)
Now we will show that this condition is also compatible with the trace condition on $\sigma_{0}$ ($\operatorname{tr}\sigma_{0}=1$). By taking the trace with respect to the second subsystem in the above equation, we obtain
$\left(\Pi_{+}-\Pi_{-}\right)(\operatorname{tr}{\sigma_{0}})\leq\left(\left|(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}-\mathbbm{1}_{AB}\right|^{-\frac{1}{2}}+\Pi_{0}\right)(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}\left(\left|(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}-\mathbbm{1}_{AB}\right|^{-\frac{1}{2}}+\Pi_{0}\right).$ (103)
Let $\{\mu_{k}\}$ be the spectrum of $(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}$, then the spectral decomposition of $(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}$ gives $(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}=\sum_{i}\mu_{i}P_{i}$, where $P_{i}$’s are the projectors onto the eigensubspaces of $(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}$. Then we have $\left|(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}-\mathbbm{1}_{AB}\right|^{-\frac{1}{2}}=\sum_{i}\left|\mu_{i}-1\right|^{-\frac{1}{2}}P_{i}$, which implies $\left(\left[(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}-\mathbbm{1}_{AB}\right]^{-\frac{1}{2}}+\Pi_{0}\right)(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}\left(\left[(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}-\mathbbm{1}_{AB}\right]^{-\frac{1}{2}}+\Pi_{0}\right)=\sum_{i}\frac{\mu_{i}}{\left|\mu_{i}-1\right|}P_{i}+\Pi_{0}$. Note that $\mu_{i}$ is positive but can be less than one while the operator $\left(\Pi_{+}-\Pi_{-}\right)$ may be nonpositive, then there is no clash of trace condition on $\sigma_{0}$, i.e., $\operatorname{tr}(\sigma_{0})=1$, with inequality in Eq. (103).
In the special case where $\mathsf{T}$ is a trace increasing map, i.e., $\operatorname{tr}(\mathsf{X})<\mathrm{tr}(\mathsf{T}(\mathsf{X}))$ for all $X\in\mathcal{L}\left(A\otimes{B}\right)$ or $\mathsf{T}^{*}(\mathbbm{1}_{CD})-\mathbbm{1}_{AB}$ is a strictly positive operator. Under this condition we have $\Pi_{+}=\mathbbm{I}_{AB}$ and $\Pi_{-}=0$ and $\widehat{\mathsf{C}}_{\mathsf{T}^{\prime}}$ is positive semidefinite if and only if the operator $\sigma_{0}\in\mathcal{L}(C\otimes D)$ satisfies
$\mathbbm{1}_{AB}\otimes\sigma_{0}\leq\left(\left[(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}-\mathbbm{1}_{AB}\right]^{-\frac{1}{2}}\otimes\mathbbm{1}_{CD}\right)\widehat{\mathsf{C}}_{\mathsf{T}}\left(\left[(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}-\mathbbm{1}_{AB}\right]^{-\frac{1}{2}}\otimes\mathbbm{1}_{CD}\right).$ (104)
Taking partial trace with respect to the subsystems $C$ and $D$ in above equation, we get
$\mathbb{I}_{AB}\operatorname{tr}(\sigma_{0})\leq\left(\left[(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}-\mathbbm{1}_{AB}\right]^{-\frac{1}{2}}\right)(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}\left(\left[(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}-\mathbbm{1}_{AB}\right]^{-\frac{1}{2}}\right)=\sum_{i}\frac{\mu_{i}}{\mu_{i}-1}P_{i}>\mathbbm{I}_{AB}.$ (105)
Thus the trace condition on $\sigma_{0}$, i.e., $\operatorname{tr}(\sigma_{0})=1$ does not create any contradiction between Eq. (104).
We will now provide two examples on how to construct $\sigma_{0}$ for a given trace increasing map $\mathsf{T}$ another map $\mathsf{M}$ that is neither trace increasing nor trace decreasing.
Example $1$:–Let us define a map $\mathsf{T}:\mathcal{L}\left(A\otimes{B}\right)\rightarrow\mathcal{L}\left({C}\otimes{D}\right)$ as follows
$\mathsf{T}(\mathsf{X})=\alpha\operatorname{tr}(\mathsf{X})\mathbbm{1}_{CD},$ (106)
where $\alpha$ is real. We are not assuming whether $\mathsf{T}$ is trace increasing or decreasing yet. The above equation also implies $\mathsf{T}^{*}(\mathbbm{1}_{CD})=\alpha n\mathbbm{1}_{AB}$, where $n:=\operatorname{tr}(\mathbbm{1}_{CD})$ and $\mathsf{T}^{*}$ is adjoint of $\mathsf{T}$. We now have
$\left(\left[(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}-\mathbbm{1}_{AB}\right]^{-\frac{1}{2}}\otimes\mathbbm{1}_{CD}\right)\widehat{\mathsf{C}}_{\mathsf{T}}\left(\left[(\mathsf{T}^{*}(\mathbbm{1}_{CD}))^{t}-\mathbbm{1}_{AB}\right]^{-\frac{1}{2}}\otimes\mathbbm{1}_{CD}\right)=\left(\frac{\alpha}{n\alpha-1}\right)\mathbbm{1}_{AB}\otimes\mathbbm{1}_{CD}.$ (107)
Notice that if we choose $\alpha>\frac{1}{n}$, then $\mathsf{T}$ becomes trace increasing map. Using Eq. (104) we obtain the following inequality
$\mathbbm{1}_{AB}\otimes\sigma_{0}\leq\left(\frac{\alpha}{n\alpha-1}\right)\mathbbm{1}_{AB}\otimes\mathbbm{1}_{CD}.$ (108)
By a suitable choice of $\alpha$, we can make the number $\left(\frac{\alpha}{n\alpha-1}\right)$ as large as possible: take $\alpha=\frac{1}{n}+\frac{1}{n^{2}L}$, where $L$ is a very large positive real number, then $\left(\frac{\alpha}{n\alpha-1}\right)=L+\frac{1}{n}$. If we choose $\sigma_{0}$ to a unit trace and diagonal matrix with the diagonal elements being less than $L+\frac{1}{n}$ (some of the diagonal entries can be chosen to be negative also), then the above inequality will hold. Thus with this choice of $\sigma_{0}$, the CP and trace increasing map $\mathsf{T}(\mathsf{X})=\alpha\operatorname{tr}(\mathsf{X})\mathbbm{1}_{CD}$ can be made a CP and trace preserving map, using Eq. (99).
Example $2$:– Let us define a map $\mathsf{T}$, whose input and output systems are qubits, by the Kraus operators $K_{1}=\alpha\outerproduct{0}{0}$ | and $K_{2}=\beta\outerproduct{1}{1}$ |, where {|$\{\ket{0},\ket{1}\}$ are computational basis and $\alpha>1$, $10<\beta<$ 1 are positive real constants. If we take the input of $\mathsf{T}$ as $\mathsf{X}=\outerproduct{0}{0}$ |, we then have $\operatorname{tr}(\mathsf{X})<\mathrm{tr}(\mathsf{T}(\mathsf{X}))$. On the other hand if we take the input of $\mathsf{T}$ as $\mathsf{X}=\outerproduct{1}{1}$ |, we then have $\operatorname{tr}(\mathsf{X})>\mathrm{tr}(\mathsf{T}(\mathsf{X}))$. Clearly the map $\mathsf{T}$ is neither trace increasing nor trace decreasing map. The Choi operator of $\mathsf{T}$ is given by
$\widehat{\mathsf{C}}_{\mathsf{T}}=\alpha^{2}\outerproduct{00}{00}+\beta^{2}\outerproduct{11}{11}.$ | . (109)
Now we will calculate the right hand side of Eq. (102). First we need the following term
[…]
We then have
$\mathsf{A}:=\left(\left|(\mathsf{T}^{*}(\mathbbm{1}))^{t}-\mathbbm{1}\right|^{-\frac{1}{2}}\otimes\mathbbm{1}\right)\widehat{\mathsf{C}}_{\mathsf{T}}\left(\left|(\mathsf{T}^{*}(\mathbbm{1}))^{t}-\mathbbm{1}\right|^{-\frac{1}{2}}\otimes\mathbbm{1}\right)=\alpha^{2}(\alpha^{2}-1)^{-1}\outerproduct{00}{00}+\beta^{2}(\beta^{2}-1)^{-1}\outerproduct{11}{11}.$ | . (111)
We can write $\mathsf{T}^{*}(\mathbbm{1})^{t}-\mathbbm{1}=M-N$, where $M=(\alpha^{2}-1)\outerproduct{0}{0}$ | and $N=(1-\beta^{2})\outerproduct{1}{1}$ |. This also implies $\Pi_{M}=\outerproduct{0}{0}$ | and $\Pi_{N}=\outerproduct{1}{1}$ |. We want to find a $\sigma_{0}$ such that $\operatorname{tr}(\sigma_{0})=1$ and $(\Pi_{M}-\Pi_{N})\otimes\sigma_{0}\leq\mathsf{A}$. Let us choose a diagonal matrix $\sigma$ with eigenvalues $\lambda_{1}$ and $\lambda_{2}$ such that $\lambda_{1}<\alpha^{2}(\alpha^{2}-1)^{-1}$, $\left|\lambda_{2}\right|<\beta^{2}(\beta^{2}-1)^{-1}$, $\lambda_{2}<0$ and $\lambda_{1}+\lambda_{2}>1$. Now if we take $\alpha=\sqrt{2}$, $\beta=\frac{1}{\sqrt{2}}$ and choose $\lambda_{1}=1.99$, $\lambda_{2}=-\frac{1}{5}$ then all the above-mentioned condition are satisfied. Now we choose $\sigma_{0}=\frac{\sigma}{\operatorname{tr}(\sigma)}$. With this choice we have $\operatorname{tr}(\sigma_{0})=1$ and $(\Pi_{M}-\Pi_{N})\otimes\sigma_{0}\leq\mathsf{A}$. This example can trivially be generalized to incorporate bipartite systems suitable in the context of representing maps.
B.5 Proof of Lemma 4
Consider the representing map $\mathsf{T}:\mathcal{L}\left(A\otimes B\right)\rightarrow\mathcal{L}\left(C\otimes D\right)$ of a supermap $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ and let $\Theta^{*}$ be the adjoint of $\Theta$. From the definition of adjoint, for all $\mathsf{X}\in\mathcal{L}\left(A\otimes B\right)$ and $\mathsf{Y}\in\mathcal{L}\left(C\otimes D\right)$, we have
$\mathsf{T}^{*}(\mathsf{Y}),\mathsf{X}\right>$ =$\mathsf{Y},T(\mathsf{X})\right>=\left<\widehat{\mathsf{C}}_{\Gamma_{\mathsf{Y}}},\widehat{\mathsf{C}}_{\Theta(\Gamma_{\mathsf{X}})}\right>=\mathrm{tr}\left[\widehat{\mathsf{C}}^{\dagger}_{\Gamma_{\mathsf{Y}}}\widehat{\mathsf{C}}_{\Theta\left(\Gamma_{\mathsf{X}}\right)}\right].$ (112)
Let $\{e^{Q}_{ij}\}$ be a canonical basis for the space $\mathcal{L}\left(Q\right)$, where $Q=\{A,B,C,D\}$. Then, the above equation leads to
$\mathrm{tr}\left[\widehat{\mathsf{C}}^{\dagger}_{\Gamma_{\mathsf{Y}}}\widehat{\mathsf{C}}_{\Theta\left(\Gamma_{\mathsf{X}}\right)}\right]$
$=\sum_{i,j}\sum_{k,l}\mathrm{tr}\left[\left({e_{ij}^{C}}^{\dagger}\otimes\left\{\Gamma_{\mathsf{Y}}\left(e_{ij}^{C}\right)\right\}^{\dagger}\right)\left({e_{kl}^{C}}\otimes\Theta\left(\Gamma_{\mathsf{X}}\right)\left({e_{kl}^{C}}\right)\right)\right]$
$=\sum_{i,j}\sum_{k,l}\delta_{ik}\delta_{jl}\mathrm{tr}\left[\left\{\Gamma_{\mathsf{Y}}\left({e_{ij}^{C}}\right)\right\}^{\dagger}\Theta\left(\Gamma_{\mathsf{X}}\right)\left({e_{kl}^{C}}\right)\right]=\sum_{i,j}\mathrm{tr}\left[\left\{\Gamma_{Y}\left({e_{ij}^{C}}\right)\right\}^{\dagger}\Theta\left(\Gamma_{\mathsf{X}}\right)\left({e_{ij}^{C}}\right)\right]=\left<\Gamma_{\mathsf{Y}},\Theta(\Gamma_{\mathsf{X}})\right>.$ (113)
Using the definition of the adjoint of a supermap $\Theta$, we obtain
$\Gamma_{\mathsf{Y}},\Theta(\Gamma_{\mathsf{X}})\right>=\left<\Theta^{*}(\Gamma_{\mathsf{Y}}),\Gamma_{\mathsf{X}}\right>$ $=\sum_{i,j}\left<\Theta^{*}\left(\Gamma_{\mathsf{Y}}\right)\left({e_{ij}^{A}}\right),\Gamma_{\mathsf{X}}\left({e_{ij}^{A}}\right)\right>$
$=\sum_{i,j}\sum_{k,l}\delta_{ik}\delta_{jl}\left<\Theta^{*}\left(\Gamma_{\mathsf{Y}}\right)\left({e_{ij}^{A}}\right),\Gamma_{\mathsf{X}}\left({e_{kl}^{A}}\right)\right>$
$=\sum_{i,j}\sum_{k,l}\left<{e_{ij}^{A}},{e_{kl}^{A}}\right>\left<\Theta^{*}\left(\Gamma_{\mathsf{Y}}\right)\left({e_{ij}^{A}}\right),\Gamma_{\mathsf{X}}\left({e_{kl}^{A}}\right)\right>$
=$\sum_{i,j}{e_{ij}^{A}}\otimes\Theta^{*}\left(\Gamma_{\mathsf{Y}}\right)\left({e_{ij}^{A}}\right),\sum_{k,l}{e_{kl}^{A}}\otimes\Gamma_{\mathsf{X}}\left({e_{kl}^{A}}\right)\right>=\left<\widehat{\mathsf{C}}_{\Theta^{*}(\Gamma_{\mathsf{Y}})},\mathsf{X}\right>.$
As one can choose $\mathsf{X}$ arbitrarily, it can be concluded that $\mathsf{T}^{*}(\mathsf{Y})=\mathsf{C}_{\Theta^{*}(\Gamma_{\mathsf{Y}})}$.
Appendix C Proofs of statements in Section III
C.1 Proof of monotonicity of relative entropy under quantum superchannel
Let $\mathcal{N}_{A^{\prime}\rightarrow A}$ and $\mathcal{M}_{A^{\prime}\rightarrow A}$ be a quantum channel and a CP map, respectively, and let $\Theta:\mathcal{L}\left(A^{\prime},A\right)\rightarrow\mathcal{L}\left(C^{\prime},C\right)$ be a superchannel. To prove that the proposed inequality holds, we start with the left-hand side, i.e., $D\left(\Theta\left(\mathcal{N}\right)\|\Theta\left(\mathcal{M}\right)\right)$. Now, considering a general superchannel $\Theta$, whose action on any CP map is defined in Eq.(13), we can write
$D\left[\Theta\left(\mathcal{N}\right)\;\middle\|\;\Theta\left(\mathcal{M}\right)\right]$
[…]
$\leq\underset{\Psi_{A^{\prime}EB}}{\mathrm{sup}}D\left(\left(V^{\prime}_{AE\rightarrow C}\hskip 1.42271pt\circ\left(\mathcal{N}_{A^{\prime}\rightarrow A}\otimes\mathrm{id}_{E}\right)\otimes\mathrm{id}_{B}\right)\Psi_{A^{\prime}EB}\;\middle\|\;\left(V^{\prime}_{AE\rightarrow C}\hskip 1.42271pt\circ\left(\mathcal{M}_{A^{\prime}\rightarrow A}\otimes\mathrm{id}_{E}\right)\otimes\mathrm{id}_{B}\right)\Psi_{A^{\prime}EB}\right),$
where the inequality in the third line follows from the fact that the set of pure states obtained by varying $\Psi_{C^{\prime}B}$ and applying $U_{C^{\prime}\rightarrow A^{\prime}E}$ is smaller in comparison to the set of all pure states in the system $A^{\prime}EB$, i.e, $\{\Psi_{A^{\prime}EB}\}$. In other words, supremum of a set is always greater than or equal to supremum over its subsets. Let us define $\left(\left(\mathcal{N}_{A^{\prime}\rightarrow A}\otimes\mathrm{id}_{E}\right)\otimes\mathrm{id}_{B}\right)\Psi_{A^{\prime}EB}:=\rho^{\Psi}_{ABE}$, which is also a density operator since $\mathcal{N}_{A^{\prime}\rightarrow A}$ is a quantum channel, and $\left(\left(\mathcal{M}_{A^{\prime}\rightarrow A}\otimes\mathrm{id}_{E}\right)\otimes\mathrm{id}_{B}\right)\Psi_{A^{\prime}EB}:=\sigma^{\Psi}_{ABE}$, which is a positive operator, because map $\mathcal{M}$ is CP but not trace preserving. Now, we have
$\underset{\Psi_{A^{\prime}EB}}{\mathrm{sup}}D\left(\left(V^{\prime}_{AE\rightarrow C}\hskip 1.42271pt\circ\left(\mathcal{N}_{A^{\prime}\rightarrow A}\otimes\mathrm{id}_{E}\right)\otimes\mathrm{id}_{B}\right)\Psi_{A^{\prime}EB}\;\middle\|\;\left(V^{\prime}_{AE\rightarrow C}\hskip 1.42271pt\circ\left(\mathcal{M}_{A^{\prime}\rightarrow A}\otimes\mathrm{id}_{E}\right)\otimes\mathrm{id}_{B}\right)\Psi_{A^{\prime}EB}\right)$
$=\underset{\Psi_{A^{\prime}EB}}{\mathrm{sup}}D\left(\left(V^{\prime}_{AE\rightarrow C}\otimes\mathrm{id}_{B}\right)\rho^{\Psi}_{ABE}\;\middle\|\;\left(V^{\prime}_{AE\rightarrow C}\otimes\mathrm{id}_{B}\right)\sigma^{\Psi}_{AEB}\right)$
$\leq\underset{\Psi_{A^{\prime}EB}}{\mathrm{sup}}D\left(\rho^{\Psi}_{ABE}\;\middle\|\;\sigma^{\Psi}_{AEB}\right)$
$=\underset{\Psi_{A^{\prime}EB}}{\mathrm{sup}}D\left(\left(\left(\mathcal{N}_{A^{\prime}\rightarrow A}\otimes\mathrm{id}_{E}\right)\otimes\mathrm{id}_{B}\right)\Psi_{A^{\prime}EB}\;\middle\|\;\left(\left(\mathcal{M}_{A^{\prime}\rightarrow A}\otimes\mathrm{id}_{E}\right)\otimes\mathrm{id}_{B}\right)\Psi_{A^{\prime}EB}\right)=D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right],$ (114)
where the third line follows from the monotonicity of relative entropy under a quantum channel, and the last line simply follows from the definition of relative entropy between two completely positive maps. Thus, we have the inequality
$D\left[\Theta\left(\mathcal{N}\right)\|\Theta\left(\mathcal{R}\right)\right]\leq D\left[\mathcal{N}\|\mathcal{R}\right].$ (115)
C.2 Proof of monotonicity of relative entropy under CP and TP preserving supermaps
Since $\mathcal{N}_{A\rightarrow A^{\prime}}$ and $\mathcal{M}_{A\rightarrow A^{\prime}}$ are quantum channels and the supermap $\Theta:\mathcal{L}\left(A^{\prime},A\right)\rightarrow\mathcal{L}\left(C^{\prime},C\right)$ is CP- and TP-preserving map, we have that $\Theta(\mathcal{N})$ and $\Theta(\mathcal{M})$ are also quantum channels. Here the quantum channels $\mathcal{M},\mathcal{N}$ and supermap $\Theta$ are such that the state that realizes the supremum of relative entropy functionals $D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]$ and $D\left[\Theta(\mathcal{N})\;\middle\|\;\Theta(\mathcal{M})\right]$ are element of $\mathsf{FRank}\left(\widetilde{A}\otimes{A}\right)\cap\mathcal{D}\left(\widetilde{A}\otimes{A}\right)$ and $\mathsf{FRank}\left(\widetilde{C}\otimes{C}\right)\cap\mathcal{D}\left(\widetilde{C}\otimes{C}\right)$, respectively. We then have $D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]=D\left({\mathsf{C}}_{\mathcal{N}}^{\Psi}\;\middle\|\;{\mathsf{C}}_{\mathcal{M}}^{\Psi}\right)$ and $D\left[\Theta(\mathcal{N})\;\middle\|\;\mathcal{M}(\Theta)\right]=D\left({\mathsf{C}}_{\Theta(\mathcal{N})}^{\Phi}\;\middle\|\;{\mathsf{C}}_{\Theta(\mathcal{M})}^{\Phi}\right)$, where $\Psi\in\mathsf{FRank}\left(\widetilde{A}\otimes{A}\right)\cap\mathcal{D}\left(\widetilde{A}\otimes{A}\right)$ and $\Phi\in\mathsf{FRank}\left(\widetilde{C}\otimes{C}\right)\cap\mathcal{D}\left(\widetilde{C}\otimes{C}\right)$ (see Appendix D for detailed calculations). Further, $\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}$ and $\mathsf{C}_{\Theta\left(\mathcal{M}\right)}^{\Phi}$ are Choi states of channels $\Theta\left(\mathcal{N}\right)$ and $\Theta\left(\mathcal{N}\right)$, respectively and $\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}$ and $\mathsf{C}_{\Theta\left(\mathcal{M}\right)}^{\Phi}$ are Choi states of channels $\Theta\left(\mathcal{N}\right)$ and $\Theta\left(\mathcal{M}\right)$, respectively. Note that the Choi states $\mathsf{C}_{\mathcal{N}}^{\Psi}$, $\mathsf{C}_{\Theta(\mathcal{N})}^{\Phi}$ and $\mathsf{C}_{\mathcal{M}}^{\Psi}$, $\mathsf{C}_{\Theta(\mathcal{M})}^{\Phi}$ are connected via the representing map $\mathfrak{T}$ of the supermap $\Theta$ as $\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}=\mathfrak{T}\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)$ and $\mathsf{C}_{\Theta\left(\mathcal{M}\right)}^{\Phi}=\mathfrak{T}\left(\mathsf{C}_{\mathcal{M}}^{\Psi}\right)$. As the supermap $\Theta$ is CP-preserving, the representing map $\mathfrak{T}$ must be a positive map. However, $\mathfrak{T}$ may not be a trace preserving map in general. In the case where it is a trace non-increasing, we can always replace it with the following trace preserving map: $\mathfrak{T}^{\prime}(\mathsf{X}):=\mathfrak{T}(\mathsf{X})+\left[\mathrm{tr}(\mathsf{X})-\mathrm{tr}(\mathfrak{T}(\mathsf{X}))\right]\sigma_{0}$, where $\sigma_{0}\in\mathcal{D}(A\otimes A^{\prime})$ is some fixed state and $\mathsf{X}\in\mathcal{L}(C\otimes C^{\prime})$, and we have $\mathfrak{T}^{\prime}(\mathsf{C}^{\Psi}_{\mathcal{N}})=\mathfrak{T}(\mathsf{C}^{\Psi}_{\mathcal{N}})$. A similar equality holds for $\mathsf{C}^{\Psi}_{\mathcal{M}}$ as well. Then (MHR, 17, Theorem 1) leads to
$D\left({\mathsf{C}}_{\mathcal{N}}^{\Psi}\;\middle\|\;{\mathsf{C}}_{\mathcal{M}}^{\Psi}\right)\geq D\left(\mathfrak{T}^{\prime}\left({\mathsf{C}}_{\mathcal{N}}^{\Psi}\right)\;\middle\|\;\mathfrak{T}^{\prime}\left(\mathsf{C}_{\mathcal{M}}^{\Psi}\right)\right),$ (116)
which implies
$D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]\geq D\left[\Theta\left(\mathcal{N}\right)\;\middle\|\;\Theta\left(\mathcal{M}\right)\right].$ (117)
If we relax the TP-preserving condition on supermap $\Theta$ and take it to be a CP-preserving map, then the representing map $\mathfrak{T}$ will be a positive map but we can not define the trace preserving map $\mathfrak{T}^{\prime}$ such that $\mathfrak{T}^{\prime}(\mathsf{C}^{\Psi}_{\mathcal{N}})=\mathfrak{T}(\mathsf{C}^{\Psi}_{\mathcal{N}})$, for all quantum channels $\mathcal{N}$. However, we can find a condition for $\mathfrak{T}$ to be a trace preserving map. The representing map $\mathsf{T}$ is trace preserving if $\Theta^{*}$ is completely depolarising preserving map (see Lemma 6). We want to find a similar condition for $\mathfrak{T}$. The map $\mathfrak{T}$ is trace preserving if and only if its adjoint is unital, i.e., $\mathfrak{T}^{*}\left(\mathbbm{1}\right)=\mathbbm{1}$, which can also be written as $\mathfrak{T}^{*}\left(\mathsf{C}^{\Psi}_{\Theta^{-1}_{\Psi}(\mathcal{R}_{A\rightarrow A^{\prime}})}\right)=\mathsf{C}^{\Phi}_{\Theta^{-1}_{\Phi}(\mathcal{R}_{C\rightarrow C^{\prime}})}$. Using Eq. (32), we obtain that the following condition on supermap $\Theta$ implies that $\mathfrak{T}$ is trace preserving:
$\Theta^{*}\circ\Theta^{*}_{\Psi}\left(\mathcal{R}_{A\rightarrow A^{\prime}}\right)=\Theta^{*}_{\Phi}\left(\mathcal{R}_{C\rightarrow C^{\prime}}\right).$ (118)
If $\mathcal{N},\mathcal{M}$ and $\Theta$ are such tha $\Psi$ and $\Phi$ are maximally entangled states, then, $\Theta^{*}_{\Psi}$ and $\Theta^{*}_{\Phi}$ becomes identity supermap (see the definition of $\Theta_{\Psi},\Theta_{\Phi}$ in Eq. (28)) and $\Theta^{*}$ becomes $\mathcal{R}$-preserving map.
C.3 Proof of Proposition 4
From our definition of relative entropy between a quantum channel and a positive map, we have
$D\left[\mathcal{N}\otimes\widetilde{\mathcal{N}}\;\middle\|\;\mathcal{M}\otimes\widetilde{\mathcal{M}}\right]$
$=\underset{\Psi_{RAB}}{\mathrm{sup}}~{}D\left(\left(\mathrm{id}_{R}\otimes\mathcal{N}\otimes\widetilde{N}\right)(\Psi_{RAB})\;\middle\|\;\left(\mathrm{id}_{R}\otimes\mathcal{M}\otimes\widetilde{\mathcal{M}}\right)(\Psi_{RAB})\right)$
$\geq\underset{\Psi_{R_{1}A}\otimes\Psi_{R_{2}B}}{\mathrm{sup}}~{}D\left(\left(\mathrm{id}_{R_{1}}\otimes\mathcal{N}\right)\Psi_{R_{1}A}\otimes\left(\mathrm{id}_{R_{2}}\otimes\widetilde{N}\right)\Psi_{R_{2}B}\;\middle\|\;\left(\mathrm{id}_{R_{1}}\otimes\mathcal{M}\right)\Psi_{R_{1}A}\otimes\left(\mathrm{id}_{R_{2}}\otimes\widetilde{M}\right)\Psi_{R_{2}B}\right)$
$=\underset{\Psi_{R_{1}A}\otimes\Psi_{R_{2}B}}{\mathrm{sup}}~{}\left(D\left(\left(\mathrm{id}_{R_{1}}\otimes\mathcal{N}\right)\Psi_{R_{1}A}\;\middle\|\;\left(\mathrm{id}_{R}\otimes\mathcal{M}\right)\Psi_{R_{1}A}\right)+D\left(\left(\mathrm{id}_{R_{2}}\otimes\widetilde{\mathcal{N}}\right)\Psi_{R_{2}B}\;\middle\|\;\left(\mathrm{id}_{R_{2}}\otimes\widetilde{\mathcal{M}}\right)\Psi_{R_{2}B}\right)\right)$
$=D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]+D[\widetilde{\mathcal{N}}\;\|\;\widetilde{\mathcal{M}}],$ (119)
where in the second line we have used the fact that supremum of some function over a bigger set is always greater than its supremum over a smaller set, and in the third line we have used the additivity of quantum relative entropy, i.e., for positive semidefinite operators $\rho_{1},\rho_{2},\sigma_{1}$, and $\sigma_{2}$, we have, $D\left(\rho_{1}\otimes\rho_{2}\|\sigma_{1}\otimes\sigma_{2}\right)=\mathrm{tr}\left(\rho_{2}\right)D\left(\rho_{1}\|\sigma_{1}\right)+\mathrm{tr}\left(\rho_{1}\right)D\left(\rho_{2}\|\sigma_{2}\right)$, and the last line follows simply from the definition of relative entropy between a quantum channel and a CP map.
C.4 Axioms for state and channel entropy
We can write von Neumann entropy in terms of relative entropy as follows
$S(A)_{\rho}=-D\left(\rho_{A}\|\mathbbm{1}_{A}\right).$ (120)
The above definition of entropy of a quantum state satisfies some physically motivated axioms:
  1. 1.
    Under the action of random unitary channels, it increases: let $\mathcal{U}(\rho)=\sum_{i}p_{i}U_{i}\rho U_{i}^{\dagger}$ be a random unitary channel with $U_{i}$’s being unitary operators. Then $S(\mathcal{U}(\rho))\geq S(\rho)$ holds.
  2. 2.
    It is additive under tensor product of quantum states: for $\rho\in\mathcal{D}(A)$ and $\sigma\in\mathcal{D}(A^{\prime})$, we have $S(\rho\otimes\sigma)=S(\rho)+S(\sigma)$.
  3. 3.
    It is maximum for the maximally mixed state and zero for all pure states.
It was advocated in GW (21) that the entropy functional for a quantum channel should satisfy the following physically motivated axioms:
  1. 1.
    It should increase under the action of random unitary superchannels.
  2. 2.
    It should be additive under tensor product of quantum channels.
  3. 3.
    It should be maximum for the completely depolarizing channel and should be zero for a replacement channel that outputs a pure state.
C.5 Proof of additivity of channel entropy
Let $\mathcal{N}_{A_{1}\rightarrow B_{1}}$ and $\mathcal{M}_{A_{2}\rightarrow B_{2}}$ be two quantum channels, and $\mathcal{R}_{A_{1}\rightarrow B_{1}}$, $\mathcal{R}_{A_{2}\rightarrow B_{2}}$ be the corresponding completely depolarising map. It can be shown that additivity will hold if the following equality also holds
$D\left[\mathcal{N}_{A_{1}\rightarrow B_{1}}\otimes{\mathcal{M}}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{1}\rightarrow B_{1}}\otimes{\mathcal{R}}_{A_{2}\rightarrow B_{2}}\right]=D\left[\mathcal{N}_{A_{1}\rightarrow B_{1}}\;\middle\|\;\mathcal{R}_{A_{1}\rightarrow B_{1}}\right]+D\left[\mathcal{M}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{2}\rightarrow B_{2}}\right].$
To show that the above equality holds, it is sufficient to prove that the quantity on the left-hand side is less than or equal to the quantity on the right-hand side and then the other way around is also true. If both inequalities are true, then, both sides must be equal. First, we will show the ``"``"`$\geq$"` ` ≥ " part, i.e.,
$D\left[\mathcal{N}_{A_{1}\rightarrow B_{1}}\otimes{\mathcal{M}}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{1}\rightarrow B_{1}}\otimes{\mathcal{R}}_{A_{2}\rightarrow B_{2}}\right]\geq D\left[\mathcal{N}_{A_{1}\rightarrow B_{1}}\;\middle\|\;\mathcal{R}_{A_{1}\rightarrow B_{1}}\right]+D\left[\mathcal{M}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{2}\rightarrow B_{2}}\right].$
Let us define the action of channel $\mathcal{M}_{A_{2}\rightarrow B_{2}}$ and the CP map $\mathcal{R}_{A_{2}\rightarrow B_{2}}$ on an arbitrary pure state $\Psi_{RA_{1}A_{2}}$ as $\mathcal{M}_{A_{2}\rightarrow B_{2}}\left(\Psi_{RA_{1}A_{2}}\right):=\rho_{RA_{1}B_{2}}$, and $\mathcal{R}_{A_{2}\rightarrow B_{2}}\left(\Psi_{RA_{1}A_{2}}\right):=\sigma_{RA_{1}B_{2}}$, where $\rho_{RA_{1}B_{2}}$ is a density operator and $\sigma_{RA_{1}B_{2}}$ is a positive operator. We then have
$D\left[\mathcal{N}_{A_{1}\rightarrow B_{1}}\otimes{\mathcal{M}}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{1}\rightarrow B_{1}}\otimes{\mathcal{R}}_{A_{2}\rightarrow B_{2}}\right]$
$=\sup_{\Psi_{RA_{1}A_{2}}}D\left(\left(\mathrm{id}_{R}\otimes\mathcal{N}_{A_{1}\rightarrow B_{1}}\otimes\mathcal{M}_{A_{2}\rightarrow B_{2}}\right)\Psi_{RA_{1}A_{2}}\;\middle\|\;\left(\mathrm{id}_{R}\otimes\mathcal{R}_{A_{1}\rightarrow B_{1}}\otimes\mathcal{R}_{A_{2}\rightarrow B_{2}}\right)\Psi_{RA_{1}A_{2}}\right)$
$=\sup_{\Psi_{RA_{1}A_{2}}}D\left(\left(\mathrm{id}_{R}\otimes\mathcal{N}_{A_{1}\rightarrow B_{1}}\otimes\mathrm{id}_{B_{2}}\right)\rho_{RA_{1}B_{2}}\;\middle\|\;\left(\mathrm{id}_{R}\otimes\mathcal{R}_{A_{1}\rightarrow B_{1}}\otimes\mathrm{id}_{B_{2}}\right)\sigma_{RA_{1}B_{2}}\right)$
$=\sup_{\Psi_{R^{\prime}A_{1}}}D\left(\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{N}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}\;\middle\|\;\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{R}_{A_{1}\rightarrow B_{1}}\right)\sigma_{R^{\prime}A_{1}}\right),$ (121)
where $R^{\prime}=RB_{2}$ is relabeling of the system $RB_{2}$. Now, using the definition of the completely depolarising map, the operator $\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{R}_{A_{1}\rightarrow B_{1}}\right)\sigma_{R^{\prime}A_{1}}$ can be simplified to $\sigma_{R^{\prime}}\otimes\mathbbm{1}_{B_{1}}$. Also, let us define $A_{R^{\prime}B_{1}}:=\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{N}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}$, then we observe $A_{R^{\prime}}=\mathrm{tr}_{B_{1}}\left(A_{R^{\prime}B_{1}}\right)=\left(\mathrm{id}_{R^{\prime}}\otimes\mathrm{tr_{B_{1}}}\right)\left(\mathrm{id_{R^{\prime}}}\otimes\mathcal{N}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}=\left(\mathrm{id}_{R^{\prime}}\otimes\mathrm{tr_{B_{1}}}\circ\mathcal{N}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}=\rho_{R^{\prime}}$. Thus, we can write
$D\left(\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{N}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}\;\middle\|\;\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{R}_{A_{1}\rightarrow B_{1}}\right)\sigma_{R^{\prime}A_{1}}\right)$ $=-S(A_{R^{\prime}B_{1}})-\mathrm{tr}\left(A_{R^{\prime}}\log\sigma_{R^{\prime}}\right),$ (122)
$D\left(\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{N}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}\;\middle\|\;\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{R}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}\right)$ $=-S\left(A_{R^{\prime}B_{1}}\right)-\mathrm{tr}\left(A_{R^{\prime}}\log\rho_{R^{\prime}}\right).$ (123)
Using the above equations, we see that
$D\left(\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{N}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}\;\middle\|\;\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{R}_{A_{1}\rightarrow B_{1}}\right)\sigma_{R^{\prime}A}\right)$
$=D\left(\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{N}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}\;\middle\|\;\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{R}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A}\right)+D\left(\rho_{R^{\prime}}\;\middle\|\;\sigma_{R^{\prime}}\right).$
Define
$X$ $:=D\left(\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{N}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}\;\middle\|\;\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{R}_{A_{1}\rightarrow B_{1}}\right)\sigma_{R^{\prime}A}\right),$ (124)
$Y$ $:=D\left(\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{N}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A_{1}}\;\middle\|\;\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{R}_{A_{1}\rightarrow B_{1}}\right)\rho_{R^{\prime}A}\right).$ (125)
That is, $X=Y+D\left(\rho_{R^{\prime}}\;\middle\|\;\sigma_{R^{\prime}}\right)$. Now exploiting the monotonicity of the relative entropy under partial trace, we have
$X$ $\leq Y+D\left(\rho_{R^{\prime}B_{2}A_{1}}\;\middle\|\;\sigma_{R^{\prime}B_{2}A_{1}}\right)$
$=Y+D\left(\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{M}_{A_{2}\rightarrow B_{2}}\otimes\mathrm{id}_{A_{1}}\right)\Psi_{R^{\prime}A_{2}A_{1}}\;\middle\|\;\left(\mathrm{id}_{R^{\prime}}\otimes\mathcal{R}_{A_{2}\rightarrow B_{2}}\otimes\mathrm{id}_{A_{1}}\right)\Psi_{R^{\prime}A_{2}A_{1}}\right)$
$=Y+D\left(\left(\mathrm{id}_{R^{\prime}}\otimes\mathrm{id}_{A_{1}}\otimes\mathcal{M}_{A_{2}\rightarrow B_{2}}\right)\Psi_{R^{\prime}A_{1}A_{2}}\;\middle\|\;\left(\mathrm{id}_{R^{\prime}}\otimes\mathrm{id}_{A_{1}}\otimes\mathcal{R}_{A_{2}\rightarrow B_{2}}\right)\Psi_{R^{\prime}A_{1}A_{2}}\right).$ (126)
Taking supremum on both sides of the above inequality with respect to state $\Psi_{R^{\prime}A_{1}A_{2}}$, we then obtain
$\sup_{\Psi_{R^{\prime}A_{1}A_{2}}}X\leq\sup_{\Psi_{R^{\prime}A_{1}A_{2}}}Y+D\left[\mathcal{M}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{2}\rightarrow B_{2}}\right],$ (127)
which also implies
$D\left[\mathcal{N}_{A_{1}\rightarrow B_{1}}\otimes\mathcal{M}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{1}\rightarrow B_{1}}\otimes\mathcal{R}_{A_{2}\rightarrow B_{2}}\right]$ $\leq D\left[\mathcal{N}_{A_{1}\rightarrow B_{1}}\;\middle\|\;\mathcal{R}_{A_{1}\rightarrow B_{1}}\right]+D\left[\mathcal{M}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{2}\rightarrow B_{2}}\right].$ (128)
By multiplying with the minus sign on both sides of the above inequality and using the definition of the entropy of the quantum channel, we obtain the desired inequality
$S\left[\mathcal{N}_{A_{1}\rightarrow B_{1}}\otimes\mathcal{M}_{A_{2}\rightarrow B_{2}}\right]\geq S\left[\mathcal{N}_{A_{1}\rightarrow B_{1}}\right]+S\left[\mathcal{M}_{A_{2}\rightarrow B_{2}}\right].$ (129)
Now, we will show the ``"``"`$\leq$"` ` ≤ " part, i.e., $S\left(\mathcal{N}_{A_{1}\rightarrow B_{1}}\otimes\mathcal{M}_{A_{2}\rightarrow B_{2}}\right)\leq S\left(\mathcal{N}_{A_{1}\rightarrow B_{1}}\right)+S\left(\mathcal{M}_{A_{2}\rightarrow B_{2}}\right)$. From the definition of relative entropy of quantum channels, we have
$D\left[\mathcal{N}_{A_{1}\rightarrow B_{1}}\otimes\mathcal{M}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{1}\rightarrow B_{1}}\otimes\mathcal{R}_{A_{2}\rightarrow B_{2}}\right]$
$=\sup_{\Psi_{EA_{1}A_{2}}}D\left((\mathrm{id}_{E}\otimes{\mathcal{N}}_{A_{1}\rightarrow B_{1}}\otimes{\mathcal{M}}_{A_{2}\rightarrow B_{2}})\Psi_{EA_{1}A_{2}}\;\middle\|\;(\mathrm{id}_{E}\otimes{\mathcal{R}}_{A_{1}\rightarrow B_{1}}\otimes{\mathcal{R}}_{A_{2}\rightarrow B_{2}})\Psi_{{EA_{1}A_{2}}}\right),$ (130)
where we can take the ancilla $E$ isomorphic to $A_{1}\otimes A_{2}$. In other words, we can take $E=R_{1}\otimes R_{2}$, such that, reference systems $R_{1}$ and $R_{2}$ are isomorphic to $A_{1}$ and $A_{2}$, respectively. Now, using the facts that $R_{1}\otimes R_{2}\otimes A_{1}\otimes A_{2}$ is isomorphic to $R_{1}\otimes A_{1}\otimes R_{2}\otimes A_{2}$, and supremum over a set is always greater than or equal to the supremum over its subsets, we obtain
$\sup_{\Psi_{R_{1}R_{2}A_{1}A_{2}}}D\left((\mathrm{id}_{R_{1}R_{2}}\otimes{\mathcal{N}}_{A_{1}\rightarrow B_{1}}\otimes{\mathcal{M}}_{A_{2}\rightarrow B_{2}})\Psi_{EA_{1}A_{2}}\;\middle\|\;(\mathrm{id}_{R_{1}R_{2}}\otimes{\mathcal{R}}_{A_{1}\rightarrow B_{1}}\otimes{\mathcal{R}}_{A_{2}\rightarrow B_{2}})\Psi_{{EA_{1}A_{2}}}\right)$
$\geq\sup_{\widetilde{\Psi}}D\left((\mathrm{id}_{R_{1}}\otimes{\mathcal{N}}_{A_{1}\rightarrow B_{1}})\Psi_{R_{1}A_{1}}\otimes(\mathrm{id}_{R_{2}}\otimes{\mathcal{M}}_{A_{2}\rightarrow B_{2}})\Psi_{R_{2}A_{2}}\;\middle\|\;(\mathrm{id}_{R_{1}}\otimes{\mathcal{R}}_{A_{1}\rightarrow B_{1}})\Psi_{R_{1}A_{1}}\otimes(\mathrm{id}_{R_{2}}\otimes{\mathcal{R}}_{A_{2}\rightarrow B_{2}})\Psi_{R_{2}A_{2}}\right)$
$=\sup_{\widetilde{\Psi}}\left\{\right.\mathrm{tr}\left[(\mathrm{id}_{R_{2}}\otimes{\mathcal{M}}_{A_{2}\rightarrow B_{2}})\Psi_{R_{2}A_{2}}\right]~{}D\left((\mathrm{id}_{R_{1}}\otimes{\mathcal{N}}_{A_{1}\rightarrow B_{1}})\Psi_{R_{1}A_{1}}\;\middle\|\;(\mathrm{id}_{R_{1}}\otimes{\mathcal{R}}_{A_{1}\rightarrow B_{1}})\Psi_{R_{1}A_{1}}\right)$
$\hskip 28.45274pt+\mathrm{tr}\left[(\mathrm{id}_{R_{1}}\otimes{\mathcal{N}}_{A_{1}\rightarrow B_{1}})\Psi_{R_{1}A_{1}}\right]~{}D\left((\mathrm{id}_{R_{2}}\otimes{\mathcal{M}}_{A_{2}\rightarrow B_{2}})\Psi_{R_{2}A_{2}}\;\middle\|\;(\mathrm{id}_{R_{2}}\otimes{\mathcal{R}}_{A_{2}\rightarrow B_{2}})\Psi_{R_{2}A_{2}}\right)\left.\right\}$
$=\sup_{\widetilde{\Psi}_{1}}D\left((\mathrm{id}_{R_{1}}\otimes{\mathcal{N}}_{A_{1}\rightarrow B_{1}})\Psi_{R_{1}A_{1}}\;\middle\|\;(\mathrm{id}_{R_{1}}\otimes{\mathcal{R}}_{A_{1}\rightarrow B_{1}})\Psi_{R_{1}A_{1}}\right)$
$\hskip 28.45274pt+\sup_{\widetilde{\Psi}_{2}}D\left((\mathrm{id}_{R_{2}}\otimes{\mathcal{M}}_{A_{2}\rightarrow B_{2}})\Psi_{R_{2}A_{2}}\;\middle\|\;(\mathrm{id}_{R_{2}}\otimes{\mathcal{R}}_{A_{2}\rightarrow B_{2}})\Psi_{R_{2}A_{2}}\right)$
$=D\left[\mathcal{N}_{A_{1}\rightarrow B_{1}}\;\middle\|\;\mathcal{R}_{A_{1}\rightarrow B_{1}}\right]+D\left[\mathcal{M}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{2}\rightarrow B_{2}}\right],$ (131)
where, $\widetilde{\Psi}:=\Psi_{R_{1}A_{1}}\otimes\Psi_{R_{2}A_{2}}$, $\widetilde{\Psi}_{1}:=\Psi_{R_{1}A_{1}}$ and $\widetilde{\Psi}_{2}:=\Psi_{R_{2}A_{2}}$. In the second line we have used the additivity of quantum relative entropy, i.e., for positive semidefinite operators $\rho_{1},\rho_{2},\sigma_{1}$, and $\sigma_{2}$, we have, $D\left(\rho_{1}\otimes\rho_{2}\|\sigma_{1}\otimes\sigma_{2}\right)=\mathrm{tr}\left(\rho_{2}\right)D\left(\rho_{1}\|\sigma_{1}\right)+\mathrm{tr}\left(\rho_{1}\right)D\left(\rho_{2}\|\sigma_{2}\right)$, and the last line simply follows from the definition of relative entropy of quantum channels. Thus, we have the following inequality,
$-D\left[\mathcal{N}_{A_{1}\rightarrow B_{1}}\otimes{\mathcal{M}}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{1}\rightarrow B_{1}}\otimes{\mathcal{R}}_{A_{2}\rightarrow B_{2}}\right]\leq-D\left[\mathcal{N}_{A_{1}\rightarrow B_{1}}\;\middle\|\;\mathcal{R}_{A_{1}\rightarrow B_{1}}\right]-D\left[\mathcal{M}_{A_{2}\rightarrow B_{2}}\;\middle\|\;\mathcal{R}_{A_{2}\rightarrow B_{2}}\right].,$ (132)
which implies that
$S\left[{\mathcal{N}}_{A_{1}\rightarrow B_{1}}\otimes{\mathcal{M}}_{A_{2}\rightarrow B_{2}}\right]\leq S\left[{\mathcal{N}}_{A_{1}\rightarrow B_{1}}]+S[{\mathcal{M}}_{A_{2}\rightarrow B_{2}}\right].$ (133)
Combining inequalities (129) and (133), we conclude that the entropy functional of a quantum channel is additive under tensor product of channels, i.e,
$S\left[{\mathcal{N}}_{A_{1}\rightarrow B_{1}}\otimes{\mathcal{M}}_{A_{2}\rightarrow B_{2}}\right]=S\left[{\mathcal{N}}_{A_{1}\rightarrow B_{1}}\right]+S\left[{\mathcal{M}}_{A_{2}\rightarrow B_{2}}\right].$ (134)
C.6 Entropy gain under positive maps
Lemma 9.
For a given pair of input and output quantum states $\{\sigma\in\mathcal{D}\left(C\right),\rho\in\mathcal{D}\left(A\right\}$, where the output state $\rho$ is always a pure state but the input state could be either a pure or a mixed state, there always exists a unital CP map $\mathcal{F}:\mathcal{L}\left(C\right)\rightarrow\mathcal{L}\left(A\right)$, such that $\rho=\mathcal{F}(\sigma)$.
Proof.
Consider a set $\mathtt{K}_{C}=\{\sigma,\mathbbm{1}\}\subset\mathcal{L}\left(C\right)$, and let $\mathcal{A}_{\mathtt{K}_{C}}$ denote the sub-algebra generated by $\mathtt{K}_{C}$. Clearly, $\mathcal{A}_{\mathtt{K}_{C}}$ is a commutative unital $C^{*}$ sub-algebra of $\mathcal{L}(C)$. Now, we define a map $\mathcal{F}:\mathcal{A}_{\mathtt{K}_{C}}\rightarrow\mathcal{L}\left(A\right)$ by $\mathcal{F}(\sigma)=\rho$, $\mathcal{F}(\sigma^{n})=\rho$, for all $n\in\mathbb{N}/\{0\}$, and $\mathcal{F}(\mathbbm{1}_{C})=\mathbbm{1}_{A}$, and extend it by linearity to the whole space. As $\sigma$ is positive and hence, hermitian, it is clear that the map $\mathcal{F}$ is a ${}^{\ast}-\text{homomorphism}$ which also implies that $\mathcal{F}$ is a positive map. As $\mathcal{A}_{\mathtt{K}_{C}}$ is a commutative unital $C^{*}$ algebra and $\mathcal{L}\left(A\right)$ is a unital $C^{*}$ algebra, we have that $\mathcal{F}$ is completely positive Pau (03). We now employ Arveson’s extension theorem to extend it to $\mathcal{L}\left(C\right)$.
Arveson’s extension theorem Pau (03): Let $\mathfrak{B}$ be a $C^{*}$-algebra and $\mathfrak{D}\subseteq\mathfrak{B}$ an operator system (a hermitian subspace containing the identity of the $C^{*}$-algebra). Let $\mathcal{F}:\mathfrak{D}\rightarrow\mathcal{L}\left(A\right)$ be a completely positive map. Then there exists a completely positive map, $\mathcal{F^{\prime}}:\mathfrak{B}\rightarrow\mathcal{L}\left(A\right)$, which extends $\mathcal{F}$ to the whole algebra $\mathfrak{B}$. Clearly, $\mathcal{A}_{\mathtt{K}_{C}}$ is an operator system. Hence, the map $\mathcal{F}$ extends to the full algebra $\mathcal{L}\left(C\right)$. ∎
Note that our proof will also hold if we take output state $\rho$ to be a mixed state if it satisfies the following condition: $\forall~{}c_{n}\in\mathbb{C}$ if $\sum_{n}c_{n}\sigma^{n}=0$, then $\sum_{n}c_{n}\rho^{n}$ must vanish. Then, we can define a map $\mathcal{F}:\mathcal{A}_{\mathtt{K}_{C}}\rightarrow\mathcal{L}\left(A\right)$ by $\mathcal{F}(\sigma^{n})=\rho^{n}$, for all $n\in\mathbb{N}/\{0\}$, and $\mathcal{F}(\mathbbm{1}_{C})=\mathbbm{1}_{A}$, and extend it by linearity to the whole space.
Next, we derive several entropic inequalities for unital CP maps, CP maps, and positive maps. Our results also reduce to well-known previous results for an appropriate choice of a parameter involved in the remainder term of these bounds. Our main result in this section is Theorem 3, which is on the entropy gain of positive maps.
Proposition 13.
Let $\rho\in\mathcal{D}\left(A\right)$, $\sigma\in\mathcal{D}\left(C\right)$ and $\mathcal{F}:\mathcal{L}\left(C\right)\rightarrow\mathcal{L}\left(A\right)$ be a completely positive unital map that connects $\rho$ and $\sigma$. Then, the von Neumann entropy of $\rho$ and $\sigma$ are related by the following equation.
$S\left(\rho\right)\geq\norm{\mathcal{F}^{*}(\mathbbm{1}_{A})}_{\infty}S\left(\sigma\right),$ (135)
where $\mathcal{F}^{*}:\mathcal{L}\left(A\right)\rightarrow\mathcal{L}\left(C\right)$ is the adjoint of $\mathcal{F}$.
Proof.
Let $\mathcal{F}:\mathcal{L}\left(C\right)\rightarrow\mathcal{L}\left(A\right)$ denote an unital and completely positive map, then we can define a completely positive trace nonincreasing map ~$\widetilde{\mathcal{F}}:\mathcal{L}\left(C\right)\rightarrow\mathcal{L}\left(A\right)$ as follows ~$\widetilde{\mathcal{F}}(\sigma):=\frac{\mathcal{F}(\sigma)}{\norm{\mathcal{F}^{*}(\mathbbm{1}_{A})}_{\infty}}$. Moreover, for any unital and completely positive map $\mathcal{F}:\mathcal{L}\left(C\right)\rightarrow\mathcal{L}\left(A\right)$, and for any positive operator $\sigma\in\mathcal{L}\left(C\right)$, the following operator inequality holds (see Wol (12)), $\mathcal{F}\left(\sigma\right)\log\mathcal{F}\left(\sigma\right)\leq\mathcal{F}\left(\sigma\log\sigma\right)=\norm{\mathcal{F}^{*}\left(\mathbbm{1}_{A}\right)}_{\infty}\widetilde{\mathcal{F}}(\sigma\log\sigma)$. By taking trace on both sides and using the fact that $\widetilde{F}$ is trace nonincreasing, we then have $-S(\left(\mathcal{F}\left(\sigma\right)\right)\leq-\norm{\mathcal{F}^{*}\left(\mathbbm{1}_{A}\right)}_{\infty}S\left(\sigma\right)$, which implies that $S\left(\rho\right)\geq\norm{\mathcal{F}^{*}\left(\mathbbm{1}_{A}\right)}_{\infty}S\left(\sigma\right)$.
Note that, in principle, there may exist more than one unital CP map connecting $\rho$ to $\sigma$  (which we also proved earlier in the beginning of this section). Thus, this lower-bound can be further tightened as $S(\rho)\geq\sup\{\norm{\mathcal{F}^{*}\left(\mathbbm{1}_{A}\right)}_{\infty}\}S\left(\sigma\right)$, where the supremum is taken over all possible maps $\mathcal{F}$ connecting states $\rho$ and $\sigma$. Further, notice that the value of $\sup\{\norm{\mathcal{F}^{*}\left(\mathbbm{1}_{A}\right)}_{\infty}\}$ decides whether the entropy gain is positive or negative. Thus, we can make the following remark: for a completely positive unital map $\mathcal{F}$, entropy gain is non-negative if $\sup\{\norm{\mathcal{F}^{*}\left(\mathbbm{1}_{A}\right)}_{\infty}\}\geq 1$, and entropy gain is negative if $\sup\{\norm{\mathcal{F}^{*}\left(\mathbbm{1}_{A}\right)}_{\infty}\}<1$. Once we assume that map $\mathcal{F}$ is also trace preserving, which implies that $\mathcal{F}^{*}$ is unital, we then have $\sup\{\norm{\mathcal{F}^{*}\left(\mathbbm{1}_{A}\right)}_{\infty}\}=1$, thereby retrieving the well-known result of AU (77). ∎
Lemma 10.
Let $\mathcal{F}:\mathcal{L}\left(C\right)\rightarrow\mathcal{L}\left(A\right)$ be a completely positive map and let the entropy gain of $\sigma\in\mathcal{D}\left({C}\right)$ under the map $\mathcal{F}$ be defined as $\Delta S\left(\sigma,\mathcal{F}\right):=S\left(\mathcal{F}\left(\sigma\right)\right)-S\left(\sigma\right)$. Then, for all $\sigma\in\mathcal{D}\left({C}\right)$ such that $\mathcal{F}\left(\sigma\right)>0$, we have
$\Delta S\left(\sigma,\mathcal{F}\right)\geq\max\left\{D\left(\sigma\;\middle\|\;\sigma_{\alpha}\right),D\left(\sigma\;\middle\|\;\hat{\sigma}_{\alpha}\right)\right\},$ (136)
where $\alpha:=\norm{\mathcal{F}^{*}(\mathbbm{1}_{A})}_{\infty}$, $\sigma_{\alpha}:=\left(\frac{1}{\alpha}\right)^{\alpha}\left(\mathcal{F}^{*}\circ\mathcal{F}(\sigma)\right)^{\alpha}$, and $\hat{\sigma}_{\alpha}:=\frac{1}{\alpha}\mathcal{F}^{*}\left(\mathcal{F}(\sigma)\right)^{\alpha}$.
Proof.
We first define a completely positive, trace nonincreasing map ~$\widetilde{\mathcal{F}}:\mathcal{L}\left(C\right)\rightarrow\mathcal{L}\left(A\right)$ from map $\mathcal{F}:\mathcal{L}\left(C\right)\rightarrow\mathcal{L}\left(A\right)$ as $\widetilde{F}(\sigma):=\frac{1}{\alpha}\mathcal{F}(\sigma)$, where $\alpha:=\norm{\mathcal{F}^{*}(\mathbbm{1}_{A})}_{\infty}$. Now using the definition of adjoint, the entropy gain can be written as $\Delta S\left(\sigma,\mathcal{F}\right)=\operatorname{tr}\left(\sigma\log\sigma\right)-\operatorname{tr}\left(\alpha\sigma\widetilde{\mathcal{F}}^{*}\left(\log\mathcal{F}\left(\sigma\right)\right)\right)$. As $\mathcal{F}\left(\sigma\right)$ is a full rank operator by assumption, we can use the fact that for any sub-unital map $\widetilde{\mathcal{F}}$, we have ~$\widetilde{\mathcal{F}}\left(\log(\sigma)\right)\leq\log\left(\widetilde{\mathcal{F}}\left(\sigma\right)\right)$ if $\mathcal{F}\left(\sigma\right)>0$. Using this, we get $\Delta S\left(\sigma,\mathcal{F}\right)\geq\operatorname{tr}\left(\sigma\log\sigma\right)-\operatorname{tr}\left(\sigma\alpha(\log\widetilde{\mathcal{F}}^{*}\circ\mathcal{F}\left(\sigma\right))\right)$, which implies $\Delta S\left(\sigma,\mathcal{F}\right)\geq D\left(\sigma\;\middle\|\;\sigma_{\alpha}\right)$, where $\sigma_{\alpha}:=\left(\frac{1}{\alpha}\right)^{\alpha}\left(\mathcal{F}^{*}\circ\mathcal{F}(\sigma)\right)^{\alpha}$. Note that we can also write entropy gain as $\Delta S\left(\sigma,\mathcal{F}\right)=\operatorname{tr}\left(\sigma\log\sigma\right)-\operatorname{tr}\left(\widetilde{\mathcal{F}}\left(\sigma\right)\log\left(\mathcal{F}\left(\sigma\right)\right)^{\alpha}\right)$. Then, by following the similar arguments as above, we obtain another bound on entropy gain as $\Delta S\left(\sigma,\mathcal{F}\right)\geq D\left(\sigma\;\middle\|\;\hat{\sigma}_{\alpha}\right)$, where $\hat{\sigma}_{\alpha}:=\frac{1}{\alpha}\mathcal{F}^{*}\left(\mathcal{F}(\sigma)\right)^{\alpha}$. Since both the remainder terms, $D\left(\sigma\;\middle\|\;\sigma_{\alpha}\right)$ and $D\left(\sigma\;\middle\|\;\hat{\sigma}_{\alpha}\right)$, are valid lower bounds on the entropy gain, we have
$\Delta S\left(\sigma,\mathcal{F}\right)\geq\max\left\{D\left(\sigma\;\middle\|\;\sigma_{\alpha}\right),D\left(\sigma\;\middle\|\;\hat{\sigma}_{\alpha}\right)\right\}.$ (137)
This completes the proof of the lemma. ∎
If we take $\mathcal{F}^{*}$ to be unital (or equivalently $\mathcal{F}$ to be a quantum channel) in the above lemma, then we have $\alpha=1$, and we obtain a well-known result BDW (16): $\Delta S\left(\sigma,\mathcal{F}\right)\geq D\left(\sigma\|\mathcal{F}^{*}\circ\mathcal{F}(\sigma)\right)$. We now state some known facts which we will be using to prove our next results.
  1. 1.
    Let $A$ and $B$ be two finite-dimensional Hilbert spaces and let $\mathcal{F}:\mathcal{L}(A)\rightarrow\mathcal{L}(B)$ be a positive linear map such that $\mathcal{F}(\mathbbm{1}_{A})\leq\mathbbm{1}_{B}$. Then for all $\mathsf{X}>0$ we have the following inequality Wol (12):
    $\mathcal{F}\left(\log(\mathsf{X})\right)\leq\log(\mathcal{F}(\mathsf{X})).$ (138)
  2. 2.
    Let $\rho\in\mathcal{D}(A)$ be a quantum state and $\sigma\in\mathcal{L}(A)_{+}$ be a positive semidefinite operator. Then the following limit holds Wil (13):
    $D\left(\rho\;\middle\|\;\sigma\right)=\lim_{\epsilon\rightarrow 0+}D\left(\rho\;\middle\|\;\sigma+\epsilon\mathbbm{1}_{A}\right).$ (139)
  3. 3.
    Let $\rho\in\mathcal{D}(A)$ be a quantum state and $\sigma,\sigma^{\prime}\in\mathcal{L}(A)_{+}$ be positive semidefinite operators such that $\sigma\leq\sigma^{\prime}$. Then we have the following inequality Wil (13) :
    $D\left(\rho\;\middle\|\;\sigma\right)\geq D\left(\rho\;\middle\|\;\sigma^{\prime}\right).$ (140)
In the following, we provide a twofold generalization of Lemma 10: (1) we consider positive maps instead of completely positive maps, (2) we drop the requirement that $\mathcal{F}(\sigma)>0$.
Theorem 3.
Let $\mathcal{F}:\mathcal{L}\left(A\right)\rightarrow\mathcal{L}\left(B\right)$ be a positive map. Then, for all $\rho\in\mathcal{D}\left(A\right)$, the change in entropy after and before applying the map is given by
$S\left(\mathcal{F}\left(\rho\right)\right)-S\left(\rho\right)\geq D\left(\rho\;\middle\|\;\rho_{\alpha}\right),$ (141)
where $\alpha:=\norm{\mathcal{F}^{*}(\mathbbm{1}_{B})}_{\infty}$ and $\rho_{\alpha}:=\frac{1}{\alpha^{\alpha}}\left(\mathcal{F}^{*}\circ\mathcal{F}(\rho)\right)^{\alpha}$.
Proof.
For the given positive map $\mathcal{F}$, we define a positive and trace nonincreasing map $\widetilde{\mathcal{F}}$ as ~$\widetilde{\mathcal{F}}(\rho):=\frac{1}{\alpha}\mathcal{F}(\rho)$, where $\alpha:=\norm{\mathcal{F}^{*}(\mathbbm{1}_{B})}_{\infty}$. Now we define an operator $\sigma:=\mathcal{F}(\rho)+\epsilon\mathbbm{1}_{B}$. Clearly for all $\epsilon>0$, we have $\sigma>0$. Let us consider the following:
$S(\sigma)-S(\rho)$ $=\operatorname{tr}(\rho\log\rho)-\operatorname{tr}(\sigma\log\sigma)$ (142)
$=\operatorname{tr}(\rho\log\rho)-\operatorname{tr}(\alpha\widetilde{\mathcal{F}}(\rho)\log\sigma)-\epsilon\operatorname{tr}(\log\sigma)$
$=\operatorname{tr}(\rho\log\rho)-\alpha\operatorname{tr}(\rho\widetilde{\mathcal{F}}^{*}\log\sigma)-\epsilon\operatorname{tr}(\log\sigma)$
$\geq\operatorname{tr}(\rho\log\rho)-\alpha\operatorname{tr}(\rho\log\widetilde{\mathcal{F}}^{*}(\sigma))-\epsilon\operatorname{tr}(\log\sigma)$
$=(1-\alpha)\operatorname{tr}(\rho\log\rho)+\alpha D\left(\rho\;\middle\|\;\widetilde{\mathcal{F}}^{*}(\sigma)\right)-\epsilon\operatorname{tr}(\log\sigma),$
As $\widetilde{\mathcal{F}}$ is trace nonincreasing we have ~$\widetilde{\mathcal{F}}^{*}(\mathbbm{1}_{B})\leq\mathbbm{1}_{A}$, then the inequality in the fourth line of above relations follows from the inequality (138). Now ~$\widetilde{\mathcal{F}}^{*}(\sigma)=\widetilde{\mathcal{F}}^{*}\circ\mathcal{F}(\rho)+\epsilon\widetilde{\mathcal{F}}^{*}(\mathbbm{1}_{B})\leq\widetilde{\mathcal{F}}^{*}\circ\mathcal{F}(\rho)+\epsilon\mathbbm{1}_{A}$. Thus, using inequality (140), we have $D\left(\rho\;\middle\|\;\widetilde{\mathcal{F}}^{*}(\sigma)\right)\geq D\left(\rho\;\middle\|\;\widetilde{\mathcal{F}}^{*}\circ\mathcal{F}(\rho)+\epsilon\mathbbm{1}_{A}\right)$. Using this, we get
$S(\sigma)-S(\rho)$ $\geq(1-\alpha)\operatorname{tr}(\rho\log\rho)+\alpha D\left(\rho\;\middle\|\;\widetilde{\mathcal{F}}^{*}\circ\mathcal{F}(\rho)+\epsilon\mathbbm{1}_{A}\right)-\epsilon\operatorname{tr}(\log\sigma).$ (143)
Now we use Eq. (139) and the continuity of relative entropy to arrive at the following:
$S(\mathcal{F}(\rho))-S(\rho)$ $\geq(1-\alpha)\operatorname{tr}(\rho\log\rho)+\alpha D\left(\rho\;\middle\|\;\widetilde{\mathcal{F}}^{*}\circ\mathcal{F}(\rho)\right)$ (144)
$=\operatorname{tr}(\rho\log\rho)-\operatorname{tr}\left(\rho\alpha\log\widetilde{\mathcal{F}}^{*}\circ\mathcal{F}(\rho)\right)$
$=\operatorname{tr}\left(\rho\log\rho\right)-\operatorname{tr}\left(\rho\log\left\{\widetilde{\mathcal{F}}^{*}\circ\mathcal{F}(\rho)\right\}^{\alpha}\right)$
$=D\left(\rho\;\middle\|\;\left\{\widetilde{\mathcal{F}}^{*}\circ\mathcal{F}(\rho)\right\}^{\alpha}\right)=D\left(\rho\;\middle\|\;\left(\frac{1}{\alpha}\right)^{\alpha}\left(\mathcal{F}^{*}\circ\mathcal{F}(\rho)\right)^{\alpha}\right).$
This completes the proof of the theorem.
If we take $\mathcal{F}^{*}$ to be unital (or equivalently $\mathcal{F}$ to be a trace preserving) in Theorem 3, then we have $\alpha=1$, and we obtain a well-known result BDW (16): $S\left(\mathcal{F}(\rho)\right)-S(\rho)\geq D\left(\rho\;\middle\|\;\mathcal{F}^{*}\circ\mathcal{F}(\sigma)\right)$.
C.7 Proof of Entropy gain under superchannels
Let us consider a channel $\mathcal{N}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left(B\right)$ such that the state that realizes the supremum in the entropy functional $S\left[\mathcal{N}\right]$ is an element of $\mathsf{FRank}\left(\widetilde{A}\otimes{A}\right)\cap\mathcal{D}\left(\widetilde{A}\otimes{A}\right)$. Then we can write
$D\left[\mathcal{N}\;\middle\|\;\mathcal{R}\right]$ $=\sup_{\rho_{\widetilde{A}A}}D\left(\mathrm{id}_{\widetilde{A}}\otimes\mathcal{N}_{A\rightarrow B}\left(\rho_{\widetilde{A}A}\right)\;\middle\|\;\mathrm{id}_{\widetilde{A}}\otimes\mathcal{R}_{A\rightarrow B}\left(\rho_{\widetilde{A}A}\right)\right)$
$=D\left(\left(\mathrm{id}_{\widetilde{A}}\otimes\mathcal{N}_{A\rightarrow B}\right)\outerproduct{\Psi}{\Psi}\;\middle\|\;\left(\mathrm{id}_{\widetilde{A}}\otimes\mathcal{R}_{A\rightarrow B}\right)\outerproduct{\Psi}{\Psi}\right),$ (145)
where |$\outerproduct{\Psi}{\Psi}\in\mathsf{FRank}\left(\widetilde{A}\otimes{A}\right)\cap\mathcal{D}\left(\widetilde{A}\otimes{A}\right)$. A simple calculation shows that $\left(\mathrm{id}_{\widetilde{A}}\otimes{R}_{A\rightarrow B}\right)\outerproduct{\Psi}{\Psi}=\Psi_{\widetilde{A}}\otimes\mathbbm{1}_{B}$, where $\Psi_{\widetilde{A}}=\mathrm{tr}_{A}\left(\outerproduct{\Psi}{\Psi}\right)$. Since $\Psi_{\widetilde{A}}$ is full rank operator, we have $\left(\mathrm{id}_{\widetilde{A}}\otimes\mathcal{N}_{A\rightarrow B}\right)\outerproduct{\Psi}{\Psi}={\mathsf{C}}_{\mathcal{N}}^{\Psi}$, where ${\mathsf{C}}_{\mathcal{N}}^{\Psi}$ is a Choi state of the quantum channel $\mathcal{N}_{A\rightarrow B}$ with respect to |$\outerproduct{\Psi}{\Psi}$ |. Thus, we can write the entropy of quantum channel $\mathcal{N}_{A\rightarrow B}$ as
$S\left[\mathcal{N}_{A\rightarrow B}\right]$ $=-D\left({\mathsf{C}}_{\mathcal{N}}^{\Psi}\;\middle\|\;\Psi_{\widetilde{A}}\otimes\mathbbm{1}_{B}\right)=S\left({\mathsf{C}}_{\mathcal{N}}^{\Psi}\right)-S\left(\Psi_{\widetilde{A}}\right).$ (146)
Now, consider a superchannel $\Theta:\mathcal{L}\left(A,B\right)\rightarrow\mathcal{L}\left(C,D\right)$ such that the state that realizes the supremum of the entropy functional $S\left[\Theta\left(\mathcal{N}_{A\rightarrow B}\right)\right]$ is an element of $\mathsf{FRank}\left(\widetilde{C}\otimes{C}\right)\cap\mathcal{D}\left(\widetilde{C}\otimes{C}\right)$. Again via a simple calculation we can show that $\left(\mathrm{id}_{\widetilde{C}}\otimes\Theta\left({R}_{A\rightarrow B}\right)\right)\outerproduct{\Phi}{\Phi}=\Phi_{\widetilde{C}}\otimes\mathbbm{1}_{D}$, where $\Phi_{\widetilde{C}}=\mathrm{tr}_{C}\left(\outerproduct{\Phi}{\Phi}\right)$. Since $\Phi_{\widetilde{C}}$ is full rank operator, we then have $\left(\mathrm{id}_{\widetilde{C}}\otimes\Theta\left(\mathcal{N}_{A\rightarrow B}\right)\right)\outerproduct{\Phi}{\Phi}=\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}$, where $\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}$ is a Choi state for quantum channel $\Theta\left(\mathcal{N}_{A\rightarrow B}\right)$. Thus, we can write the entropy of quantum channel $\Theta\left(\mathcal{N}_{A\rightarrow B}\right)$ as
$S\left[\Theta\left(\mathcal{N}_{A\rightarrow B}\right)\right]$ $=-D\left({\mathsf{C}}_{\Theta\left(\mathcal{N}\right)}^{\Phi}\;\middle\|\;\Phi_{\widetilde{C}}\otimes\mathbbm{1}_{D}\right)=S\left({\mathsf{C}}_{\Theta\left(\mathcal{N}\right)}^{\Phi}\right)-S\left(\Phi_{\widetilde{C}}\right).$ (147)
Now using Eqs. (146) and (147), the difference in the entropy of the quantum channel after and before applying the superchannel becomes
$S\left[\Theta\left(\mathcal{N}_{A\rightarrow B}\right)\right]-S\left[\mathcal{N}_{A\rightarrow B}\right]=\Delta\left[\mathsf{C}_{\Theta(\mathcal{N})}^{\Phi},\mathsf{C}_{\mathcal{N}}^{\Psi}\right]+\Delta^{\prime}\left[\Psi_{\widetilde{A}},\Phi_{\widetilde{C}}\right],$ (148)
where $\Delta\left[\mathsf{C}_{\Theta(\mathcal{N})}^{\Phi},\mathsf{C}_{\mathcal{N}}^{\Psi}\right]:=S\left(\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}\right)-S\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)$ and $\Delta^{\prime}\left[\Psi_{\widetilde{A}},\Phi_{\widetilde{C}}\right]:=S\left(\Psi_{\widetilde{A}}\right)-S\left(\Phi_{\widetilde{C}}\right)$.
In the following, we aim to find a lower bound for the entropy gain. One may calculate a lower bound on the functions $\Delta$ and $\Delta^{\prime}$ separately or together. Let $\mathfrak{T}$ be the representing for superchannel $\Theta$. Since Choi states $\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}$ and $\mathsf{C}_{\mathcal{N}}^{\Psi}$ are connected via the representing map $\mathfrak{T}$ (see Eq. (II) and Remark 3), we can find a lower bound on $\Delta\left[\mathsf{C}^{\Phi}_{\Theta\left(\mathcal{N}\right)},\mathsf{C}_{\mathcal{N}}^{\Psi}\right]$ in terms of relative entropy between Choi states $\mathsf{C}_{\mathcal{N}}^{\Psi}$ and $\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)_{\alpha}$ using Theorem 3, as follows:
$\Delta\left[\mathsf{C}_{\Theta(\mathcal{N})}^{\Phi},\mathsf{C}_{\mathcal{N}}^{\Psi}\right]$ $=S\left(\mathfrak{T}\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)\right)-S\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)\geq D\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\;\middle\|\;\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)_{\alpha}\right),$ (149)
where $\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)_{\alpha}=\frac{1}{\alpha^{\alpha}}\left(\mathfrak{T}^{*}\circ\mathfrak{T}\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)\right)^{\alpha}=\frac{1}{\alpha^{\alpha}}\left(\mathsf{C}^{\Psi}_{{\hat{\mathrm{\Theta}}_{\Psi}}^{-1}\circ\mathrm{\Theta}^{*}\circ\hat{\mathrm{\Theta}}_{\Phi}\circ\mathrm{\Theta}(\mathcal{N})}\right)^{\alpha}$ and $\alpha=\norm{\mathfrak{T}^{*}(\mathbbm{1})}_{\infty}=\norm{\mathsf{C}^{\Psi}_{{\hat{\mathrm{\Theta}}_{\Psi}}^{-1}\circ\mathrm{\Theta}^{*}\circ\hat{\mathrm{\Theta}}_{\Phi}\circ\mathrm{\Theta}^{-1}_{\Phi}(\mathcal{R})}}_{\infty}$ with $\mathcal{R}$ being the completely depolarizing map in $\mathcal{L}\left(C,D\right)$. The resulting inequality provides a lower bound of $\Delta\left[\mathsf{C}_{\Theta(\mathcal{N})}^{\Phi},\mathsf{C}_{\mathcal{N}}^{\Psi}\right]$. With this, we conclude that the following inequality holds
$S\left[\Theta\left(\mathcal{N}_{A\rightarrow B}\right)\right]-S\left[\mathcal{N}_{A\rightarrow B}\right]\geq D\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\;\middle\|\;\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)_{\alpha}\right)+\Delta^{\prime}\left[\Psi_{\widetilde{A}},\Phi_{\widetilde{C}}\right],$ (150)
which completes our proof.
Appendix D Refinement of data processing inequality
Let us consider a pair of quantum channels $\mathcal{N}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left(A^{\prime}\right)$ and $\mathcal{M}:\mathcal{L}\left({A}\right)\rightarrow\mathcal{L}\left(A^{\prime}\right)$ such that the state that realizes the supremum in the relative entropy functional $D\left[\mathcal{N}\|\mathcal{M}\right]$ is an element of $\mathsf{FRank}\left(\widetilde{A}\otimes{A}\right)\cap\mathcal{D}\left(\widetilde{A}\otimes{A}\right)$. Then we can write
$D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]$ $=\sup_{\rho_{\widetilde{A}A}}D\left(\mathrm{id}_{\widetilde{A}}\otimes\mathcal{N}\left(\rho_{\widetilde{A}A}\right)\;\middle\|\;\mathrm{id}_{\widetilde{A}}\otimes\mathcal{M}\left(\rho_{\widetilde{A}A}\right)\right)$
$=D\left(\left(\mathrm{id}_{\widetilde{A}}\otimes\mathcal{N}\right)\outerproduct{\Psi}{\Psi}\;\middle\|\;\left(\mathrm{id}_{\widetilde{A}}\otimes\mathcal{M}\right)\outerproduct{\Psi}{\Psi}\right),$ (151)
where |$\outerproduct{\Psi}{\Psi}\in\mathsf{FRank}\left(\widetilde{A}\otimes{A}\right)\cap\mathcal{D}\left(\widetilde{A}\otimes{A}\right)$, and $\mathrm{rank}(\mathrm{tr}_{A}\left(\outerproduct{\Psi}{\Psi}\right))=|\widetilde{A}|$ |. We then have $\left(\mathrm{id}_{\widetilde{A}}\otimes\mathcal{N}\right)\outerproduct{\Psi}{\Psi}={\mathsf{C}}_{\mathcal{N}}^{\Psi}$ and $\left(\mathrm{id}_{\widetilde{A}}\otimes\mathcal{N}\right)\outerproduct{\Psi}{\Psi}={\mathsf{C}}_{\mathcal{M}}^{\Psi}$, where ${\mathsf{C}}_{\mathcal{N}}^{\Psi}$ and ${\mathsf{C}}_{\mathcal{M}}^{\Psi}$ are Choi states for quantum channels $\mathcal{N}$ and $\mathcal{M}$, respectively. Therefore, in this case, we can write the relative entropy between quantum channels $\mathcal{N}$ and $\mathcal{M}$ as
$D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]=D\left({\mathsf{C}}_{\mathcal{N}}^{\Psi}\;\middle\|\;{\mathsf{C}}_{\mathcal{M}}^{\Psi}\right).$ (152)
Now, consider a superchannel $\Theta:\mathcal{L}\left(A,A^{\prime}\right)\rightarrow\mathcal{L}\left(C,C^{\prime}\right)$ such that the state that realizes the supremum in the relative entropy functional $D\left[\Theta(\mathcal{N})\;\middle\|\;\Theta(\mathcal{M})\right]$ is an element of $\mathsf{FRank}\left(\widetilde{C}\otimes{C}\right)\cap\mathcal{D}\left(\widetilde{C}\otimes{C}\right)$. In particular, let |$\ket{\Phi}$ be the state that realizes the supremum of the relative entropy $D\left[\Theta(\mathcal{N})\;\middle\|\;\Theta(\mathcal{M})\right]$. Since $\mathrm{rank}\left(\mathrm{tr}_{C}\left(\outerproduct{\Phi}{\Phi}\right)\right)=|\widetilde{C}|$ |, we have $\left(\mathrm{id}_{\widetilde{C}}\otimes\Theta\left(\mathcal{N}\right)\right)\outerproduct{\Phi}{\Phi}=\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}$ and $\left(\mathrm{id}_{\widetilde{C}}\otimes\Theta\left(\mathcal{M}\right)\right)\outerproduct{\Phi}{\Phi}=\mathsf{C}_{\Theta\left(\mathcal{M}\right)}^{\Phi}$, where $\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}$ and $\mathsf{C}_{\Theta\left(\mathcal{M}\right)}^{\Phi}$ are Choi states of channels $\Theta\left(\mathcal{N}\right)$ and $\Theta\left(\mathcal{M}\right)$, respectively. Thus, again, we can write the relative entropy between quantum channels $\Theta(\mathcal{N})$ and $\Theta(\mathcal{M})$ as
$D\left[\Theta(\mathcal{N})\;\middle\|\;\Theta(\mathcal{M})\right]=D\left({\mathsf{C}}_{\Theta(\mathcal{N})}^{\Phi}\;\middle\|\;{\mathsf{C}}_{\Theta(\mathcal{M})}^{\Phi}\right).$ (153)
Also, from the data processing inequality, we have $D\left[\mathcal{N}\|\mathcal{M}\right]\geq D\left[\Theta\left(\mathcal{N}\right)\|\Theta\left(\mathcal{M}\right)\right]$, which then implies the following inequality
$D\left({\mathsf{C}}_{\mathcal{N}}^{\Psi}\;\middle\|\;{\mathsf{C}}_{\mathcal{M}}^{\Psi}\right)\geq D\left({\mathsf{C}}_{\Theta(\mathcal{N})}^{\Phi}\;\middle\|\;{\mathsf{C}}_{\Theta(\mathcal{M})}^{\Phi}\right).$ (154)
Since, the Choi states $\mathsf{C}_{\mathcal{N}}^{\Psi}$, $\mathsf{C}_{\mathcal{M}}^{\Psi}$, $\mathsf{C}_{\Theta(\mathcal{N})}^{\Phi}$ and $\mathsf{C}_{\Theta(\mathcal{N})}^{\Phi}$ are connected via representing map $\mathfrak{T}$ of superchannel $\Theta$ as $\mathsf{C}_{\Theta\left(\mathcal{N}\right)}^{\Phi}=\mathfrak{T}\left(\mathsf{C}_{\mathcal{N}}^{\Psi}\right)$ and $\mathsf{C}_{\Theta\left(\mathcal{M}\right)}^{\Phi}=\mathfrak{T}\left(\mathsf{C}_{\mathcal{M}}^{\Psi}\right)$. Using this, we can rewrite the above entropic inequality as
$D\left({\mathsf{C}}_{\mathcal{N}}^{\Psi}\;\middle\|\;{\mathsf{C}}_{\mathcal{M}}^{\Psi}\right)\geq D\left(\mathfrak{T}\left({\mathsf{C}}_{\mathcal{N}}^{\Psi}\right)\;\middle\|\;\mathfrak{T}\left(\mathsf{C}_{\mathcal{M}}^{\Psi}\right)\right).$ (155)
Recall that the representing map $\mathfrak{T}$ of superchannel $\Theta$ is, in general, a CP map that becomes trace preserving if the adjoint of $\Theta$ becomes completely depolarising map preserving (see Lemma 6). But, in general, we can replace $\mathfrak{T}$ with the following trace preserving map: $\mathfrak{T}^{\prime}(\mathsf{X}):=\mathfrak{T}(\mathsf{X})+\left[\mathrm{tr}(\mathsf{X})-\mathrm{tr}(\mathfrak{T}(\mathsf{X}))\right]\sigma_{0}$, where $\mathsf{X}\in\mathcal{L}(C\otimes C^{\prime})$ and $\sigma_{0}\in\mathcal{L}(A\otimes A^{\prime})$ is chosen such that $\mathfrak{T}^{\prime}$ is a CPTP map, which further satisfy $\mathfrak{T}^{\prime}(\mathsf{C}^{\Psi}_{\mathcal{N}})=\mathfrak{T}(\mathsf{C}^{\Psi}_{\mathcal{N}})$ and $\mathsf{C}^{\Psi}_{\mathcal{M}}$ (see Appendix B.4). This is possible as $\Theta\in\mathsf{SCT}$. Thus, we have
$D\left({\mathsf{C}}_{\mathcal{N}}^{\Psi}\;\middle\|\;{\mathsf{C}}_{\mathcal{M}}^{\Psi}\right)\geq D\left(\mathfrak{T}^{\prime}\left({\mathsf{C}}_{\mathcal{N}}^{\Psi}\right)\;\middle\|\;\mathfrak{T}^{\prime}\left(\mathsf{C}_{\mathcal{M}}^{\Psi}\right)\right).$ (156)
Since $\mathfrak{T}^{\prime}$ is a quantum channel and Choi states are quantum states, we can write the refined version of the above inequality using the result (JRS+, 18, Theorem 2.1) as
$D\left({\mathsf{C}}_{\mathcal{N}}^{\Psi}\;\middle\|\;{\mathsf{C}}_{\mathcal{M}}^{\Psi}\right)-D\left(\mathfrak{T}\left({\mathsf{C}}_{\mathcal{N}}^{\Psi}\right)\;\middle\|\;\mathfrak{T}\left(\mathsf{C}_{\mathcal{M}}^{\Psi}\right)\right)\geq-\log F\left(\mathsf{C}^{\Psi}_{\mathcal{N}},\left(\mathcal{P}^{\mathrm{R}}\circ\mathfrak{T}^{\prime}\right)\mathsf{C}^{\Psi}_{\mathcal{N}}\right).$ (157)
Therefore, we have
$D\left[\mathcal{N}\;\middle\|\;\mathcal{M}\right]-D\left[\Theta\left(\mathcal{N}\right)\;\middle\|\;\Theta\left(\mathcal{M}\right)\right]\geq-\log F\left(\mathsf{C}^{\Psi}_{\mathcal{N}},\left(\mathcal{P}^{\mathrm{R}}\circ\mathfrak{T}^{\prime}\right)\mathsf{C}^{\Psi}_{\mathcal{N}}\right),$ (158)
which concludes our proof.
Appendix E Proof of monotonicity of $\mathbf{D}^{(n)}$
We restrict ourselves to the class of $(n+1)^{\text{th}}$ order processes which satisfy the following equation.
$\Phi^{(n+1)}\left(\Theta^{(n)}\right)=\Gamma^{(n)}_{\text{post}}\circ\left(\operatorname{id}\otimes~{}\Theta^{(n)}\right)\circ\Lambda^{(n)}_{\text{pre}}.$ (159)
For this class of $(n+1)^{\text{th}}$ order processes, our definition of generalized divergence between two $n^{\text{th}}$ order processes satisfies the monotonicity relation, Eq.(87). The proof goes as follows. Consider two $n^{\text{th}}$ order processes $\Theta^{(n)}$ and $\Gamma^{(n)}$ such that action of any $(n+1)^{\text{th}}$ order process, say $\Phi^{(n+1)}$, can be written in terms of concatenation of $n^{\text{th}}$ order processes, given by Eq. (159). Now, we show that in this scenario, data processing inequality holds, i.e., $\mathbf{D}^{(n)}\left[\Theta^{(n)}\;\middle\|\;\Gamma^{(n)}\right]\geq\mathbf{D}^{(n)}\left[\Phi^{(n+1)}\left(\Theta^{(n)}\right)\;\middle\|\;\Phi^{(n+1)}\left(\Gamma^{(n)}\right)\right]$. Using Eq. (86), we can write
$\mathbf{D}^{(n)}\left[\Phi^{(n+1)}\left(\Theta^{(n)}\right)\;\middle\|\;\Phi^{(n+1)}\left(\Gamma^{(n)}\right)\right]$
$=\sup_{\Omega^{(n-1)}\in\left(\mathcal{W}^{(n-1)}\otimes\mathcal{V}^{(n-1)}\right)_{+}}\mathbf{D}^{(n-1)}\left[\left(\mathrm{id}\otimes\Phi^{(n+1)}\left(\Theta^{(n)}\right)\right)\left(\Omega^{(n-1)}\right)\;\middle\|\;\left(\mathrm{id}\otimes\Phi^{(n+1)}\left(\Gamma^{(n)}\right)\right)\left(\Omega^{(n-1)}\right)\right].$ (160)
Using Eq. (159), we can write $\left(\mathrm{id}\otimes\Phi^{(n+1)}\left(\Theta^{(n)}\right)\right)\left(\Omega^{(n-1)}\right)=\left(\mathrm{id}\otimes\Gamma_{\text{post}}^{(n)}\right)\left(\mathrm{id}\otimes\left(\mathrm{id}\otimes\Theta^{(n)}\right)\right)\left(\widetilde{\Omega}^{(n-1)}\right)$, where we have defined $\widetilde{\Omega}^{(n-1)}:=\left(\mathrm{id}\otimes\Lambda_{\text{pre}}^{(n)}\right)\Omega^{(n-1)}$, and similarly we also have $\left(\mathrm{id}\otimes\Phi^{(n+1)}\left(\Gamma^{(n)}\right)\right)\left(\Omega^{(n-1)}\right)=\left(\mathrm{id}\otimes\Gamma_{\text{post}}^{(n)}\right)\left(\mathrm{id}\otimes\left(\mathrm{id}\otimes\Gamma^{(n)}\right)\right)\left(\widetilde{\Omega}^{(n-1)}\right)$. Now, using these on the right-hand side of Eq. (160), we obtain
$\mathbf{D}^{(n)}\left[\Phi^{(n+1)}\left(\Theta^{(n)}\right)\;\middle\|\;\Phi^{(n+1)}\left(\Gamma^{(n)}\right)\right]$
$\leq\sup_{\widetilde{\Omega}^{(n-1)}}\mathbf{D}^{(n-1)}\left[\left(\mathrm{id}\otimes\Gamma_{\text{post}}^{(n)}\right)\left(\mathrm{id}\otimes\left(\mathrm{id}\otimes\Theta^{(n)}\right)\right)\left(\widetilde{\Omega}^{(n-1)}\right)\;\middle\|\;\left(\mathrm{id}\otimes\Gamma_{\text{post}}^{(n)}\right)\left(\mathrm{id}\otimes\left(\mathrm{id}\otimes\Gamma^{(n)}\right)\right)\left(\widetilde{\Omega}^{(n-1)}\right)\right].$ (161)
In the above inequality, $\mathbf{D}^{(n-1)}\left[\cdot\;\middle\|\;\cdot\right]$ is a generalized divergence by our definition of generalized divergence of $n^{\text{th}}$ order process, thus using the fact that if $\Gamma_{\text{post}}^{(n)}$ is an $n^{\text{th}}$ order physical process then $\mathrm{id}\otimes\Gamma_{\text{post}}^{(n)}$ is also a valid $n^{\text{th}}$ order physical process, we have:
$\mathbf{D}^{(n-1)}\left[\left(\mathrm{id}\otimes\Gamma_{\text{post}}^{(n)}\right)\left(\mathrm{id}\otimes\left(\mathrm{id}\otimes\Theta^{(n)}\right)\right)\left(\widetilde{\Omega}^{(n-1)}\right)\;\middle\|\;\left(\mathrm{id}\otimes\Gamma_{\text{post}}^{(n)}\right)\left(\mathrm{id}\otimes\left(\mathrm{id}\otimes\Gamma^{(n)}\right)\right)\left(\widetilde{\Omega}^{n-1}\right)\right]$
$\leq\mathbf{D}^{(n-1)}\left[\left(\mathrm{id}\otimes\left(\mathrm{id}\otimes\Theta^{(n)}\right)\right)\left(\widetilde{\Omega}^{(n-1)}\right)\;\middle\|\;\left(\mathrm{id}\otimes\left(\mathrm{id}\otimes\Gamma^{(n)}\right)\right)\left(\widetilde{\Omega}^{(n-1)}\right)\right].$ (162)
Using this inequality in (161), we finally have
$\mathbf{D}^{(n)}\left[\Theta^{(n)}\;\middle\|\;\Gamma^{(n)}\right]\geq\mathbf{D}^{(n)}\left[\Phi^{(n+1)}\left(\Theta^{(n)}\right)\;\middle\|\;\Phi^{(n+1)}\left(\Gamma^{(n)}\right)\right].$